### v1

# The moduli spaces of left-invariant statistical structures on Lie groups

Abstract.
In the context of information geometry, the concept known as left-invariant statistical structure on Lie groups is defined by Furuhata–Inoguchi–Kobayashi [Inf. Geom. (2021)]. In this paper, we introduce the notion of the moduli space of left-invariant statistical structures on a Lie group. We study the moduli spaces for three particular Lie groups, each of which has a moduli space of left-invariant Riemannian metrics that is a singleton. As applications, we classify left-invariant conjugate symmetric statistical structures and left-invariant dually flat structures (which are equivalently left-invariant Hessian structures) on these three Lie groups. A characterization of the Amari–Chentsov $\alpha$-connections on the Takano Gaussian space is also given.
Key words and phrases:
statistical manifold; statistical Lie group; Hessian manifold; solvable Lie group
2020 Mathematics Subject Classification:
Primary 53C30; Secondary 53B12, 22E25, 53A15.
2020 Mathematics Subject Classification:
Primary 53C30, Secondary 53B12, 22E25, 53A15
The first author is supported by JST SPRING, Grant Number JPMJSP2132. The second author is supported by JST SPRING, Grant Number JPMJSP2119. The third author is supported by JSPS KAKENHI Grant Number JP24K06714. The fourth author is supported by JSPS KAKENHI Grant Numbers 23K22395, 24K21193. This work was partly supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165 and Osaka Metropolitan University Strategic Research Promotion Project (Development of International Research Hubs). This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
1. Introduction
In differential geometry, classifying geometric structures on manifolds is a fundamental problem. When the manifolds are equipped with additional geometric structures or suitable group actions, it becomes natural to consider compatible geometric structures. A particular area of interest is left-invariant geometric structures on Lie groups, which have emerged as a significant field of study. It is crucial to examine whether given Lie groups admit particular left-invariant geometric structures and, furthermore, to classify those structures on the given Lie groups. In this context, the concept of moduli spaces of left-invariant geometric structures plays a vital role, providing insight into their classification and deformation theory (e.g., [3, 5, 12, 18, 19, 21, 23, 32]).
In this paper, we focus on the statistical structures on manifolds (cf. [22, 24]). A statistical structure originates from information geometry, and is defined as a pair consisting of a Riemannian metric and a torsion-free affine connection that satisfies suitable conditions (see Section 2.2 for details). In the theory of statistical structures, two distinguished subclasses have received particular attention. The first is the class of dually flat structures, defined by the property that the associated affine connection is flat (cf. [1]). Dually flat structures play an important role in information geometry (see [1] for details). The second subclass is the class of conjugate symmetric statistical structures (cf. [24]), which is a broader concept that includes dually flat structures. This structure has significant connections with affine hypersurface theory (see [31]) and constant curvature statistical manifolds (see [16]). It is an important problem to examine whether given manifolds admit these structures, and to classify such structures on the manifolds.
A statistical structure on a Lie group is said to be left-invariant if both the metric and the connection are left-invariant (cf. [9, 14]). In this paper, we focus on this type of structure, and our concern is the following problem:
Problem A.
Given a Lie group $G$, find all left-invariant dually flat structures and left-invariant conjugate symmetric statistical structures on $G$.
Here, we review some previous works related to Problem A. We note that a dually flat structure is equivalent to a Hessian structure (see [35]). Shima [33] deeply studied this structure, and proved that left-invariant Hessian structures are admissible only on solvable Lie groups. Note that not every solvable Lie group admits a left-invariant Hessian structure, and a complete classification is still far from being achieved. For left-invariant conjugate symmetric statistical structures on Lie groups, a typical example is provided by the space $\mathcal{N}$ of univariate normal distributions, which can be identified with the Lie group $\mathbb{R}_{>0}\ltimes\mathbb{R}$. Furuhata–Inoguchi–Kobayashi [9] showed that the Amari–Chentsov $\alpha$-connections on the space $\mathcal{N}$ are the only left-invariant affine connections that are conjugate symmetric with respect to the Fisher metric on $\mathcal{N}$. Furthermore, Kobayashi and the second author [17] proved that a similar statement holds for the Lie group of multivariate normal distributions. Recently, Inoguchi and the second author [14] classified left-invariant conjugate symmetric statistical structures on $3$-dimensional Lie groups. Moreover, for the classification of a certain classes of left-invariant statistical structures on two- and three-dimensional Lie groups, see also [26]. For higher-dimensional Lie groups, aside from the Lie group of multivariate normal distributions studied in [17], the classification problem becomes more complicated and remains widely open.
In this paper, we introduce a new framework for investigating Problem A, which can be applied to Lie groups of any dimension. In the study of left-invariant Riemannian metrics on Lie groups, the notion of moduli spaces plays an important role. This concept was introduced for the Iwasawa manifold in [6] and further developed by Kodama, Takahara, and the third author in [18]. Consider the space of all left-invariant Riemannian metrics on a Lie group $G$, and the natural action of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ on it. The moduli space $\mathfrak{PM}(G)$ of left-invariant Riemannian metrics on $G$ is defined as the orbit space of this action. Inspired by this concept, let us denote by $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ the space of all left-invariant statistical structures on $G$, and consider the natural group action on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ by $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$. We define the moduli space of left-invariant statistical structures on $G$ as the orbit space of the action, and denoted by $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$.
We will see that two left-invariant statistical structures $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ belongs to an $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$-orbit if and only if there exists a scaling statistical isomorphism compatible with the group structure (see Definition 4.2.1) between statistical Lie groups $(G,g,\nabla)$ and $(G,g^{\prime},\nabla^{\prime})$. In particular, the action of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ preserves the properties of statistical structures, such as dually flatness and conjugate symmetry. Consequently, we define two subspaces of $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$, denoted by $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$, as the moduli spaces of left-invariant conjugate symmetric statistical structures and left-invariant dually flat structures, respectively. In this paper, motivated by Problem A, we address the following problem:
Problem B.
Given a Lie group $G$, determine the moduli spaces $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$.
For example, if $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)=\emptyset$, then the Lie group $G$ does not admit any left-invariant dually flat structures. If $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$ is a singleton, then the uniqueness of such structure holds. We note that $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)\neq\emptyset$ always holds, since the pair of a left-invariant Riemannian metric and its Levi-Civita connection is conjugate symmetric. Therefore, regarding $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$, we are interested in the existence of nontrivial left-invariant conjugate symmetric statistical structures.
It is natural to first consider Problem B for Lie groups $G$ where the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ is small. Recall that a left-invariant statistical structure on a Lie group $G$ is defined as a pair of a left-invariant Riemannian metric and a left-invariant affine connection. We are thus led to consider Lie groups $G$ whose moduli space $\mathfrak{PM}(G)$ of left-invariant Riemannian metrics is small, since these provide natural candidates for the study of Problem B. In fact, the simplest case has been classified. Specifically, for a connected and simply-connected Lie group, the moduli space $\mathfrak{PM}(G)$ is a singleton (that is, a left-invariant Riemannian metric is essentially unique) if and only if it is isomorphic to one of the following three series (cf. [23, 18]):
(1.1) $\mathbb{R}^{n},\quad G_{\mathbb{R}\mathrm{H}^{n}}\ (n\geq 2),\quad H^{3}\times\mathbb{R}^{n-3}\ (n\geq 3).$
Note that $\mathbb{R}^{n}$ is the abelian Lie group, $G_{\mathbb{R}\mathrm{H}^{n}}$ is the Lie group associated with the real hyperbolic space $\mathbb{R}\mathrm{H}^{n}$ (the solvable part of the Iwasawa decomposition of the identity component $SO^{0}(n,1)$ of $SO(n,1)$, and acts simply-transitively on $\mathbb{R}\mathrm{H}^{n}$), and $H^{3}$ is the three dimensional Heisenberg group. It is well-known that their unique metrics are flat on $\mathbb{R}^{n}$, negative constant sectional curvature on $G_{\mathbb{R}\mathrm{H}^{n}}$, and Ricci soliton on $H^{3}\times\mathbb{R}^{n-3}$, respectively.
The main result of this paper addresses Problem B for the three series of Lie groups shown in (1.1). For each of these Lie groups $G$, Table 1 summarizes the sets of representatives for the moduli spaces $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$.
For these Lie groups $G$, let us denote by $S^{3}(\mathfrak{g}^{\ast})$ the space of all symmetric $(0,3)$-tensors on the Lie algebra $\mathfrak{g}$ of $G$. Then the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ can be written as an orbit space of $S^{3}(\mathfrak{g}^{\ast})$ under a certain Lie group (see Section 4.3). A set of representatives for the moduli spaces $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$, resp. $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$, refers to a subset of $S^{3}(\mathfrak{g}^{\ast})$ whose projection onto these moduli spaces is surjective. If this projection is bijective, we call it a complete set of representatives. The notations $V^{+}$, $v_{0}$, and $w_{0}$ appearing in Table 1 will be explained in Theorems 1.0.1, 1.0.2, and 1.0.3, respectively.
Table 1. Sets of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$ in (1.1)
$G$ $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$
$\mathbb{R}^{n}$ $S^{3}(\mathfrak{g}^{\ast}_{\mathbb{R}^{n}})$ $V^{+}$
$G_{\mathbb{R}\mathrm{H}^{n}}$ $\mathbb{R}v_{0}$ $\{v_{0},-v_{0}\}$
$H^{3}$ $\{0\}$ $\emptyset$
$H^{3}\times\mathbb{R}^{n-3}$ $(\mathbb{R}w_{0}\odot S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast}))\oplus S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$ $\emptyset$
In order to express the moduli spaces, we need to fix the basis of the Lie algebras. For each of our three Lie algebras $\mathfrak{g}_{\mathbb{R}^{n}}$, $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ and $\mathfrak{h}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$, the standard basis $\{e_{1},e_{2},\ldots,e_{n}\}$ is a basis whose nonzero bracket relations are given as follows:
$\mathfrak{g}_{\mathbb{R}^{n}}$ $:\mbox{(none)},$
$\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ $:[e_{1},e_{i}]=e_{i}\quad(i=2,\ldots,n),$
$\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ $:[e_{1},e_{2}]=e_{3}.$
Let $\{x_{1},\ldots,x_{n}\}\subset\mathfrak{g}^{\ast}$ be the dual basis of the standard basis. We will now describe the results for each of the three Lie groups in detail. The first case is the abelian Lie group $\mathbb{R}^{n}$.
Theorem 1.0.1.
For the abelian Lie group $\mathbb{R}^{n}$, the following holds::
  • (1)
    All left-invariant statistical structures on $\mathbb{R}^{n}$ are conjugate symmetric.
  • (2)
    The following is a set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(\mathbb{R}^{n})$::
    $V^{+}:=\left\{\sum_{i=1}^{n}\lambda_{i}x_{i}^{3}~\middle|~\lambda_{1}\geq\dots\geq\lambda_{n}\geq 0\right\}.$
    The moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(\mathbb{R}^{n})$ is homeomorphic to the orbit space of the action of $\mathbb{R}_{>0}$ on $V^{+}$, which implies the existence of an (n1)(n-1)-dimensional family of left-invariant dually flat structures on $\mathbb{R}^{n}$.
The second case is the Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ of the real hyperbolic space.
Theorem 1.0.2.
For the Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$, we denote by
$v_{0}:=4x_{1}^{3}+6\sum_{i=2}^{n}x_{1}x_{i}^{2}\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast}).$
  • (1)
    $\mathbb{R}v_{0}$ is a complete set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G_{\mathbb{R}\mathrm{H}^{n}})$. Specifically, the moduli space, equipped with the natural quotient topology, is homeomorphic to the subspace $\mathbb{R}v_{0}$ of $S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})$. This implies that there exists a one-dimensional family of left-invariant conjugate symmetric statistical structures on $G_{\mathbb{R}\mathrm{H}^{n}}$.
  • (2)
    $\{v_{0},-v_{0}\}$ is a complete set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G_{\mathbb{R}\mathrm{H}^{n}})$. Specifically, the moduli space, equipped with the natural quotient topology, is homeomorphic to a two-point discrete space. This implies that there exist exactly two left-invariant dually flat structures on $G_{\mathbb{R}\mathrm{H}^{n}}$ up to scaling statistical isomorphisms compatible with group structure.
The third case is the Lie group $H^{3}\times\mathbb{R}^{n-3}$, the direct product of the 3-dimensional Heisenberg group and an abelian Lie group. We denote by $\odot$ the symmetric tensor product.
Theorem 1.0.3.
For the Lie group $H^{3}\times\mathbb{R}^{n-3}$, we denote by
$w_{0}:=x_{1}x_{1}+x_{2}x_{2}+x_{3}x_{3}\in S^{2}(\mathfrak{h}_{3}^{\ast}).$
  • (1)
    $(\mathbb{R}w_{0}\odot S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast}))\oplus S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$ is a set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(H^{3}\times\mathbb{R}^{n-3})$. In particular, if $n=3$, then the set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(H^{3})$ is a singleton, which implies that a left-invariant conjugate symmetric statistical structure on $H^{3}$ is essentially unique; this structure consists of a left-invariant Riemannian metric and its Levi-Civita connection. If $n\geq 4$, then the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(H^{3}\times\mathbb{R}^{n-3})$ is homeomorphic to the orbit space of $(\mathbb{R}w_{0}\odot S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast}))\oplus S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$ under the action of O(n3)O(n-3).
  • (2)
    There do not exist any left-invariant dually flat structures on $H^{3}\times\mathbb{R}^{n-3}$.
We note that the statement regarding $H^{3}$ has been established in [14]. Our argument provides an alternative approach and generalizes this result to the higher-dimensional case. In particular, unlike the case of $n=3$, the Lie group $H^{3}\times\mathbb{R}^{n-3}$ with $n\geq 4$ admits nontrivial left-invariant conjugate symmetric statistical structures.
Finally, in this section, we present one application of our theorems. By using an explicit expression for the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G_{\mathbb{R}\mathrm{H}^{n}})$ of left-invariant conjugate symmetric statistical structures on $G_{\mathbb{R}\mathrm{H}^{n}}$, we obtained a characterization of the Amari–Chentsov $\alpha$-connections on the Takano Gaussian space. This result is analogous to the characterization of the Amari–Chentsov $\alpha$-connections on the family of univariate normal distributions (cf. [9]), and on the family of multivariate normal distributions (cf. [17]).
Corollary 1.0.4 (see also Theorem 6.3.1).
The Amari–Chentsov $\alpha$-connection $\nabla$ on the Takano Gaussian space $(\mathcal{N}_{T},g^{F})$ is characterized by the following two properties:
  • (1)
    The statistical structure $(g^{F},\nabla)$ is conjugate symmetric.
  • (2)
    $\nabla$ is $\mathbb{R}_{>0}\ltimes\mathbb{R}^{n}$-invariant.
This paper is organized as follows: In Section 2, we briefly recall the basic facts of statistical manifolds and introduce several important classes and examples of statistical structures. We begin in Section 3 with a study of left-invariant statistical structures on Lie groups. Section 3.3 contains several computational formulas that will be used later, in Section 5 and the subsequent sections. The notion of the moduli space of left-invariant statistical structures is then introduced in Section 4. Finally, Sections 5, 6, and 7 present classification results together with the corresponding moduli spaces for the Lie groups listed in (1.1).
2. Preliminaries
Throughout this section, we fix a smooth manifold $M$.
2.1. Notations
In this subsection, we fix the terminology used throughout this paper.
For each affine connection $\nabla$ on $M$, we use the symbols $R^{\nabla}$ and $T^{\nabla}$ for the curvature tensor field of type $(1,3)$ and the torsion tensor field of type $(1,2)$ of $\nabla$, respectively, that is,
$R^{\nabla}(X,Y)Z$ $:=(\nabla_{X}\nabla_{Y}-\nabla_{Y}\nabla_{X}-\nabla_{[X,Y]})Z,$
$T^{\nabla}(X,Y)$ $:=\nabla_{X}Y-\nabla_{Y}X-[X,Y].$
Let us fix a Riemannian metric $g$ and an affine connection $\nabla$ on $M$. We write $\nabla^{g}$ for the Levi-Civita connection of $g$ on $M$. The dual connection of $\nabla$ with respect to $g$ will be denoted by $\overline{\nabla}$, that is,
$Xg(Y,Z)=g(\nabla_{X}Y,Z)+g(Y,\overline{\nabla}_{X}Z).$
The following relation between $\nabla$ and $\overline{\nabla}$ is well-known.
Proposition 2.1.1 ([25]).
If two of the following four conditions are satisfied, then the remaining two will follow:
  1. (1)
    $T^{\nabla}\equiv 0$,
  2. (2)
    $T^{\overline{\nabla}}\equiv 0$,
  3. (3)
    $(\nabla+\overline{\nabla})/2=\nabla^{g}$,
  4. (4)
    A $(0,3)$-tensor field $\nabla g$ be totally symmetric.
The following formula is also well-known (see [28]):
(2.1) $g(R^{\nabla}(X,Y)Z,W)+g(R^{\overline{\nabla}}(X,Y)W,Z)=0.$
In particular, we have
(2.2) $R^{\nabla}\equiv 0\ \iff\ R^{\overline{\nabla}}\equiv 0.$
2.2. Statistical manifolds
In this subsection, we recall some basic facts about statistical manifolds, see [1, 24] for details.
In this paper, we say that a triple $(M,g,\nabla)$ is called a statistical manifold if $g$ is a Riemannian metric and $\nabla$ is an affine connection on $M$ satisfying all conditions in Proposition 2.1.1. For a statistical manifold $(M,g,\nabla)$, the pair $(g,\nabla)$ is called a statistical structure on $M$. In particular, $\nabla$ is called a statistical connection on the Riemannian manifold $(M,g)$.
For each Riemannian metric $g$ on $M$, the pair $(g,\nabla^{g})$ is a statistical structure on $M$ by the definition of the Levi-Civita connection, and $(g,\nabla^{g})$ is called the trivial statistical structure (see [8]). Thus, statistical manifolds can be considered as a generalization of Riemannian manifolds.
For each statistical structure $(g,\nabla)$, the pair $(g,\overline{\nabla})$ is also a statistical structure on $M$. The pair $(g,\overline{\nabla})$ is called the dual statistical structure of the statistical structure $(g,\nabla)$, and the triple $(M,g,\overline{\nabla})$ is called the dual statistical manifold of $(M,g,\nabla)$.
We define a certain equivalence relation between statistical manifolds as follows.
Definition 2.2.1.
Let $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ be statistical manifolds. If there exists a diffeomorphism $f\vcentcolon M\to M^{\prime}$ and a positive real number $r>0$ satisfying the following two conditions (1) and (2), then $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ are said to be statistically isomorphic up to scaling by $(f,r)$:
  1. (1)
    $f^{\ast}\nabla^{\prime}=\nabla$,
  2. (2)
    $f^{\ast}g^{\prime}=r\cdot g$.
And then, the pair $(f,r)$ is called a scaling statistical isomorphism between statistical manifolds $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$. In particular, in the case $r=1$, the map $f$ is called a statistical isomorphism from $(M,g,\nabla)$ to $(M^{\prime},g^{\prime},\nabla^{\prime})$, and we say that $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ are statistically isomorphic to each other.
The following proposition will be applied in Theorem 2.3.4.
Proposition 2.2.2.
Let $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ be statistical manifolds that are statistically isomorphic up to scaling by $(f,r)$. Then the following equality holds:
$f^{\ast}\overline{\nabla^{\prime}}=\overline{\nabla}.$
Next, we introduce the difference tensor and the cubic form, which are equivalent to statistical connections on Riemannian manifolds. Let $(g,\nabla)$ be a statistical structure on $M$. We define the $(1,2)$-tensor field $K$ by $K(X,Y)=\nabla_{X}Y-\nabla^{g}_{X}Y$. This tensor field $K$ is called the difference tensor of $(g,\nabla)$. Since the affine connections $\nabla$ and $\nabla^{g}$ are both torsion-free, the difference tensor $K$ is symmetric, that is, $K(X,Y)=K(Y,X)$. We define the $(0,3)$-tensor field $C$ by $C(X,Y,Z)=(\nabla_{X}g)(Y,Z)$. This tensor field $C$ is called the cubic form of $(g,\nabla)$. Note that C(X,Y,Z)=2g(K(X,Y),Z)C(X,Y,Z)=-2g(K(X,Y),Z) holds.
We introduce the following notions. Let $\mathcal{A}_{\mathrm{Stat}}(M,g)$ be the set of all statistical connections on the Riemannian manifold $(M,g)$, and $\mathcal{K}(M,g)$ the set of all symmetric $(1,2)$-tensor fields on $M$ satisfying the following equalities,
$g(K(X,Y),Z)=g(K(Y,Z),X)=g(K(Z,X),Y).$
In addition, we denote by $S^{3}(T^{\ast}M)$ the set of all symmetric $(0,3)$-tensor fields on $M$. The following fact is folklore known.
Proposition 2.2.3.
The following maps are both bijective:
$\mathcal{A}_{\mathrm{Stat}}(M,g)\ni\nabla$ $\longmapsto K^{(g,\nabla)}\in\mathcal{K}(M,g),$
$\mathcal{K}(M,g)\ni K$ $\longmapsto(\nabla^{g}+K)g\in S^{3}(T^{\ast}M).$
Here, $K^{(g,\nabla)}$ is the difference tensor of $(g,\nabla)$. The composition of the above maps can be expressed as follows:
$\mathcal{A}_{\mathrm{Stat}}(M,g)\ni\nabla\longmapsto\nabla g\in S^{3}(T^{\ast}M).$
By Proposition 2.2.3, a pair consisting of the Riemannian metric $g$ and a cubic form $C\in S^{3}(T^{\ast}M)$, or a pair consisting of $g$ and a difference tensor $K\in\mathcal{K}(M,g)$, can also be called a statistical structure on $M$. Thus, we will also call $(g,C)$ and $(g,K)$ statistical structures on $M$ as well. We see that the statistical structure $(g,0)$ on $M$ corresponds to the statistical structure $(g,\nabla^{g})$. Here 0 is a origin of $\mathcal{K}(M,g)$ or $S^{3}(T^{\ast}M)$ as the liner spaces.
We note that if two statistical manifolds $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ are statistically isomorphic up to scaling by $(f,r)$, then the following equations hold:
$f^{\ast}K^{\prime}=K,\quad f^{\ast}C^{\prime}=r\cdot C.$
Here, $K$, $C$, $K^{\prime}$, and $C^{\prime}$ are the difference tensors and cubic forms on $M$ and $M^{\prime}$, respectively.
Throughout this paper, for each $K\in\mathcal{K}(M,g)$ and each vector field $X$ on $M$, we define the $(1,1)$-tensor field $K_{X}$ on $M$ by putting $K_{X}(Y):=K(X,Y)$ for each vector field $Y$. For a pair $X,Y$ of vector fields on $M$, the Lie bracket $K_{X},K_{Y}]$ of $K_{X}$ and $K_{Y}$ is defined as $K_{X},K_{Y}](Z)=K_{X}(K_{Y}(Z))-K_{Y}(K_{X}(Z))$. The following proposition, proved by Opozda [30, Corollary 3.8], will be applied in Section 5.2:
Proposition 2.2.4.
Fix a statistical structure $(g,\nabla)$ and a point $p\in M$. We denote by $K$ the difference tensor of $(g,\nabla)$. Then the following two conditions on $(g,\nabla,p)$ are equivalent:
  1. (1)
    $K_{X},K_{Y}]_{p}=0$ for any pair $(X,Y)$ of vector fields on $M$.
  2. (2)
    There exists an orthonormal basis $e_{1},\dots,e_{n}$ of $(T_{p}M,g_{p})$ and $c_{1},\dots c_{n}\in\mathbb{R}$ such that
    $K_{p}(e_{i},e_{i})=c_{i}e_{i},\quad K_{p}(e_{i},e_{j})=0$
    for distinct $i,j=1,\dots,n$.
Let $(M_{i},g_{i},\nabla_{i})$ be statistical manifolds for $i=1,2$. The direct product manifold $M_{1}\times M_{2}$ carries a natural statistical structure given by the direct product metric $g_{1}\oplus g_{2}$ and the direct product connection $\nabla_{1}\oplus\nabla_{2}$. The cubic form $C$ of this statistical structure is given by
$C(X,Y,Z)=C_{1}(X_{1},Y_{1},Z_{1})+C_{2}(X_{2},Y_{2},Z_{2}),$
for all vector fields $X,Y,Z$ on $M_{1}\times M_{2}$, where $C_{i}$ denotes the cubic form of $(M_{i},g_{i},\nabla_{i})$, and $X_{i}$ denotes the $M_{i}$-component of the vector field $X$ $(i=1,2)$.
2.3. Some classes of statistical structures
In this subsection, we introduce four classes of statistical structures: conjugate symmetric, constant curvature, dually flat, and constant Hessian curvature (abbreviated as CHC). Note that these classes are nested in the order listed above; that is, every CHC structure is dually flat, every dually flat structure has constant curvature, and every constant curvature statistical structure is conjugate symmetric.
In [24], Lauritzen defined a statistical manifold $(M,g,\nabla)$ to be conjugate symmetric if $R^{\nabla}\equiv R^{\overline{\nabla}}$. We also call a statistical structure $(g,\nabla)$ conjugate symmetric if the statistical manifold $(M,g,\nabla)$ is conjugate symmetric. Note that for each Riemannian metric $g$ on a manifold $M$, the trivial statistical structure $(g,\nabla^{g})$ is clearly conjugate symmetric. Moreover, the property of being conjugate symmetric is closed under taking direct products in the sense of Section 2.2.
The conjugate symmetry of a statistical structure can be characterized as follows:
Proposition 2.3.1 (Proposition 2.4 in [16]).
Let $(g,\nabla)$ be a statistical structure on $M$, and let $K$ and $C$ be the difference tensor and cubic form of $(g,\nabla)$ respectively. Then, the following four conditions are equivalent:
  1. (i)
    The statistical structure $(g,\nabla)$ is conjugate symmetric.
  2. (ii)
    A $(0,4)$-tensor field $\nabla C$ is totally symmetric.
  3. (iii)
    A $(0,4)$-tensor field $\nabla^{g}C$ is totally symmetric.
  4. (iv)
    A $(1,3)$-tensor field $\nabla^{g}K$ is totally symmetric, that is,
    $(\nabla^{g}_{X_{1}}K)(X_{2},X_{3})=(\nabla^{g}_{X_{\sigma(1)}}K)(X_{\sigma(2)},X_{\sigma(3)})$
    for vector fields $X_{1},X_{2},X_{3}$ and $\sigma\in\mathfrak{S}_{3}$.
Note that, by Proposition 2.3.1, the following set is a linear subspace of $S^{3}(T^{\ast}M)$:
(2.3) $\left\{C\in S^{3}(T^{\ast}M)\ |\ (g,C)\text{ is conjugate symmetric}\right\}.$
Throughout this paper, for a conjugate symmetric statistical manifold $(M,g,\nabla)$, the $(0,4)$-tensor field $R^{\nabla}_{g}$ is defined by putting
$R^{\nabla}_{g}(X,Y,Z,W):=g(R^{\nabla}(X,Y)Z,W)$
for vector fields $X,Y,Z,W$ on $M$. Then the following holds:
Proposition 2.3.2 (cf. [29]).
For a conjugate symmetric statistical manifold $(M,g,\nabla)$, the equality below holds:
(2.4) $R^{\nabla}(X,Y)=R^{\nabla^{g}}(X,Y)+[K_{X},K_{Y}]$
for all vector fields $X$, $Y$ on $M$. In particular, the $(0,4)$-tensor field $R^{\nabla}_{g}$ satisfies
$R^{\nabla}_{g}(X,Y,Z,W)=-R^{\nabla}_{g}(Y,X,Z,W)=-R^{\nabla}_{g}(X,Y,W,Z)=R^{\nabla}_{g}(Z,W,X,Y)$
and
$R^{\nabla}_{g}(X,Y,Z,W)+R^{\nabla}_{g}(Y,Z,X,W)+R^{\nabla}_{g}(Z,X,Y,W)=0$
for vector fields $X,Y,Z,W$ on $M$.
For conjugate symmetric statistical manifolds, “sectional curvatures of the statistical structure” can be defined in the natural sense as below: Let $(M,g,\nabla)$ be a conjugate symmetric statistical manifold and fix a point $p$. For each linearly independent pair $\{v,w\}$ of vectors in $T_{p}M$, we put
$\mathrm{Sect}^{\nabla}_{g}(v,w):=\frac{(R^{\nabla}_{g})_{p}(v,w,w,v)}{g_{p}(v,v)g_{p}(w,w)-g_{p}(v,w)^{2}}.$
Then $\mathrm{Sect}^{\nabla}_{g}(v,w)$ is known to depend only on $\mathrm{Span}\{v,w\}$. Thus $\mathrm{Sect}^{\nabla}_{g}(v,w)$ is called the sectional curvature of the $2$-dimensional tangential plane $\mathrm{Span}\{v,w\}\subset T_{p}M$ on $(M,g,\nabla)$ at $p$ (see [24] for the details).
In this paper, a statistical manifold $(M,g,\nabla)$ is said to have constant curvature if there exists a constant $k\in\mathbb{R}$ such that
$R^{\nabla}(X,Y)Z=k(g(Y,Z)X-g(X,Z)Y)$
(see [22]) or equivalently, $(M,g,\nabla)$ is conjugate symmetric and $\mathrm{Sect}^{\nabla}_{g}(v,w)=k$ for any $p\in M$ and any linearly independent pair $(v,w)$ of vectors in $T_{p}M$.
Remark 2.3.3.
Opozda [29] gave a definition of sectional curvatures for general statistical manifolds.
A statistical manifold $(M,g,\nabla)$ is called a dually flat statistical manifold, or simply dually flat, when $R^{\nabla}=0$ (or equivalently, $R^{\overline{\nabla}}=0$ by Equation (2.2)) holds. For a dually flat statistical manifold $(M,g,\nabla)$, we call $(g,\nabla)$ dually flat structure on $M$. Note that dually flatness is preserved under taking direct products, in the sense of Section 2.2 (see [38]). As is well-known, the dually flat statistical manifold and the dually flat structures are equivalent to the notion of Hessian manifold and Hessian structures, respectively, see [35] for the details. The dually flat structures play an important role in information geometry, see for examples [1].
Let $(M,g,\nabla)$ be a Hessian manifold (i.e. a dually flat statistical manifold). A Hessian structure $(g,\nabla)$ is said to be of CHC (of constant Hessian sectional curvature) if there exists a constant $c\in\mathbb{R}$ such that the equation
$(\nabla_{X}K)(Y,Z)=-\frac{c}{2}\{g(X,Y)Z+g(X,Z)Y\}$
holds for any vector fields $X,Y,Z$ on $M$, and then $(g,\nabla)$ is also called of CHC $c$ (cf. [10, 34]). In [10], Furuhata and Kurose gave a classification of connected and simply-connected Hessian manifolds of non-positive CHC.
It should be emphasized that by the arguments in Section 2.2, the following holds.
Theorem 2.3.4.
The classes of conjugate symmetric, constant curvature, dually flat, and CHC statistical manifolds are preserved under scaling statistical isomorphisms.
2.4. Some examples of statistical manifolds
In this subsection, we introduce some examples of statistical manifolds.
Example 2.4.1 ($n$-variable normal distribution family).
Let $n\in\mathbb{Z}$ with $n\geq 1$, and $\mathrm{Sym}^{+}(n,\mathbb{R})$ be the space of all positive definite symmetric matrices of order $n$. The $n$-variate normal distribution family is a statistical model $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$ with $\mathcal{N}^{n}=\mathrm{Sym}^{+}(n,\mathbb{R})\times\mathbb{R}^{n}$ and
[…]
(see [4] for the definition of statistical model). Here, $dx$ denotes the Lebesgue measure on $\mathbb{R}^{n}$, and $\mathcal{P}(\mathbb{R}^{n})$ denotes the space of all probability measures on $\mathbb{R}^{n}$. The Fisher metric on $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$ is a Riemannian metric on $\mathcal{N}^{n}$ which has the form
(2.5) $g^{F}_{\hat{\theta}}(X,Y)=\int_{x\in\mathbb{R}^{n}}(X\log N(x\mid\theta))(Y\log N(x\mid\theta))N(x\mid\hat{\theta})dx\quad(\hat{\theta}\in\mathcal{N}^{n}).$
For $\alpha\in\mathbb{R}$, the Amari–Chentsov $\alpha$-tensor field on $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$ is a symmetric $(0,3)$-tensor field on $\mathcal{N}^{n}$ which has the form
(2.6) $C^{A(\alpha)}_{\hat{\theta}}(X,Y,Z)=\alpha\int_{x\in\mathbb{R}^{n}}(X\log N(x\mid\theta))(Y\log N(x\mid\theta))(Z\log N(x\mid\theta))N(x\mid\hat{\theta})dx\quad(\hat{\theta}\in\mathcal{N}^{n}).$
Moreover, a statistical connection $\nabla^{A(\alpha)}$ on $(\mathcal{N}^{n},g^{F})$ correspondence to $(g^{F},C^{A(\alpha)})$ in the sense of Proposition 2.2.3 is called the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$.
The following fact for the statistical manifold $(\mathcal{N}^{n},g^{F},C^{A(\alpha)})$ is well-known.
Fact 2.4.2 ([2, 24]).
For any $\alpha\in\mathbb{R}$, the statistical manifold $(\mathcal{N}^{n},g^{F},\nabla^{A(\alpha)})$ is conjugate symmetric. Moreover, it is dually flat if and only if $\alpha=\pm 1$.
Example 2.4.3 (Takano Gaussian space).
Let $n\in\mathbb{Z}$ with $n\geq 1$, $\mathcal{N}_{T}^{n}:=\mathbb{R}_{>0}\times\mathbb{R}^{n}$ and
$\mathbf{p}_{T}=\mathbf{p}_{N}\circ\iota:\mathcal{N}^{n}_{T}\to\mathcal{P}(\mathbb{R}^{n}),\ \ (\Sigma,\mu)\mapsto N(x\mid\mathrm{diag}(\Sigma,\dots,\Sigma),\mu)dx.$
Here,
$\iota:\mathcal{N}^{n}_{T}\hookrightarrow\mathcal{N}^{n},\ \ (\Sigma,\mu)\mapsto(\mathrm{diag}(\Sigma,\dots,\Sigma),\mu).$
Then the triple $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$ is a statistical model in the sense of [4]. On $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$, the Fisher metric $g^{F}$, the Amari–Chentsov $\alpha$-tensor field $C^{A(\alpha)}$, and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ are defined in the similar way as in Example 2.4.1, respectively. In [7], the Riemannian manifold $(\mathcal{N}_{T}^{n},g^{F})$ is called the Takano Gaussian space, and we use this terminology in this paper as well.
Remark 2.4.4.
For $n=1$, the statistical model $(\mathcal{N}_{T}^{1},\mathbb{R},\mathbf{p}_{T})$ coincides with the univariate normal distribution family $(\mathcal{N}^{1},\mathbb{R},\mathbf{p}_{\mathrm{N}})$.
For the statistical manifolds $(\mathcal{N}^{n}_{T},g^{F},\nabla^{A(\alpha)})$, the following holds.
Proposition 2.4.5.
For any $\alpha\in\mathbb{R}$, the statistical manifold $(\mathcal{N}^{n}_{T},g^{F},C^{A(\alpha)})$ has constant curvature (cf. [36]). Moreover, it is dually flat if and only if $\alpha=\pm 1$ (cf. [2, 24]).
For other examples of statistical manifolds obtained from statistical models, see [1, 24] for example.
Remark 2.4.6.
The statistical structure on $\mathcal{N}^{n}$ defined above is related to statistical inference on the multivariate normal distribution family $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$ (see [1] for details).
3. Preliminaries for left-invariant geometric structures on Lie groups
The purpose of this section is to fix our terminologies for left-invariant Riemannian metrics and left-invariant statistical structures on Lie groups.
3.1. Homogeneous statistical manifolds
Let $(g,\nabla)$ be a statistical structure on a manifold $M$ equipped with a smooth action of a Lie group $G$. We say that $(g,\nabla)$ is $G$-invariant if $g$ and $\nabla$ are both $G$-invariant. In particular, when $M$ is $G$-homogeneous space, the triple $(M,g,\nabla)$ is called the $G$-homogeneous statistical manifold.
We shall introduce some examples of homogeneous statistical manifolds obtained from statistical models. We use the notions introduced in Section 2.4.
Example 3.1.1.
The statistical manifold $(\mathcal{N}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.4.1) is an $\mathrm{Aff}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$. Note that the isotropy subgroup of the $\mathrm{Aff}(n,\mathbb{R})$-action on $\mathcal{N}^{n}$ at the point $(I_{n},0)$ is $O(n)\ltimes\{0\}$. Here, $I_{n}$ denotes the identity matrix of size $n$.
Example 3.1.2.
Let $\mathrm{Aff}^{d+}(n,\mathbb{R})\coloneqq\mathbb{R}_{>0}\ltimes\mathbb{R}^{n}$. The statistical manifold $(\mathcal{N}_{T}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.4.3) is an $\mathrm{Aff}^{d+}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$. Note that the $\mathrm{Aff}^{d+}(n,\mathbb{R})$-action on $\mathcal{N}_{T}^{n}$ is simply-transitive.
The nontrivial point in the above two examples is the invariance of the Fisher metric and the Amari–Chentsov $\alpha$-connection. As the invariance of the Fisher metric and the Amari–Chentsov $\alpha$-tensor field on a statistical model, it is described in detail in [4] (see the generalization of Chentsov’s theorem, Corollary 5.3 in [4]).
3.2. Left-invariant Riemannian metrics and left-invariant statistical structures on Lie groups
Let $G$ be a Lie group. We consider $G$ is a $G$-homogeneous space equipped with the left translations. In this subsection, we introduce some basic facts about the left-invariant statistical structures on $G$.
A left-invariant Riemannian metric on $G$ is a Riemannian metric that is invariant under left translations. Let $\mathfrak{g}$ denote the Lie algebra of $G$, that is, the Lie algebra consisting of all left-invariant vector fields on $G$. As is well-known that the space $\mathop{\widetilde{\mathfrak{M}}}(G)$ of all left-invariant Riemannian metrics on $G$ can be regarded as the space $\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ of all inner products on the vector space $\mathfrak{g}$ in a canonical sense (see [15] for the details). Throughout this paper, the space $\mathop{\widetilde{\mathfrak{M}}}(G)\cong\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ is considered as a smooth manifold in the usual sense. In particular, for each $\langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$, we have an diffeomorphism $\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})\cong GL(\mathfrak{g})/{O(\mathfrak{g},\langle,\rangle)}$, where $GL(\mathfrak{g})$ denotes the general linear group of the vector space $\mathfrak{g}$ and $O(\mathfrak{g},\langle,\rangle)$ the orthogonal group of the inner product space $(\mathfrak{g},\langle,\rangle)$. Note that such manifold $\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ can be considered as a symmetric space.
A left-invariant affine connection on a Lie group $G$ is an affine connection that is invariant under left translations. A statistical structure $(g,\nabla)$ on $G$ is said to be left-invariant if $g$ and $\nabla$ are both left-invariant. For such $(g,\nabla)$, the triple $(G,g,\nabla)$ is called a statistical Lie group, see [9].
Throughout this paper the set of all left-invariant statistic structures on $G$ will be denoted by
$\mathop{\mathcal{L}\mathrm{Stat}}(G):=\left\{\text{ left-invariant statistic structures }(g,\nabla)\text{ on }G\right\}.$
Throughout this paper, we write $S^{k}(\mathfrak{g}^{\ast})=\overbrace{\mathfrak{g}^{*}\odot\dots\odot\mathfrak{g}^{*}}^{k}$ for the $k$-th order symmetric tensors of the dual space $\mathfrak{g}^{\ast}$ of $\mathfrak{g}$. For $x_{1},\dots,x_{k}\in\mathfrak{g}^{\ast}$, we realize the element $x_{1}x_{2}\cdots x_{k}\in S^{k}(\mathfrak{g}^{\ast})$ in the following sense:
$(x_{1}x_{2}\cdots x_{k})(X_{1},X_{2},\cdots,X_{k}):=\frac{1}{k!}\sum_{\sigma\in\mathfrak{S}_{k}}(x_{\sigma(1)}(X_{1}))\cdot(x_{\sigma(2)}(X_{2}))\cdots(x_{\sigma(k)}(X_{k}))\quad(X_{1},\dots,X_{k}\in\mathfrak{g}).$
In particular, if we put $n$ the dimension of $G$ and take a basis $\{x_{1},\dots,x_{n}\}$ of $\mathfrak{g}^{\ast}$, the sets $\{x_{i_{1}}x_{i_{2}}\cdots x_{i_{k}}\mid 1\leq i_{1}\leq i_{2}\leq\cdots\leq i_{k}\leq n\}$ forms a basis of $S^{k}(\mathfrak{g}^{\ast})$, and thus the real vector space $S^{k}(\mathfrak{g}^{\ast})$ is $\binom{n+k-1}{k}-$dimensional. In the following sections, we shall often identify $S^{k}(\mathfrak{g}^{\ast})$ with the vector space $\mathrm{Homog}_{k}(x_{1},\dots,x_{n})$ of homogeneous polynomials of degree $k$.
We note that for $k=3$, the space $S^{3}(\mathfrak{g}^{\ast})$ is nothing but the space of all left-invariant cubic forms on $G$.
For each $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$, we also define
  • $\mathcal{LA}_{\mathrm{Stat}}(G,g):=\{\nabla\in\mathcal{A}_{\mathrm{stat}}(G,g)\mid\text{$\nabla$ is left-invariant}\}\subset\mathcal{A}_{\mathrm{stat}}(G,g)$ and
  • $\mathcal{LK}(G,g):=\{K\in\mathcal{K}(G,g)\mid\text{$K$ is left-invariant}\}\subset\mathcal{K}(G,g)$.
For each $(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)$, we denote by $K^{(g,\nabla)}$ the difference tensor of $(g,\nabla)$. Furthermore, for $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$ and $K\in\mathcal{LK}(G,g)$, let us write $C^{(g,K)}$ for the cubic form of $(g,K)$ in the sense of $C^{(g,K)}(X,Y,Z):=-2g(K(X,Y),Z)$ as in Section 2.2.
We also note that for a fixed $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$, the space $\mathcal{LK}(G,g)$ is identified with the space $\mathcal{K}(\mathfrak{g},g)$ consists of the symmetric bilinear form $K:\mathfrak{g}\odot\mathfrak{g}\rightarrow\mathfrak{g}$ satisfying the following equalities,
$\langle K(X,Y),Z\rangle=\langle K(Y,Z),X\rangle=\langle K(Z,X),Y\rangle.$
where $\langle,\rangle$ denotes the inner product on the vector space $\mathfrak{g}$ corresponding to the left-invariant metric $g$.
For the sake of computations in sections later, we summarize some correspondences in the form of the following proposition, which is a reformulation of Proposition 2.2.3.
Proposition 3.2.1.
Let $g$ be a left-invariant Riemannian metric on $G$, and $\langle,\rangle$ denotes the inner product on $\mathfrak{g}$ corresponding to $g$. The following maps are bijective:
$\mathcal{LA}_{\mathrm{Stat}}(G,g)\ni\nabla$ $\longmapsto K^{(g,\nabla)}\in\mathcal{LK}(G,g)\cong\mathcal{K}(\mathfrak{g},g),$
$\mathcal{LK}(G,g)\cong\mathcal{K}(\mathfrak{g},g)\ni K$ $\longmapsto C^{K}\in S^{3}(\mathfrak{g}^{\ast}).$
Here, $K^{(g,\nabla)}$ is the difference tensor of $(g,\nabla)$, and $C^{K}(X,Y,Z):=-2\langle K(X,Y),Z\rangle$. The composition of the above maps can be expressed as follows:
$\mathcal{LA}_{\mathrm{Stat}}(G,g)\ni\nabla\longmapsto\nabla g\in S^{3}(\mathfrak{g}^{\ast}).$
In particular, the composition defines a bijection
$\mathop{\mathcal{L}\mathrm{Stat}}(G)=\bigsqcup_{g\in\mathop{\widetilde{\mathfrak{M}}}(G)}\mathcal{LA}_{\mathrm{Stat}}(G,g)\to\mathop{\widetilde{\mathfrak{M}}}(G)\times S^{3}(\mathfrak{g}^{\ast}),\quad(g,\nabla)\mapsto(g,C^{(g,K^{(g,\nabla)})}=\nabla g),$
and hence $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ is a trivial vector bundle on the symmetric space $\mathop{\widetilde{\mathfrak{M}}}(G)$ with the fiber $S^{3}(\mathfrak{g}^{\ast})$.
The proof is straightforward.
Throughout this paper, we also use the following terminologies:
$\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)$ $\coloneqq\{(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)\mid(g,\nabla)\text{ is conjugate symmetric}\},$
[…]
Let us fix $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$ and put:
  • $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g):=\{C\in S^{3}(\mathfrak{g}^{\ast})\ |\ (g,C)\text{ is conjugate symmetric}\}$,
  • $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g):=\{C\in S^{3}(\mathfrak{g}^{\ast})\ |\ (g,C)\text{ is dually flat}\}$.
Note that $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$ is always a linear subspace of $S^{3}(\mathfrak{g}^{\ast})$ but $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g)$ is not.
3.3. Some remarks on computations
Let $G$ be an $n$-dimensional Lie group and $g$ a left-invariant Riemannian metric on $G$. In this subsection, we introduce several computational formulas for left-invariant statistical connections on $(G,g)$ that will be used in later sections.
We denote by $\mathfrak{g}$ the Lie algebra of $G$, that is, the Lie algebra of all left-invariant vector fields on $G$, and $\langle,\rangle$ the inner product on $\mathfrak{g}$ corresponding to $g$. Let us fix an orthonormal basis $\{e_{1},\dots,e_{n}\}$ of $\mathfrak{g}$ with respect to $\langle,\rangle$. The structure constant for $\{e_{1},\dots,e_{n}\}$ is denoted by $a_{ij}^{k}:=\langle[e_{i},e_{j}],e_{k}\rangle$.
For the Levi-Civita connection $\nabla^{g}$ on $(G,g)$, the (generalized) Christoffel symbol with respect to the global frame $\{e_{1},\dots,e_{n}\}$ is denoted by $\Gamma_{ij}^{k}:=\langle\nabla^{g}_{e_{i}}e_{j},e_{k}\rangle$. Note that $\nabla^{g}_{e_{i}}e_{j}$ is left-invariant on $G$ and thus $\Gamma_{ij}^{k}$ is constant on $G$.
The following is a formula for computing $\Gamma^{k}_{ij}$ from the structure constants.
Proposition 3.3.1.
$\Gamma_{ij}^{k}=\frac{1}{2}(a_{ij}^{k}+a_{ki}^{j}+a_{kj}^{i})$.
Proof.
It is well-known that for each $X,Y\in\mathfrak{g}$,
$\nabla^{g}_{X}Y=\frac{1}{2}[X,Y]+U(X,Y)$
holds, where $U$ is a $\mathfrak{g}$-valued symmetric bilinear form on $\mathfrak{g}$ defined by
$2\langle U(X,Y),Z\rangle=\langle[Z,X],Y\rangle+\langle X,[Z,Y]\rangle\quad(X,Y,Z\in\mathfrak{g})$
(cf. [15], Chapter X). By the definition,
$\langle U(e_{i},e_{j}),e_{k}\rangle=\frac{1}{2}(a_{ki}^{j}+a_{kj}^{i}),$
and thus
$\Gamma_{ij}^{k}$ $=\langle\nabla^{g}_{e_{i}}e_{j},e_{k}\rangle$
$=\langle\frac{1}{2}[e_{i},e_{j}]+U(e_{i},e_{j}),e_{k}\rangle$
$=\frac{1}{2}\langle[e_{i},e_{j}],e_{k}\rangle+\langle U(e_{i},e_{j}),e_{k}\rangle$
$=\frac{1}{2}(a_{ij}^{k}+a_{ki}^{j}+a_{kj}^{i}).$
We write $\{x_{1},\dots,x_{n}\}$ for the dual basis of the orthonormal basis $\{e_{1},\dots,e_{n}\}$ of $\mathfrak{g}$. Recall that $\{x_{i_{1}}x_{i_{2}}x_{i_{3}}\}_{i_{1}\leq i_{2}\leq i_{3}}$ forms a basis of $S^{3}(\mathfrak{g}^{\ast})$ (see Section 3.2). The following is a formula for computing the tensor $\nabla^{g}(x_{i_{1}}x_{i_{2}}x_{i_{3}})\in S^{3}(\mathfrak{g}^{\ast})\otimes\mathfrak{g}^{\ast}$.
Proposition 3.3.2.
Let $1\leq i_{1},i_{2},i_{3}\leq n$. Then the tensor $\nabla^{g}(x_{i_{1}}x_{i_{2}}x_{i_{3}})\in S^{3}(\mathfrak{g}^{\ast})\otimes\mathfrak{g}^{\ast}$ can be written as
$\nabla^{g}(x_{i_{1}}x_{i_{2}}x_{i_{3}})=-\sum_{t,u}(\Gamma_{tu}^{i_{1}}(x_{u}x_{i_{2}}x_{i_{3}})+\Gamma_{tu}^{i_{2}}(x_{u}x_{i_{1}}x_{i_{3}})+\Gamma_{tu}^{i_{3}}(x_{u}x_{i_{1}}x_{i_{2}}))\otimes x_{t}.$
The proof is straight forward.
For each $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$, we simply write $R^{C}_{g}$ for the left-invariant $(0,4)$ tensor field $R_{g}^{\nabla^{(g,C)}}$ on $G$, that is,
$R^{C}_{g}(X,Y,Z,W)=g(R^{\nabla^{(g,C)}}(X,Y)Z,W),$
where $\nabla^{(g,C)}$ denotes the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.2.1. By Proposition 2.3.2, we shall regard the left-invariant $(0,4)$-tensor field $R^{C}_{g}$ on $G$ as an element of $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$, where $\wedge^{2}\mathfrak{g}^{\ast}$ denotes the space of alternating $2$-tensors on $\mathfrak{g}^{\ast}$, and $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ the space of symmetric $2$-tensors on $(\wedge^{2}\mathfrak{g}^{\ast})$. Let us put $\omega_{ij}:=x_{i}\otimes x_{j}-x_{j}\otimes x_{i}$ for each $i,j$. Then $\{\omega_{ij}\}_{i<j}$ forms a basis of the space $\wedge^{2}\mathfrak{g}^{\ast}$ of alternating $2$-tensors on $\mathfrak{g}^{\ast}$. We shall write $(i,j)\leq(k,l)$ if $i<k$ or “$i=k$ and $j\leq k$”. Then the family $\{\omega_{ij}\odot\omega_{kl}\}_{i<j,k<l,(i,j)\leq(k,l)}$ forms a basis of linear space $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$.
We note that $\nabla^{(g,0)}=\nabla^{g}$ and thus $R_{g}^{0}$ is the $(0,4)$-curvature tensor of the Riemannian manifold $(G,g)$. Such curvature tensor can be written as
$R^{0}_{g}=\sum_{i<j,k<l,(i,j)\leq(k,l)}r_{ijkl}(\omega_{ij}\odot\omega_{kl})$
with
$r_{ijkl}=2R^{0}_{g}(e_{i},e_{j},e_{k},e_{l})=2\sum_{u}(\Gamma_{jk}^{u}\Gamma_{iu}^{l}-\Gamma_{ik}^{u}\Gamma_{ju}^{l}-a_{ij}^{u}\Gamma_{uk}^{l})$
for $i<j$, $k<l$, $(i,j)<(k,l)$ and
$r_{ijij}=R^{0}_{g}(e_{i},e_{j},e_{i},e_{j})=\sum_{u}(\Gamma_{ji}^{u}\Gamma_{iu}^{j}-\Gamma_{ii}^{u}\Gamma_{ju}^{j}-a_{ij}^{u}\Gamma_{ui}^{j})$
for $i<j$.
Throughout this paper, for each $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$, we denote the $(0,4)$-tensor $R_{g}^{C}-R_{g}^{0}\in(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ by $K^{(g,C)},K^{(g,C)}]$, or simply $[K,K]$ when the context is clear. Then by Proposition 2.3.2,
$K,K](X,Y,Z,W)=\langle[K_{X},K_{Y}]Z,W\rangle$
holds for $X,Y,Z,W\in\mathfrak{g}$.
By direct calculation, we obtain the following:
Theorem 3.3.3.
Let $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$. For each $u=1,\dots,n$, we define the matrix $K_{u}\in M(n,\mathbb{R})$ by putting
$(K_{u})_{ij}:=\langle K^{(g,C)}_{e_{u}}(e_{i}),e_{j}\rangle=-\frac{1}{2}C(e_{u},e_{i},e_{j}).$
Then
$[K^{(g,C)},K^{(g,C)}]=2\sum_{i<j,k<l,(i,j)<(k,l)}[K_{l},K_{k}]_{ij}(\omega_{ij}\odot\omega_{kl})+\sum_{i<j}[K_{j},K_{i}]_{ij}(\omega_{ij}\odot\omega_{ij}),$
where $K_{u},K_{v}]:=K_{u}K_{v}-K_{v}K_{u}$ denotes the Lie bracket of the matrices $K_{u}$ and $K_{v}$.
4. Moduli spaces
Throughout this section, let $G$ be an $n$-dimensional Lie group with finitely many connected components, and $\mathfrak{g}$ the Lie algebra of $G$. The automorphism group of $G$ and that of $\mathfrak{g}$ are denoted by $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$, respectively. Note that $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$ are both Lie groups. The moduli space of all left-invariant statistical structures on $G$ is defined in this section.
4.1. The moduli space of left-invariant Riemannian metrics
In this subsection, we recall the definition of moduli spaces of left-invariant Riemannian metrics on Lie groups (cf. [18]).
Throughout this paper, we say that two Riemannian manifolds $(M,g)$ and $(M^{\prime},g^{\prime})$ are isometric up to scaling if there exist a scalar $r>0$ and a diffeomorphism $f:M\to M^{\prime}$ such that $r\cdot f^{*}g^{\prime}=g$.
As is well-known, the Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ acts smoothly on the space of left-invariant Riemannian metrics $\mathop{\widetilde{\mathfrak{M}}}(G)$ (see Section  3.2 for more details on $\mathop{\widetilde{\mathfrak{M}}}(G)$) by
(4.1) $(r,\varphi).g\coloneqq r\cdot(\varphi^{-1})^{\ast}g=r\cdot g(\varphi^{-1}_{\ast}\ \cdot\ ,\varphi^{-1}_{\ast}\ \cdot\ )\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),\ g\in\mathop{\widetilde{\mathfrak{M}}}(G)).$
If left-invariant Riemannian metrics $g_{1}$ and $g_{2}$ are related by the group action in (4.1), then the Riemannian manifolds $(G,g_{1})$ and $(G,g_{2})$ are isometric up to scaling, see Appendix B for details. We call the quotient space $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))\backslash\mathop{\widetilde{\mathfrak{M}}}(G)$ the moduli space of left-invariant Riemannian metrics on $G$, and denote by $\mathfrak{PM}(G)$. In what follows, we regard $\mathfrak{PM}(G)$ as a topological space equipped with the natural quotient topology.
The Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g})$ also acts smoothly on $\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ as below:
(4.2) $(r,\varphi).\langle,\rangle\coloneqq r\cdot\langle\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ \rangle\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g}),\ \langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})).$
Note that if $G$ is connected and simply-connected, then $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$ are isomorphic to each other as Lie groups, and furthermore, the homeomorphic correspondence $\mathop{\widetilde{\mathfrak{M}}}(G)\to\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ is equivariant regarding the actions of $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$. Thus the moduli space $\mathfrak{PM}(G)$ is homeomorphic to the quotient space $(\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g}))\backslash\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ in such a situation.
Remark 4.1.1.
In [18], for a Lie group $G$, the orbit space of $\mathop{\widetilde{\mathfrak{M}}}(G)$ under a certain action of the group $\mathbb{R}^{\times}\times\mathrm{Aut}(G)$ is referred to as the moduli space of left-invariant Riemannian metrics on $G$. It is straightforward to see that the action of $\mathbb{R}^{\times}\times\mathrm{Aut}(G)$ on $\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ factors through the group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ via the natural surjection
$\mathbb{R}^{\times}\times\mathrm{Aut}(G)\to\mathbb{R}_{>0}\times\mathrm{Aut}(G),\quad(r,\varphi)\mapsto(r^{2},\varphi).$
Therefore, the moduli space $\mathfrak{PM}(G)$ obtained via these two quotient constructions coincide.
The connected and simply-connected Lie group with $\mathfrak{PM}(G)$ being a singleton have been classified by [23].
Proposition 4.1.2 ([23]; see also [18]).
Let $G$ be a connected and simply-connected Lie group which $\mathfrak{PM}(G)$ be a singleton. Then, $G$ is isomorphic to one of the following Lie groups:
(4.3) $\mathbb{R}^{n},~G_{\mathbb{R}\mathrm{H}^{n}}\ (n\geq 2)~\text{ or }~H^{3}\times\mathbb{R}^{n-3}\ (n\geq 3).$
The Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ is the solvable part of the Iwasawa decomposition of $SO^{0}(n,1)$. Here, $SO^{0}(n,1)$ is the identity component of $SO(n,1)$. Particularly, $G_{\mathbb{R}\mathrm{H}^{n}}$ acts simply-transitively on $n$-dimensional real hyperbolic space $\mathbb{R}\mathrm{H}^{n}\cong\mathbb{R}_{>0}\times\mathbb{R}^{n-1}$. The Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ is known as the Lie group of $n$-dimensional real hyperbolic space $\mathbb{R}\mathrm{H}^{n}$. The Lie group $H^{3}$ is the $3$-dimensional Heisenberg group.
4.2. The moduli space of left-invariant statistical structures
In this subsection, we define the moduli space of left-invariant statistical structures on a Lie group $G$.
The Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ acts smoothly on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ (see Section 3.2 for the notation of $\mathop{\mathcal{L}\mathrm{Stat}}(G)$) by
(4.4) $(r,\varphi).(g,\nabla)\coloneqq(r\cdot(\varphi^{-1})^{\ast}g,(\varphi^{-1})^{\ast}\nabla)\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),\ (g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)).$
We mentioned the smooth structure on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ in Section 3.2 as the trivial vector bundle $\mathop{\mathcal{L}\mathrm{Stat}}(G)\cong\mathop{\widetilde{\mathfrak{M}}}(G)\times S^{3}(\mathfrak{g}^{\ast})$. In this paper, the equivalence relation on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ defined by the group action (4.4) is denoted by $\approx$, that is $(g,\nabla)\approx(g^{\prime},\nabla^{\prime})$ means that there exists a scalar $r>0$ and $\varphi\in\mathrm{Aut}(G)$ which satisfies $(r,\varphi).(g,\nabla)=(g^{\prime},\nabla^{\prime})$.
We shall introduce the following notion:
Definition 4.2.1.
Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both left-invariant statistical structures on $G$. We say that a scaling statistical isomorphism $(f,r)$ from $(G,g,\nabla)$ to $(G,g^{\prime},\nabla^{\prime})$ (see Definition 2.2.1 for the notation) is compatible with the group structure if there exists $\varphi\in\mathrm{Aut}(G)$ such that $f(gx)=\varphi(g)f(x)$ for all $g,x\in G$.
Then the next proposition holds:
Proposition 4.2.2.
Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both left-invariant statistical structures on $G$. Then $(g,\nabla)\approx(g^{\prime},\nabla^{\prime})$ on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ holds if and only if there exists a scaling statistical isomorphism compatible with the group structure between $(G,g,\nabla)$ and $(G,g^{\prime},\nabla^{\prime})$.
The proof of Proposition 4.2.2 is postponed to Appendix B (as Theorem B.0.3).
As a corollary to Proposition 4.2.2, one sees that the group action (4.4) preserves properties such as conjugate symmetry, constant curvature, dually flatness, and CHC for left-invariant statistical structures on $G$ (see Section 2.3 for the definitions of these classes).
Definition 4.2.3.
[…]
It should be remarked that the moduli spaces, equipped with the quotient topology, might not be Hausdorff in general. In fact, the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ is not Hausdorff if $G=\mathbb{R}^{n}$ (see Remark 5.1.6).
We also define the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ of left-invariant conjugate symmetric statistical structure on $G$ and the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$ of left-invariant dually flat structure on $G$ by the quotient space as follows:
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ $:=(\mathbb{R}_{>0}\times\mathrm{Aut}(G))\backslash\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)\text{ and }$
[…]
Note that the relative topology of $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ induced from $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ coincides the quotient topology defined by the surjective map $\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)\to\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$. The topology of $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$ is similar.
4.3. For the cases where $\mathfrak{PM}(G)\cong\{\ast\}$
In this subsection, let us assume that the moduli space $\mathfrak{PM}(G)$ of left-invariant Riemannian metrics on $G$ is a singleton. Then $\mathop{\widetilde{\mathfrak{M}}}(G)$ is a homogeneous space of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$, and $\mathop{\mathcal{L}\mathrm{Stat}}(G)\cong\mathop{\widetilde{\mathfrak{M}}}(G)\times S^{3}(\mathfrak{g}^{\ast})$ is an $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))$-equivariant trivial vector bundle on the homogeneous space $\mathop{\widetilde{\mathfrak{M}}}(G)$. In such a situation, $\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)$ [resp. $\mathcal{L}\mathrm{Stat}^{\mathrm{DF}}(G)$] is an $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))$-equivariant sub-vector-bundle [resp. sub-fiber-bundle] of the trivial bundle $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ on the homogeneous space $\mathop{\widetilde{\mathfrak{M}}}(G)$ with its fiber $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$ [resp. $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g)$] at each point $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$ (see Section 3.2 for the definitions of $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$ and $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g)$).
By the general theory of the topology of equivariant fiber bundles on homogeneous spaces (cf. Theorem A.0.1 in Appendix A), we obtain the following:
Proposition 4.3.1.
Let $G$ be a Lie group with $\mathfrak{PM}(G)\cong\{\ast\}$ (see Section 4.1 for the details). Fix $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$. Then the following holds:
  1. (1)
    The map
    $F_{g}:S^{3}(\mathfrak{g}^{\ast})\to\mathop{\mathcal{ML}\mathrm{Stat}}(G),~C\mapsto(\mathbb{R}_{>0}\times\mathrm{Aut}(G)).(g,\nabla^{(g,C)})$
    is continuous, open and surjective. Here, $\nabla^{(g,C)}$ is the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.2.1.
  2. (2)
    The moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ is homeomorphic to the quotient space
    (4.5) $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}\backslash S^{3}(\mathfrak{g}^{\ast}).$
Here, $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$ denotes the isotropy subgroup of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ with respect to the action (4.1) at the point $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$, and the Lie group $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$ acts on $S^{3}(\mathfrak{g}^{\ast})$ as below:
(4.6) $(r,\varphi).C\coloneqq r\cdot C(\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ )\quad((r,\varphi)\in(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g},\ C\in S^{3}(\mathfrak{g}^{\ast})),$
where $\varphi^{-1}\cdot$ denotes the induced action of $\varphi^{-1}\in\mathrm{Aut}(G)$ on $\mathfrak{g}$. Similarly, the moduli spaces $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$ are homeomorphic to the quotient spaces
$(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}\backslash S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g),$
and
$(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}\backslash S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g),$
respectively.
We also note that for the isotropy subgroup $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$, the following is known:
Proposition 4.3.2 ([37], Lemma 4.2).
Assume that $G$ is connected and simply-connected (and thus $\mathrm{Aut}(G)$ can be identified with $\mathrm{Aut}(\mathfrak{g})$). Fix $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$. We denote by $\langle,\rangle$ the corresponding inner product on $\mathfrak{g}$. Then the claims below hold:
  1. (1)
    If the Lie algebra $\mathfrak{g}$ is abelian, the isotropy subgroup $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$ is isomorphic to the Lie group $\mathbb{R}_{>0}\times O(\mathfrak{g},\langle,\rangle)$.
  2. (2)
    If the Lie algebra $\mathfrak{g}$ is non-abelian, the isotropy subgroup $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$ coincides with $\{1\}\times(\mathrm{Aut}(\mathfrak{g})\cap O(\mathfrak{g},\langle,\rangle))\subset\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g})$.
For each Lie group appearing in (4.3), the detail of the isotropy subgroup $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$ is described in Sections 5.2, 6.2 and 7.2, respectively.
5. The case of $\mathbb{R}^{n}$
In this section, we consider the commutative Lie group $\mathbb{R}^{n}$ with $n\geq 1$. Then $\mathfrak{PM}(\mathbb{R}^{n})\cong\{\ast\}$ (see Proposition 4.1.2).
Throughout this section, we denote the Lie algebra of $\mathbb{R}^{n}$ by $\mathfrak{g}_{\mathbb{R}^{n}}$. Let us fix the standard inner product $\langle,\rangle$ on $\mathfrak{g}_{\mathbb{R}^{n}}\cong\mathbb{R}^{n}$, and we denote by $g^{E}$ the left-invariant Riemannian metric on $\mathbb{R}^{n}$ corresponding to $\langle,\rangle$, that is, $g^{E}$ denotes the usual Euclidean metric on $\mathbb{R}^{n}$.
5.1. Main theorem for $\mathbb{R}^{n}$
The goal of this section is to give a proof of the following theorem:
Theorem 5.1.1.
The following hold.
  1. (1)
    All left-invariant statistical structures on $\mathbb{R}^{n}$ are conjugate symmetric, that is $\mathop{\mathcal{L}\mathrm{Stat}}(\mathbb{R}^{n})=\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(\mathbb{R}^{n})$ holds.
  2. (2)
    Let $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Then the following two conditions on $(\langle,\rangle,C)$ are equivalent:
    1. (i)
      The left-invariant statistical structure $(\langle,\rangle,C)$ on $\mathbb{R}^{n}$ is dually flat.
    2. (ii)
      There exists an orthonormal basis $\{e_{1},\dots,e_{n}\}$ on $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ such that $C$ is expressed as
      $C=\sum_{i=1}^{n}\lambda_{i}x_{i}^{3}\quad(\lambda_{i}\in\mathbb{R}).$
      Here, $\{x_{1},\dots,x_{n}\}\subset\mathfrak{g}_{\mathbb{R}^{n}}^{\ast}$ is the dual basis of $\{e_{1},\dots,e_{n}\}$.
  3. (3)
[…]
  4. (4)
    Let $\mathcal{B}=\{e_{1},\dots,e_{n}\}$ be an orthonormal basis of $(\mathfrak{g}_{\mathbb{R}},\langle,\rangle)$, and we denote by $\{x_{1},\dots,x_{n}\}$ the dual basis of $\mathcal{B}$. Put
    $V_{\mathcal{B}}:=\left\{\sum_{i=1}^{n}\lambda_{i}x_{i}^{3}\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})\biggm|\lambda_{i}\in\mathbb{R}\right\}\subset S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast}).$
    Then
[…]
As a corollary to Theorem 5.1.1 (2), we have the following proposition:
Proposition 5.1.2.
Let $g$ be a left-invariant Riemannian metric on $\mathbb{R}^{n}$ with $n\geq 2$. Then a left-invariant dually flat structure $(g,\nabla)$ on $\mathbb{R}^{n}$ is of CHC (see Section 2.3 for the notation) if and only if $\nabla$ is the Levi-Civita connection of $g$.
Proof.
The Hessian structure $(g,\nabla^{g})$ is known to be of dually flat of CHC 0 on $\mathbb{R}^{n}$ (see [34]). Conversely, we suppose that $(g,\nabla)$ is a left-invariant dually flat structure on $\mathbb{R}^{n}$ of CHC, and let us prove that $\nabla=\nabla^{g}$. By Proposition 2.3.2 and $R^{\nabla}=R^{\nabla^{g}}=0$, we have $K_{X},K_{Y}]=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}$. Here, $K$ is the difference tensor of $(g,\nabla)$. We also note that $(\mathbb{R}^{n},g)$ has the constant sectional curvature 0. Thus by [34, Corollary 2] and [10, Proposition 3.1], for each $X,Y,Z\in\mathfrak{g}_{\mathbb{R}^{n}}$, the equality below holds:
(5.1) $K(K(Z,X),Y)$ $=0\text{ for }X,Y,Z\in\mathfrak{g}_{\mathbb{R}^{n}}.$
Take the orthonormal basis $e_{1},\dots,e_{n}$ of $\mathfrak{g}_{\mathbb{R}^{n}}$ as in Theorem 5.1.1 (2) and write $C=\sum_{i}\lambda_{i}x_{i}^{3}$. Note that $K(e_{i},e_{i})=\lambda_{i}e_{i}$ for each $i$. Therefore, we have
$0=K(K(Z,X),Y)=\lambda_{i}^{2}e_{i}.$
This proves that $C=0$ and hence $\nabla=\nabla^{g}$. ∎
Let us denote by $g^{E_{1}}$ the Euclidean metric on the Euclidean line $\mathbb{R}$. We fix a unit vector $e\in\mathfrak{g}_{\mathbb{R}}$ with respect to the inner-product on $\mathfrak{g}_{\mathbb{R}}$ induced by $g^{E_{1}}$, and $t\in\mathfrak{g}_{\mathbb{R}}^{\ast}$ is defined by $t(e)=1$. Note that for each $a\in\mathbb{R}$, $at^{3}$ defines a left-invariant symmetric $(0,3)$-tensor fields on $\mathbb{R}$, and $(\mathbb{R},g^{E_{1}},at^{3})$ defines a one-dimensional dually flat statistical manifold.
Theorem 5.1.3.
For each $(g,C)\in\mathcal{L}\mathrm{Stat}^{\mathrm{DF}}(\mathbb{R}^{n})$, there exists $(a_{1},\dots,a_{n})\in\mathbb{R}^{n}$ such that $(\mathbb{R}^{n},g,C)$ is statistically isomorphic to $\prod_{i=1}^{n}(\mathbb{R},g^{E_{1}},a_{i}t^{3})$, where $\prod_{i=1}^{n}(\mathbb{R},g^{E_{1}},a_{i}t^{3})$ denotes the direct product of the family $\{(\mathbb{R},g^{E_{1}},a_{i}t^{3})\}_{i}$ of one-dimensional dually flat statistical manifolds in the sense of Section 2.2.
For each $p,r\in\mathbb{Z}_{\geq 0}$ with $p+r=n$, and each $c=(c_{1},\dots,c_{n})\in\mathbb{R}^{n}$ with $c_{1}\geq c_{2}\geq\dots c_{p}>0$ and $c_{p+1}=\dots=c_{p+r}=0$, let us define the Hessian manifold (i.e. a dually flat statistical manifold) $(\Omega_{c},g_{c},\nabla_{c})$ as below. The open domain $\Omega_{c}$ of the Euclidean space $\mathbb{R}^{n}$ is defined by
$\Omega_{c}:=\{(x_{1},\dots,x_{n})\mid x_{i}>0\text{ for each }i\leq p\},$
and $\nabla_{c}$ denotes the Euclidean connection on $\Omega_{c}$. The Riemannian metric $g_{c}$ on $\Omega_{c}$ is defined as the pullback of the Euclidean metric $g^{E}$ on $\mathbb{R}^{n}$ by the diffeomorphism
$\phi_{c}:\Omega_{c}\rightarrow\mathbb{R}^{n},~(x_{1},\dots,x_{n})\mapsto(\psi_{c_{1}}(x_{1}),\dots,\psi_{c_{n}}(x_{n})),$
where we put
$\psi_{c_{i}}(x_{i}):=\begin{cases}\frac{2}{c_{i}}\log(\frac{c_{i}}{2}x_{i})\text{ if }i\leq p\\ x_{i}\text{ if }p<i.\end{cases}$
Corollary 5.1.4.
For each $(g,\nabla)\in\mathcal{L}\mathrm{Stat}^{\mathrm{DF}}(\mathbb{R}^{n})$, there exists $c\in\{\lambda\in\mathbb{R}^{n}\mid\lambda_{1}\geq\dots\lambda_{n}\geq 0\}$ such that the Hessian manifold $(\mathbb{R}^{n},g,\nabla)$ is statistically isomorphic to the Hessian manifold $(\Omega_{c},g_{c},\nabla_{c})$ up to scaling.
Proof.
By Proposition 4.1.2, and Theorem 5.1.1, there exists an orthogonal basis $\mathcal{B}=\{e_{1},\dots,e_{n}\}$ on $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c=(c_{1},\dots,c_{n})\in\{\lambda\in\mathbb{R}^{n}\mid\lambda_{1}\geq\dots\lambda_{n}\geq 0\}$ such that by putting $C=\sum_{i}c_{i}x_{i}^{3}$, the dually flat statistical manifold $(\mathbb{R}^{n},g,\nabla)$ is statistically isomorphic to $(\mathbb{R}^{n},g^{E},C)$ up to scaling. Then, it can be directly verified that the diffeomorphism $\phi_{c}$ gives a scaling statistical isomorphism from the dually flat statistical manifold $(\Omega_{c},g_{c},\nabla_{c})$ onto $(\mathbb{R}^{n},g^{E},C)$. ∎
Remark 5.1.5.
It should be noted that, in the context of Hessian geometry, the claim of Corollary 5.1.4 can essentially be found in [33, 35], where it appears as the classification of homogeneous Hessian domains associated with commutative normal Hessian algebras.
Remark 5.1.6.
The moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(\mathbb{R}^{n})$ is not Hausdorff. In fact, the orbit of the Levi-Civita connections of left-invariant Riemannian metrics and other any orbit can not be separated in the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(\mathbb{R}^{n})$ because of the scaling actions of $\mathbb{R}_{>0}$. Note that, if we remove the point represented by Levi-Civita connections in the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(\mathbb{R}^{n})$, such subspace is Hausdorff.
5.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $\mathbb{R}^{n}$
In this subsection, we give a proof of Theorem 5.1.1.
Proof of (1) in Theorem 5.1.1.
By Proposition 4.1.2, our goal is to show that $S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)=S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Fix any $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Since $\nabla^{g^{E}}_{Y}X=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}^{\ast}$, we have $\nabla^{g^{E}}C=0$. This proves that $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)$ by Proposition 2.3.1. ∎
Proof of (2) in Theorem 5.1.1.
Let $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$. Throughout this proof, the difference tensor of the left-invariant statistical structure $(\langle,\rangle,C)$ is denoted by $K$, and the affine connection by $\nabla$ in the sense of Proposition 3.2.1. Moreover, we denote the $(1,3)$-curvature tensor of $\nabla$ by $R$, and the $(1,3)$-curvature tensor of Levi-Civita connection of $g^{E}$ by $R_{g^{E}}\equiv 0$.
Proof of (ii) $\Rightarrow$ (i):
Assume (ii). Then by Proposition 3.2.1, the difference tensor $K$ satisfies
(5.2) $K(e_{i},e_{j})=-2\delta_{ij}\lambda_{i}e_{i}\quad(i,j=1,\dots,n),$
where $\delta_{ij}$ denotes the Kronecker delta. We aim to show that $R$ coincides with 0. By combining $R_{g^{E}}\equiv 0$, Theorem 5.1.1 (1) with the equality Proposition 2.3.2, we see the following.
$R(e_{i},e_{j})=[K_{e_{i}},K_{e_{j}}]\quad(i,j=1,\dots,n).$
By Equation (5.2) and Proposition 2.2.4, we see that
$R(e_{i},e_{j})=0\quad(i,j=1,\dots,n).$
Thus, $R\equiv 0$.
Proof of (i) $\Rightarrow$ (ii):
Assume (i). Recall that any dually flat structure is conjugate symmetric. Thus, by Proposition 2.3.2 and $R_{g^{E}}\equiv 0$, we have the equality
$R(X,Y)=[K_{X},K_{Y}]$
for all vector fields $X$ and $Y$. By the assumption (i),
$K_{u},K_{v}]\equiv 0$
holds for all $u$, $v\in\mathfrak{g}_{\mathbb{R}^{n}}$. By Proposition 2.2.4, there exists an orthonormal basis $e_{1},\dots,e_{n}$ of $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c_{1},\dots c_{n}\in\mathbb{R}$ such that
$K(e_{i},e_{i})=c_{i}e_{i},\quad K(e_{i},e_{j})=0$
for distinct $i,j=1,\dots,n$. Since $C(u,v,w)=-2\langle K(u,v),w\rangle$ for any $u$, $v$ and $w\in\mathfrak{g}_{\mathbb{R}^{n}}$, the following holds:
$C(e_{i},e_{j},e_{k})=\begin{cases}-2c_{i}&(i=j=k)\\ 0&(\text{otherwise}).\end{cases}$
Therefore, by setting $\lambda_{i}=-2c_{i}$, we see that $C$ is expressed as follows:
$C=\sum_{i=1}^{n}\lambda_{i}x_{i}^{3}.$
This completes the proof.
Proofs of (3) and (4) in Theorem 5.1.1.
The claim
[…]
follows from Proposition 4.3.1, Proposition 4.3.2 (1) and Theorem 5.1.1 (1). The claim
[…]
is obtained by applying Proposition 4.3.1, Proposition 4.3.2 (1), Theorem 5.1.1 (2), and Theorem C.0.1 in Appendix C. ∎
6. The case of $G_{\mathbb{R}\mathrm{H}^{n}}$
In this section, we consider the connected and simply-connected Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ with $n\geq 2$ (see Section 4.1 for the definition of $G_{\mathbb{R}\mathrm{H}^{n}}$). Then $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$ (see Proposition 4.1.2).
Throughout this section, we denote the Lie algebra of $G_{\mathbb{R}\mathrm{H}^{n}}$ by $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$. Let us fix a canonical basis $\{e_{1},\dots,e_{n}\}$ on $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ (see [21]), that is, $\{e_{1},\dots,e_{n}\}$ is a basis on $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ with
(6.1) $e_{1},e_{i}]=e_{i}\quad(i=2,\dots,n),$
and the other bracket products are trivial. The dual basis of $\{e_{1},\dots,e_{n}\}$ is denoted by $\{x_{1},\dots,x_{n}\}$. Let us consider the inner product $\langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})$ such that the canonical basis is orthonormal. We denote by $g$ the left-invariant Riemannian metric on $G_{\mathbb{R}\mathrm{H}^{n}}$ corresponding to $\langle,\rangle$. It is well-known (see [21]) that the $(0,4)$-curvature tensor $R^{0}_{g}$ of $g$ can be written as
$R^{0}_{g}=\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$
where we put $\omega_{ij}:=x_{i}\otimes x_{j}-x_{j}\otimes x_{i}$. The Riemannian manifold $(G_{\mathbb{R}H^{n}},g)$ has constant sectional curvature.
6.1. Main theorem for $G_{\mathbb{R}\mathrm{H}^{n}}$
The goal of this section is to give a proof of the following theorem:
Theorem 6.1.1.
Let us define
$C^{\alpha}:=\alpha\bigg(4x_{1}^{3}+6\sum_{i=2}^{n}x_{1}x_{i}^{2}\bigg)\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})\cong\mathrm{Homog}_{3}(x_{1},\dots,x_{n})$
for each $\alpha\in\mathbb{R}$. Then
$\nabla^{g}C^{\alpha}$ $=-6\alpha(\sum_{i\geq 2}x_{i}^{2})^{2}\in S^{4}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast}),$
$R^{C^{\alpha}}_{g}$ $=(1-\alpha^{2})R^{0}_{g}=(1-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij},$
$S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast},g)$ $=\{C^{\alpha}\mid\alpha\in\mathbb{R}\},~\text{ and }$
$S^{3}_{\mathrm{DF}}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast},g)$ $=\{C^{1},C^{-1}\}.$
Furthermore,
$\mathop{\mathcal{ML}\mathrm{Stat}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong(SO(1)\times O(n-1))\backslash\mathrm{Homog}_{3}(x_{1},\dots,x_{n}),$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong\{C^{\alpha}\mid\alpha\in\mathbb{R}\}\cong\mathbb{R},~\text{ and }$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong\{C^{\alpha}\mid\alpha\in\pm 1\}\cong\{\pm 1\}.$
Remark 6.1.2.
The dually flat structure $(g,C^{1})$ can be found in [8, Example 2.3], as an example of Hessian structures of CHC $4$ on hyperbolic spaces. Notice that its dual structure $(\langle,\rangle,C^{-1})$ on $G_{\mathbb{R}\mathrm{H}^{n}}$ is dually flat but not CHC.
As supplementary remarks to Theorem 6.1.1, we will also discuss the following points in Section 6.3 (see Theorem 6.3.1):
Proposition 6.1.3.
For any $\alpha\in\mathbb{R}$, the left-invariant statistical structure $(g,C^{\alpha})$ has constant curvature. Furthermore, any left-invariant conjugate symmetric statistical structure on $G_{\mathbb{R}\mathrm{H}^{n}}$ is a statistical structure of constant sectional curvature in the sense of Section 2.3.
6.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $G_{\mathbb{R}\mathrm{H}^{n}}$
In this subsection, we give a proof of Theorem 6.1.1.
We shall define the subgroup $L$ of $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})$ by
$L:=\mathrm{Aut}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cap O(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}},\langle,\rangle)\subset GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}).$
Since $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$, by applying Proposition 4.3.1 and 4.3.2 (2), we have
$\mathop{\mathcal{ML}\mathrm{Stat}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong L\backslash S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast}),$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong L\backslash S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast},\langle,\rangle),$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong L\backslash S^{3}_{\mathrm{DF}}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast},\langle,\rangle).$
Therefore Theorem 6.1.1 follows from the three propositions below.
Proposition 6.2.1.
We shall identify $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})$ with $GL(n,\mathbb{R})$ by considering matrices of linear transformations on $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ with respect to the canonical basis. Then the subgroup $L:=\mathrm{Aut}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cap O(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}},\langle,\rangle)$ in $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cong GL(n,\mathbb{R})$ can be written as
$L=\left\{\begin{pmatrix}1&0\\ 0&h_{0}\end{pmatrix}\ \middle|~\ h_{0}\in O(n-1)\right\}\cong SO(1)\times O(n-1).$
Furthermore, $C^{\alpha}\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})$ is $L$-invariant for each $\alpha\in\mathbb{R}$.
Proposition 6.2.2.
Let $C\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})\cong\mathrm{Homog}_{3}(x_{1},\dots,x_{n})$. Then $\nabla^{g}C$ is symmetric if and only if $C=C^{\alpha}$ for some $\alpha\in\mathbb{R}$. Furthermore, $\nabla^{g}C^{\alpha}=-6\alpha(\sum_{i\geq 2}x_{i}^{2})^{2}\in S^{4}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})$.
Proposition 6.2.3.
$R^{C^{\alpha}}_{g}=(1-\alpha^{2})R^{0}_{g}$ for each $\alpha\in\mathbb{R}$.
First, we shall give a proof of Proposition 6.2.1:
Proof of Proposition 6.2.1.
Under the identification $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cong GL(n,\mathbb{R})$, it is well-known that the automorphism group $\mathrm{Aut}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})$ can be written as
$\mathrm{Aut}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})=\left\{\begin{pmatrix}1&0\\ \ast&h_{0}\end{pmatrix}\ \middle|~\ h_{0}\in GL(n-1,\mathbb{R})\right\}.$
Since the canonical basis is orthonormal with respect to the inner product $\langle,\rangle$, under the identification $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cong GL(n,\mathbb{R})$ fixed as above, we have $O(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}},\langle,\rangle)\cong O(n)$. Hence
$L=\mathrm{Aut}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cap O(n)=\left\{\begin{pmatrix}1&0\\ 0&h_{0}\end{pmatrix}\ \middle|~\ h_{0}\in O(n-1)\right\}\cong SO(1)\times O(n-1).$
It is readily seen that the polynomial function $\sum_{i=2}^{n}x_{i}^{2}$ is invariant under the action of the group $L=SO(1)\times O(n-1)$. Therefore, the polynomial function
$C^{\alpha}=\alpha\left(4x_{1}^{3}+6x_{1}\sum_{i=2}^{n}x_{i}^{2}\right)$
is also $L$-invariant. ∎
In order to prove Proposition 6.2.2, we prepare the following lemma:
Lemma 6.2.4.
Let us put $x_{ijkl}:=x_{i}\otimes x_{j}\otimes x_{k}\otimes x_{l}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 4}$ for each $1\leq i,j,k,l\leq n$. Then
$\nabla^{g}(x_{1}^{3})$ $=-\sum_{t\geq 2}(x_{11tt}+x_{1t1t}+x_{t11t}),$
$\nabla^{g}(x_{1}^{2}x_{i})$ $=x_{111i}-\frac{1}{3}\sum_{t\geq 2}(x_{t1it}+x_{ti1t}+x_{1tit}+x_{1itt}+x_{it1t}+x_{i1tt})\quad(\text{for }i\geq 2),$
$\nabla^{g}(x_{1}x_{i}x_{j})$ $=\frac{1}{3}(x_{11ji}+x_{1j1i}+x_{j11i}+x_{11ij}+x_{1i1j}+x_{i11j})$
$\quad\quad-\frac{1}{6}\sum_{t\geq 2}(x_{tijt}+x_{tjit}+x_{itjt}+x_{ijtt}+x_{jtit}+x_{jitt})\quad(\text{for }i,j\geq 2),\text{ and }$
$\nabla^{g}(x_{i}x_{j}x_{k})$ $=(x_{1}\odot x_{i}\odot x_{j}\odot x_{k})-\frac{1}{6}(x_{ijk1}+x_{ikj1}+x_{jik1}+x_{jki1}+x_{kij1}+x_{kji1})\quad(\text{for }i,j,k\geq 2)$
hold.
Proof of Lemma 6.2.4.
By definition, the structure constants of $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ with respect to the canonical basis are given by $a_{1t}^{t}=1$, $a_{t1}^{t}=-1$ ($t\geq 2$), with all other $a_{ij}^{k}$ equal to zero. Then, by applying Proposition 3.3.1, the generalized Christoffel symbols can be computed as
$\Gamma_{t1}^{t}$ $=-1\quad(t\geq 2),$
$\Gamma_{tt}^{1}$ $=1\quad(t\geq 2),$
with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 6.2.4 can be obtained by applying Proposition 3.3.2. ∎
Let us give a proof of Proposition 6.2.2.
Proof of Proposition 6.2.2.
First, let us assume that $\nabla^{g}C$ is symmetric. Put
$\nabla^{g}C=\sum_{u,v,s,t}((\nabla^{g}C)_{uvst})(x_{uvst})$
and $C=\sum_{u,v,s}C_{uvs}(x_{uvs})$, where we put $x_{uvs}:=x_{u}\otimes x_{v}\otimes x_{s}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 3}$ and $x_{uvst}:=x_{u}\otimes x_{v}\otimes x_{s}\otimes x_{t}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 4}$. Then by Lemma 6.2.4,
$0$ $=(\nabla^{g}C)_{i11i}-(\nabla^{g}C)_{i1i1}=-C_{111}+\frac{2}{3}C_{1ii}\quad(\text{for }i\geq 2),$
$0$ $=(\nabla^{g}C)_{111i}-(\nabla^{g}C)_{11i1}=C_{11i}\quad(\text{for }i\geq 2),$
$0$ $=(\nabla^{g}C)_{i11j}-(\nabla^{g}C)_{i1j1}=\frac{1}{3}C_{1ij}\quad(\text{for }2\leq i<j),\text{ and }$
$0$ $=(\nabla^{g}C)_{ijk1}-(\nabla^{g}C)_{ij1k}=-\frac{1}{6}C_{ijk}\quad(\text{for }2\leq i<j<k).$
Thus $C=C_{111}(x_{111}+\frac{3}{2}\sum_{i\geq 2}x_{1}x_{i}^{2})=C^{\alpha}$ for $\alpha=C_{111}/4$. Conversely, by Lemma 6.2.4, one can see that
$\nabla^{g}C^{\alpha}$ $=-2\alpha\sum_{i\geq 2}\sum_{t\geq 2}(x_{tiit}+x_{itit}+x_{iitt})$
$=-2\alpha\bigg(\sum_{i\geq 2}3x_{iiii}+\sum_{2\leq i<t}(x_{iitt}+x_{itit}+x_{itti}+x_{tiit}+x_{titi}+x_{ttii})\bigg)$
$=-6\alpha\bigg(\sum_{i\geq 2}x_{i}^{4}+\sum_{2\leq i<t}2x_{i}^{2}x_{t}^{2}\bigg)$
$=-6\alpha\bigg(\sum_{i\geq 2}x_{i}^{2}\bigg)^{2}$
and in particular, $\nabla^{g}C^{\alpha}$ is symmetric. ∎
Finally, we shall prove Proposition 6.2.3:
Proof of Proposition 6.2.3.
Fix $\alpha\in\mathbb{R}$, and put $C:=C^{\alpha}$. Recall that $R^{0}_{g}=\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$. Thus by applying Proposition 2.3.2, our goal is to show that
$K,K]=[K^{(g,C)},K^{(g,C)}]=(-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$
(see Section 3.3 for the definition of $[K,K]$ and $K^{(g,C)},K^{(g,C)}]$). Put
$\sum_{i,j,l}C_{ijl}x_{ijl}=C=\alpha\bigg(4x_{111}+2\sum_{i\geq 2}(x_{1ii}+x_{i1i}+x_{ii1})\bigg).$
Then for each $l=1,\dots,n$, the matrix $K_{l}=-\frac{1}{2}(C_{lij})_{ij}$ can be written as
$K_{1}=-4\alpha\cdot\mathrm{diag}(2,1,\dots,1),\text{ and }K_{u}=-4\alpha(E_{1u}+E_{u1})\quad\text{ for }u\geq 2,$
where we write $E_{ij}$ for the $n\times n$ matrix unit whose $(i,j)$-entry is $1$ and all other entries are zero. In particular, for $u\geq 2$, $v\geq 1$,
$[K_{u},K_{v}]$ $=\begin{cases}16\alpha^{2}(E_{u1}-E_{1u})\quad\text{ if }v=1,\\ 16\alpha^{2}(E_{uv}-E_{vu})\quad\text{ if }v\geq 2,\end{cases}$
$=16\alpha^{2}(E_{uv}-E_{vu}).$
Therefore by Theorem 3.3.3,
$[K,K]$ $=(-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}=(-\alpha^{2})R^{0}_{g}.$
This completes the proof. ∎
6.3. Amari–Chentsov $\alpha$-connection on the Takano Gaussian space
Throughout this subsection, let $m\in\mathbb{Z}_{\geq 1}$. In this subsection, we present a characterization of the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ on $(\mathcal{N}_{T}^{m},\mathbb{R}^{m},\mathbf{p}_{T})$ (see Sections 2.4 and 3.1 for notation used in this subsection).
Theorem 6.3.1.
Let $\nabla$ be a statistical connection on the Takano Gaussian space $(\mathcal{N}_{T}^{m},g^{F})$. Then, the following three conditions on $(g^{F},\nabla)$ are equivalent:
  1. (1)
    $\nabla$ is $\mathrm{Aff}^{d+}(m,\mathbb{R})$-invariant and $(g^{F},\nabla)$ is conjugate symmetric.
  2. (2)
    $\nabla$ is $\mathrm{Aff}^{d+}(m,\mathbb{R})$-invariant and $(g^{F},\nabla)$ has constant curvature.
  3. (3)
    $\nabla=\nabla^{A(\alpha)}$ holds for some $\alpha\in\mathbb{R}$.
Proof.
One sees that the Lie group $G_{\mathbb{R}\mathrm{H}^{m+1}}$ is isomorphic to $\mathrm{Aff}^{d+}(m,\mathbb{R})$. Moreover, $\mathrm{Aff}^{d+}(m,\mathbb{R})$ has a simply-transitive action on $\mathcal{N}_{T}^{m}$ under which the Fisher metric $g^{F}$ and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ on $\mathcal{N}_{T}^{m}$ are invariant, see Example 3.1.2. Thus, by Theorem 6.1.1 and Proposition 2.4.5, we obtain the claim. ∎
In [17], the Amari–Chentsov $\alpha$-connection on the $m$-variate normal distribution family $(\mathcal{N}^{m},\mathbb{R}^{m},\mathbf{p}_{N})$ is characterized in the following sense:
Proposition 6.3.2 ([17]).
Let
[…]
$\mathrm{Aff}^{s}(m,\mathbb{R}):=R\ltimes\mathbb{R}^{m}$, and $\nabla$ be an statistical connection on $(\mathcal{N}^{m},g^{F})$. Then, the following two conditions on $(g^{F},\nabla)$ are equivalent:
  1. (1)
    $\nabla$ is $\mathrm{Aff}^{s}(m,\mathbb{R})$-invariant and $(g^{F},\nabla)$ is conjugate symmetric.
  2. (2)
    The affine connection $\nabla$ is the Amari–Chentsov $\alpha$-connection for some $\alpha\in\mathbb{R}$.
Theorem 6.3.1 provides a characterization similar to the above Proposition 6.3.2 for the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}_{T}^{m},\mathbb{R}^{m},\mathbf{p}_{T})$.
Let us consider the case where $m=2$. In [13], the Lie group $\mathrm{Aff}^{d+}(1,\mathbb{R})=\mathbb{R}_{>0}\ltimes\mathbb{R}$ is called the Lie group of (univariate) normal distributions. Note that $\mathrm{Aff}^{d+}(1,\mathbb{R})\cong G_{\mathbb{R}\mathrm{H}^{2}}$. By Theorem 6.1.1, for the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(\mathrm{Aff}^{d+}(1,\mathbb{R}))$, the following theorem holds.
Theorem 6.3.3.
The following is a complete set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}}(\mathrm{Aff}^{d+}(1,\mathbb{R}))$.
$S:=\left\{(a,b,c,d)\mid b>0\right\}\ \cup\ \left\{(a,0,c,d)\mid d\geq 0\right\}\subset S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}^{\ast}).$
Here, $(a,b,c,d):=ax_{1}^{3}+bx_{1}^{2}x_{2}+cx_{1}x_{2}^{2}+dx_{2}^{3}\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}^{\ast})$, and $\{x_{1},x_{2}\}$ is the dual basis of the canonical basis on $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}$ (see Equation (6.1)). We note that $(a,b,c,d)=(4\alpha,0,6\alpha,0)$ corresponds to $(g^{F},\nabla^{A(\alpha)})$.
7. The case of $H^{3}\times\mathbb{R}^{n-3}$
In this section, we consider the product Lie group of the $3$-dimensional Heisenberg group $H^{3}$ and the commutative Lie group $\mathbb{R}^{n-3}$, where $n\geq 3$. Then $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$ (see Proposition 4.1.2).
Throughout this section, we denote the Lie algebra of $H^{3}$ by $\mathfrak{h}_{3}$. Let us fix a canonical basis $\{e_{1},e_{2},e_{3}\}$ on $\mathfrak{h}_{3}$ (see [21]), that is, $\{e_{1},e_{2},e_{3}\}$ is a basis on $\mathfrak{h}_{3}$ with
$e_{1},e_{2}]=e_{3},$
and the other bracket products are trivial. Furthermore, we fix a basis $\{e_{4},\dots,e_{n}\}$ on the Lie algebra $\mathfrak{g}_{\mathbb{R}^{n-3}}$ of $\mathbb{R}^{n-3}$, and we also call the basis $\{e_{1},e_{2},e_{3},e_{4},\dots,e_{n}\}$ on $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n}}$ canonical basis. The dual basis of $\{e_{1},e_{2},e_{3},\dots,e_{n}\}$ is denoted by $\{x_{1},x_{2},x_{3},\dots,x_{n}\}$. Let us consider the inner product $\langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})$ that the canonical basis to be the orthonormal basis on $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$. It is well-known (see [27]) that the $(0,4)$-curvature tensor $R^{0}_{g}$ of $g$ on $H^{3}\times\mathbb{R}^{n-3}$ can be written as
$R^{0}_{g}=\frac{3}{4}\omega_{12}\odot\omega_{12}-\frac{1}{4}\omega_{13}\odot\omega_{13}-\frac{1}{4}\omega_{23}\odot\omega_{23}$
where we put $\omega_{ij}:=x_{i}\otimes x_{j}-x_{j}\otimes x_{i}$.
7.1. Main theorem for $H^{3}\times\mathbb{R}^{n}$
The goal of this section is to give a proof of the following theorem:
Theorem 7.1.1.
Let us define
$w_{0}:=x_{1}x_{1}+x_{2}x_{2}+x_{3}x_{3}\in S^{2}(\mathfrak{h}_{3}^{\ast})\cong\mathrm{Homog}_{2}(x_{1},x_{2},x_{3}).$
Then
$S^{3}_{\mathrm{CS}}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast},g)$ $=\{w_{0}p_{1}+p_{3}\mid p_{1}\in S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast}),~p_{3}\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})\},$
$S^{3}_{\mathrm{DF}}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast},g)$ $=\emptyset.$
Moreover, for each $C\in S^{3}_{\mathrm{CS}}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast},g)$, we have $\nabla^{g}C=0$ and $\mathrm{Sect}^{\nabla}_{g}(e_{1},e_{3})\geq 1/4>0$, where $\nabla$ is the statistical connection corresponding to $(g,C)$. Furthermore,
$\mathop{\mathcal{ML}\mathrm{Stat}}(H^{3}\times\mathbb{R}^{n-3})$ $\cong(S(O(2)\times O(1))\times O(n-3))\backslash\mathrm{Homog}_{3}(x_{1},\dots,x_{n}),$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(H^{3}\times\mathbb{R}^{n-3})$ $\cong O(n-3)\backslash(\mathrm{Homog}_{1}(x_{4},\dots,x_{n})\oplus\mathrm{Homog}_{3}(x_{4},\dots,x_{n})),$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(H^{3}\times\mathbb{R}^{n-3})$ $=\emptyset.$
Remark 7.1.2.
For the case of $H^{3}$, by Inoguchi and third author [14], there is no non-trivial left-invariant conjugate symmetric statistical structure on $H^{3}$, that is $S^{3}_{\mathrm{CS}}(\mathfrak{h}_{3}^{\ast},g)=\{0\}$ holds for any left-invariant Riemannian metric $g\in\mathop{\widetilde{\mathfrak{M}}}(H^{3})$. Moreover, as is well-known, any left-invariant Riemannian metric on $H^{3}$ is not flat (see COROLLARY 4.6 in [27]). Thus, we can see that there is no left-invariant dually flat structure on $H^{3}$. That is, $\mathcal{L}\mathrm{Stat}^{\mathrm{DF}}(H^{3})=\emptyset$.
Remark 7.1.3.
One can also see that there is no left-invariant constant curvature statistical structure on the $3$-dimensional Heisenberg group $H^{3}$. Note that a constant curvature statistical structure is conjugate symmetric. According to [27], any left-invariant Riemannian metric on a non-abelian nilpotent Lie group is not Einstein, and hence not of constant curvature.
As supplementary remarks to Theorem 7.1.1, we will also discuss the following points:
Proposition 7.1.4.
Any left-invariant conjugate symmetric statistical structure on $H^{3}\times\mathbb{R}^{n}$ does not have constant curvature.
7.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $H^{3}\times\mathbb{R}^{n-3}$
In this subsection, we give proofs of Theorem 7.1.1 and Proposition 7.1.4.
We shall define the subgroup $L^{\prime}$ of $GL(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})$ by
$L^{\prime}:=\mathrm{Aut}(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})\cap O(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}},\langle,\rangle)\subset GL(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}).$
Since $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$, by applying Proposition 4.3.1 and 4.3.2 (2), we have
$\mathop{\mathcal{ML}\mathrm{Stat}}(H^{3}\times\mathbb{R}^{n-3})$ $\cong L^{\prime}\backslash S^{3}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast}),$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(H^{3}\times\mathbb{R}^{n-3})$ $\cong L^{\prime}\backslash S^{3}_{\mathrm{CS}}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast},\langle,\rangle),$
$\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(H^{3}\times\mathbb{R}^{n-3})$ $\cong L^{\prime}\backslash S^{3}_{\mathrm{DF}}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast},\langle,\rangle).$
Therefore Theorem 7.1.1 and Proposition 7.1.4 follows from the three propositions below.
Proposition 7.2.1.
We shall identify $GL(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})$ with $GL(n,\mathbb{R})$ by considering matrices of linear transformations on $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ with respect to the canonical basis. Then the subgroup $L^{\prime}$ in $GL(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})\cong GL(n,\mathbb{R})$ can be written as
(7.1) $L^{\prime}=\left\{\begin{pmatrix}a&b&0&0&\cdots&0\\ c&d&0&0&\cdots&0\\ 0&0&ad-bc&0&\cdots&0\\ \vdots&\vdots&0&&&\\ \vdots&\vdots&\vdots&&\text{\huge{$\alpha$}}&\\ 0&0&0&&&\end{pmatrix}~\middle|~\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in O(2),\ \alpha\in O(n-3)\right\}\cong S(O(2)\times O(1))\times O(n-3).$
Furthermore, $w_{0}p_{1}\in S^{3}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast})$ is $S(O(2)\times O(1))$-invariant for any $p_{1}\in S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$.
Proposition 7.2.2.
Let $C\in S^{3}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast})\cong\mathrm{Homog}_{3}(x_{1},\dots,x_{n})$. Then $\nabla^{g}C$ is symmetric if and only if $C=w_{0}p_{1}+p_{3}$ for some $p_{1}\in S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})=\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast}$ and $p_{3}\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$. Furthermore, $\nabla^{g}(w_{0}p_{1}+p_{3})=0$.
Proposition 7.2.3.
For each
$C=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(c_{4}x_{4}+\dots c_{n}x_{n})+\sum_{4\leq i\leq j\leq k\leq n}C_{ijk}x_{i}x_{j}x_{k}\in S^{3}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast})$
where $c_{4},\dots,c_{n},C_{ijk}\in\mathbb{R}$,
$R^{C}_{g}(e_{1},e_{2},e_{2},e_{1})$ $=-\frac{3}{4}+\frac{1}{36}\sum_{k\geq 4}c_{k}^{2},\text{ and }$
$R^{C}_{g}(e_{1},e_{3},e_{3},e_{1})$ $=\frac{1}{4}+\frac{1}{36}\sum_{k\geq 4}c_{k}^{2}.$
First, we shall give a proof of Proposition 7.2.1:
Proof of Proposition 7.2.1.
Recall that the automorphism group of $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ has been known (see e.g. [18, 20, 21]). Equation (7.1) follows directly from the expressions. Note that we use the bracket relation $e_{1},e_{2}]=e_{3}$, which may have a different numbering of the basis from the previous papers. It is readily seen that the polynomial function $w_{0}=x_{1}x_{1}+x_{2}x_{2}+x_{3}x_{3}\in S^{2}(\mathfrak{h}_{3}^{\ast})$ is invariant under the action of the group $S(O(2)\times O(1))$. Therefore, the polynomial function $w_{0}p_{1}$ is also $S(O(2)\times O(1))$-invariant for any $p_{1}\in S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$. ∎
In order to prove Proposition 7.2.2, we prepare the following lemma:
Lemma 7.2.4.
The following equalities hold.
$\nabla^{g}(x_{1}^{3})$ $=-\frac{3}{2}(x_{1}^{2}x_{2})\otimes x_{3}-\frac{3}{2}(x_{1}^{2}x_{3})\otimes x_{2},$
$\nabla^{g}(x_{2}^{3})$ $=\frac{3}{2}(x_{2}^{2}x_{1})\otimes x_{3}+\frac{3}{2}(x_{2}^{2}x_{3})\otimes x_{1},$
$\nabla^{g}(x_{3}^{3})$ $=-\frac{3}{2}(x_{3}^{2}x_{2})\otimes x_{1}+\frac{3}{2}(x_{3}^{2}x_{1})\otimes x_{2},$
$\nabla^{g}(x_{1}^{2}x_{2})$ $=-(x_{1}x_{2}x_{3})\otimes x_{2}-(x_{1}x_{2}^{2})\otimes x_{3}+\frac{1}{2}(x_{1}^{2}x_{3})\otimes x_{1}+\frac{1}{2}(x_{1}^{3})\otimes x_{3},$
$\nabla^{g}(x_{1}^{2}x_{3})$ $=-(x_{1}x_{3}^{2})\otimes x_{2}-(x_{1}x_{2}x_{3})\otimes x_{3}-\frac{1}{2}(x_{1}^{2}x_{2})\otimes x_{1}+\frac{1}{2}(x_{1}^{3})\otimes x_{2},$
$\nabla^{g}(x_{2}^{2}x_{1})$ $=(x_{1}x_{2}x_{3})\otimes x_{1}+(x_{1}^{2}x_{2})\otimes x_{3}-\frac{1}{2}(x_{2}^{2}x_{3})\otimes x_{2}-\frac{1}{2}(x_{2}^{3})\otimes x_{3},$
$\nabla^{g}(x_{2}^{2}x_{3})$ $=(x_{2}x_{3}^{2})\otimes x_{1}+(x_{1}x_{2}x_{3})\otimes x_{3}-\frac{1}{2}(x_{2}^{3})\otimes x_{1}+\frac{1}{2}(x_{2}^{2}x_{1})\otimes x_{2},$
$\nabla^{g}(x_{3}^{2}x_{1})$ $=-(x_{1}x_{2}x_{3})\otimes x_{1}+(x_{1}^{2}x_{3})\otimes x_{2}-\frac{1}{2}(x_{3}^{3})\otimes x_{2}-\frac{1}{2}(x_{2}x_{3}^{2})\otimes x_{3},$
$\nabla^{g}(x_{3}^{2}x_{2})$ $=-(x_{2}^{2}x_{3})\otimes x_{1}+(x_{1}x_{2}x_{3})\otimes x_{2}+\frac{1}{2}(x_{3}^{3})\otimes x_{1}+\frac{1}{2}(x_{1}x_{3}^{2})\otimes x_{3},\text{ and }$
$\nabla^{g}(x_{1}x_{2}x_{3})$ $=-\frac{1}{2}(x_{2}^{2}x_{3})\otimes x_{3}-\frac{1}{2}(x_{2}x_{3}^{2})\otimes x_{2}+\frac{1}{2}(x_{1}^{2}x_{3})\otimes x_{3}+\frac{1}{2}(x_{1}x_{3}^{2})\otimes x_{1}$
$\ \ \ +\frac{1}{2}(x_{1}^{2}x_{2})\otimes x_{2}-\frac{1}{2}(x_{1}x_{2}^{2})\otimes x_{1}.$
Furthermore for $4\leq i,j,k\leq n$, the following equalities also hold.
$\nabla^{g}(x_{1}^{2}x_{i})$ $=-(x_{1}x_{2}x_{i})\otimes x_{3}-(x_{1}x_{3}x_{i})\otimes x_{2},$
$\nabla^{g}(x_{2}^{2}x_{i})$ $=(x_{1}x_{2}x_{i})\otimes x_{3}+(x_{2}x_{3}x_{i})\otimes x_{1},$
$\nabla^{g}(x_{3}^{2}x_{i})$ $=(x_{1}x_{3}x_{i})\otimes x_{2}-(x_{2}x_{3}x_{i})\otimes x_{1},$
$\nabla^{g}(x_{1}x_{2}x_{i})$ $=-\frac{1}{2}(x_{2}^{2}x_{i})\otimes x_{3}-\frac{1}{2}(x_{2}x_{3}x_{i})\otimes x_{2}+\frac{1}{2}(x_{1}^{2}x_{i})\otimes x_{3}+\frac{1}{2}(x_{1}x_{3}x_{i})\otimes x_{1},$
$\nabla^{g}(x_{1}x_{3}x_{i})$ $=-\frac{1}{2}(x_{2}x_{3}x_{i})\otimes x_{3}-\frac{1}{2}(x_{3}^{2}x_{i})\otimes x_{2}+\frac{1}{2}(x_{1}^{2}x_{i})\otimes x_{2}-\frac{1}{2}(x_{1}x_{2}x_{i})\otimes x_{1},$
$\nabla^{g}(x_{2}x_{3}x_{i})$ $=\frac{1}{2}(x_{1}x_{3}x_{i})\otimes x_{3}+\frac{1}{2}(x_{3}^{2}x_{i})\otimes x_{1}+\frac{1}{2}(x_{1}x_{2}x_{i})\otimes x_{2}-\frac{1}{2}(x_{2}^{2}x_{i})\otimes x_{1},$
$\nabla^{g}(x_{1}x_{i}x_{j})$ $=-\frac{1}{2}(x_{2}x_{i}x_{j})\otimes x_{3}-\frac{1}{2}(x_{3}x_{i}x_{j})\otimes x_{2},$
$\nabla^{g}(x_{2}x_{i}x_{j})$ $=\frac{1}{2}(x_{1}x_{i}x_{j})\otimes x_{3}+\frac{1}{2}(x_{3}x_{i}x_{j})\otimes x_{1},$
$\nabla^{g}(x_{3}x_{i}x_{j})$ $=-\frac{1}{2}(x_{2}x_{i}x_{j})\otimes x_{1}+\frac{1}{2}(x_{1}x_{i}x_{j})\otimes x_{2},\text{ and }$
$\nabla^{g}(x_{i}x_{j}x_{k})$ $=0.$
Proof of Lemma 7.2.4.
By definition, the structure constants of $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ with respect to the canonical basis are given by $a_{12}^{3}=1$, $a_{21}^{3}=-1$, with all other $a_{ij}^{k}$ equal to zero. Then, by applying Proposition 3.3.1, the generalized Christoffel symbols can be computed as
$\Gamma_{12}^{3}$ $=\Gamma_{23}^{1}=\Gamma_{32}^{1}=\tfrac{1}{2},$
$\Gamma_{13}^{2}$ $=\Gamma_{21}^{3}=\Gamma_{31}^{2}=-\tfrac{1}{2},$
with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 7.2.4 can be obtained by applying Proposition 3.3.2. ∎
Let us give a proof of Proposition 7.2.2.
Proof of Proposition 7.2.2.
First, let us assume that $\nabla^{g}C$ is symmetric. Put
$\nabla^{g}C=\sum_{u,v,s,t}((\nabla^{g}C)_{uvst})(x_{u}\otimes x_{v}\otimes x_{s}\otimes x_{t})$
and $C=\sum_{u,v,s}C_{uvs}(x_{u}\otimes x_{v}\otimes x_{s})$. Then by Lemma 7.2.4, the following claims hold for each $i,j\geq 4$:
  1. (1)
    $C_{1ij}=0$ since $(\nabla^{g}C)_{2ij3}=(\nabla^{g}C)_{2i3j}$, $(\nabla^{g}C)_{2ij3}=\alpha C_{1ij}$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{2i3j}=0$.
  2. (2)
    $C_{2ij}=0$ since $(\nabla^{g}C)_{1ij3}=(\nabla^{g}C)_{1i3j}$, $(\nabla^{g}C)_{1ij3}=\alpha C_{2ij}$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{1i3j}=0$.
  3. (3)
    $C_{3ij}=0$ since $(\nabla^{g}C)_{1ij2}=(\nabla^{g}C)_{1i2j}$, $(\nabla^{g}C)_{1ij2}=\alpha C_{3ij}$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{1i2j}=0$.
  4. (4)
    $C_{12i}=0$ since $(\nabla^{g}C)_{22i3}=(\nabla^{g}C)_{223i}$, $(\nabla^{g}C)_{22i3}=\alpha C_{12i}$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{223i}=0$.
  5. (5)
    $C_{13i}=0$ since $(\nabla^{g}C)_{23i3}=(\nabla^{g}C)_{233i}$, $(\nabla^{g}C)_{23i3}=\alpha C_{13i}$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{233i}=0$.
  6. (6)
    $C_{23i}=0$ since $(\nabla^{g}C)_{13i3}=(\nabla^{g}C)_{133i}$, $(\nabla^{g}C)_{13i3}=\alpha C_{23i}$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{133i}=0$.
  7. (7)
    $C_{11i}=C_{22i}$ since $(\nabla^{g}C)_{12i3}=(\nabla^{g}C)_{123i}$, $(\nabla^{g}C)_{12i3}=\alpha(C_{11i}-C_{22i})$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{123i}=0$.
  8. (8)
    $C_{11i}=C_{33i}$ since $(\nabla^{g}C)_{13i2}=(\nabla^{g}C)_{132i}$, $(\nabla^{g}C)_{13i2}=\alpha(C_{11i}-C_{33i})$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{132i}=0$.
  9. (9)
    $C_{22i}=C_{33i}$ since $(\nabla^{g}C)_{23i1}=(\nabla^{g}C)_{231i}$, $(\nabla^{g}C)_{23i1}=\alpha(C_{22i}-C_{33i})$ (for some $\alpha\neq 0$) and $(\nabla^{g}C)_{231i}=0$.
  10. (10)
    $C_{123}=0$ since $(\nabla^{g}C)_{1122}=(\nabla^{g}C)_{1221}$, $(\nabla^{g}C)_{1122}=\alpha C_{123}$ (for some $\alpha>0$) and $(\nabla^{g}C)_{1221}=\beta C_{123}$ (for some $\beta<0$).
  11. (11)
    $C_{112}=0$ since $(\nabla^{g}C)_{1131}=(\nabla^{g}C)_{1113}$, $(\nabla^{g}C)_{1131}=\frac{1}{6}C_{112}$ and $(\nabla^{g}C)_{1113}=\frac{1}{2}C_{112}$.
  12. (12)
    $C_{113}=0$ since $(\nabla^{g}C)_{1121}=(\nabla^{g}C)_{1112}$, $(\nabla^{g}C)_{1121}=-\frac{1}{6}C_{113}$ and $(\nabla^{g}C)_{1112}=\frac{1}{2}C_{113}$.
  13. (13)
    $C_{221}=0$ since $(\nabla^{g}C)_{2232}=(\nabla^{g}C)_{2223}$, $(\nabla^{g}C)_{2232}=-\frac{1}{6}C_{221}$ and $(\nabla^{g}C)_{2223}=-\frac{1}{2}C_{221}$.
  14. (14)
    $C_{223}=0$ since $(\nabla C)_{2212}=(\nabla C)_{2221}$, $(\nabla C)_{2212}=\frac{1}{6}C_{223}$ and $(\nabla C)_{2221}=-\frac{1}{2}C_{223}$.
  15. (15)
    $C_{133}=0$ since $(\nabla^{g}C)_{3323}=(\nabla^{g}C)_{3332}$, $(\nabla^{g}C)_{3323}=-\frac{1}{6}C_{133}$ and $(\nabla^{g}C)_{3332}=-\frac{1}{2}C_{133}$.
  16. (16)
    $C_{233}=0$ since $(\nabla^{g}C)_{3313}=(\nabla^{g}C)_{3331}$, $(\nabla^{g}C)_{3313}=\frac{1}{6}C_{233}$ and $(\nabla^{g}C)_{3331}=\frac{1}{2}C_{233}$.
  17. (17)
    $C_{111}=0$ since $(\nabla^{g}C)_{1123}=(\nabla^{g}C)_{1231}$, $(\nabla^{g}C)_{1123}=-\frac{1}{2}C_{111}+\frac{1}{3}C_{122}$, $(\nabla^{g}C)_{1231}=-\frac{1}{6}C_{133}$, $C_{122}=0$ and $C_{133}=0$.
  18. (18)
    $C_{222}=0$ since $(\nabla^{g}C)_{2213}=(\nabla^{g}C)_{2132}$, $(\nabla^{g}C)_{2213}=\frac{1}{2}C_{222}-\frac{1}{3}C_{112}$, $(\nabla^{g}C)_{2132}=\frac{1}{3}C_{233}$, $C_{112}=0$ and $C_{233}=0$.
  19. (19)
    $C_{333}=0$ since $(\nabla^{g}C)_{3321}=(\nabla^{g}C)_{3213}$, $(\nabla^{g}C)_{3321}=-\frac{1}{2}C_{333}+\frac{1}{3}C_{223}$, $(\nabla^{g}C)_{3213}=-\frac{1}{3}C_{113}$, $C_{113}=0$ and $C_{223}=0$.
Thus $C$ can be written as $C=w_{0}p_{1}+p_{3}$ for some $p_{1}\in S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$ and $p_{3}\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$. Conversely, let
$C=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(c_{4}x_{4}+\dots c_{n}x_{n})+\sum_{4\leq i\leq j\leq k\leq n}C_{ijk}x_{i}x_{j}x_{k}.$
Then, it immediately follows from Lemma 7.2.4 that $\nabla^{g}C=0$. ∎
Finally, we shall prove Proposition 7.2.3:
Proof of Proposition 7.2.3.
Let
$C=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(c_{4}x_{4}+\dots c_{n}x_{n})+\sum_{4\leq i\leq j\leq k\leq n}C_{ijk}x_{i}x_{j}x_{k}.$
Recall that $R^{0}_{g}=\frac{3}{4}\omega_{12}\odot\omega_{12}-\frac{1}{4}\omega_{13}\odot\omega_{13}-\frac{1}{4}\omega_{23}\odot\omega_{23}$ (see [27]). Thus by applying Proposition 2.3.2, our goal is to show that
(7.2) $[K,K](e_{1},e_{2},e_{2},e_{1})$ $=[K^{(g,C)},K^{(g,C)}](e_{1},e_{2},e_{2},e_{1})=\frac{1}{36}\sum_{k\geq 4}c_{k}^{2},$
(7.3) $[K,K](e_{1},e_{3},e_{3},e_{1})$ $=[K^{(g,C)},K^{(g,C)}](e_{1},e_{3},e_{3},e_{1})=\frac{1}{36}\sum_{k\geq 4}c_{k}^{2}$
(see Section 3.3 for the definitions of $[K,K]$ and $K^{(g,C)},K^{(g,C)}]$). Recall that
$K,K]=2\sum_{i<j,k<l,(i,j)<(k,l)}[K_{l},K_{k}]_{ij}(\omega_{ij}\odot\omega_{kl})+\sum_{i<j}[K_{j},K_{i}]_{ij}(\omega_{ij}\odot\omega_{ij})$
(cf. Theorem 3.3.3). Thus
$[K,K](e_{1},e_{2},e_{2},e_{1})=-[K,K](e_{1},e_{2},e_{1},e_{2})$ $=-[K_{2},K_{1}]_{12}=[K_{1},K_{2}]_{12},$
$[K,K](e_{1},e_{3},e_{3},e_{1})=-[K,K](e_{1},e_{3},e_{1},e_{3})$ $=-[K_{3},K_{1}]_{13}=[K_{1},K_{3}]_{13}.$
By directly computing, one can see that
$K_{l}=-\frac{1}{6}\sum_{k\geq 4}c_{k}(E_{lk}+E_{kl})$
for each $l=1,2,3$, where we write $E_{ij}$ for the $n\times n$ matrix unit whose $(i,j)$-entry is $1$ and all other entries are zero. Therefore, we obtain
$[K_{1},K_{2}]$ $=\frac{1}{36}\bigg(\sum_{k\geq 4}c_{k}^{2}\bigg)(-E_{21}+E_{12}),$
$[K_{1},K_{3}]$ $=\frac{1}{36}\bigg(\sum_{k\geq 4}c_{k}^{2}\bigg)(-E_{31}+E_{13}),$
and hence Equations (7.2) and (7.3) hold. ∎
Acknowledgements.
We would like to acknowledge Hajime Fujita, Hitoshi Furuhata, Shintaro Hashimoto, Hideyuki Ishi, Daisuke Kazukawa, Fumiyasu Komaki, Akira Kubo, Yuichiro Sato, Koichi Tojo, and Masaki Yoshioka for their valuable comments and suggestions.
Appendix A Remarks on topologies of equivariant fiber bundles
In this section, we recall a theorem (Theorem A.0.1) concerning topologies of equivariant fiber bundles, which is applied in Section 4.3.
Let $E,X,F$ be all locally-compact Hausdorff spaces, and $\pi:E\rightarrow X$ is an topological $F$-bundle. We also fix a locally-compact Hausdorff group $G$, and a continuous $G$-action on $E$ and that on $X$ such that $\pi$ is $G$-equivariant. Suppose that the $G$-action on $X$ is transitive and the map $\varpi_{x}:G\rightarrow X,~g\mapsto g.x$ is open for each $x\in X$. Let us fix $x_{0}\in X$. We write $H$ for the isotropy subgroup of $G$ at $x_{0}$, and $E_{0}:=\pi^{-1}(x_{0})$ the fiber of $\pi$ at $x_{0}$. Then $H$ acts on $E_{0}$ naturally.
[…]
$\psi:E_{0}$ $\rightarrow G\backslash E,~c\mapsto[c]_{G},$
$\phi:H\backslash E_{0}$ $\rightarrow G\backslash E,~[c]_{H}\mapsto[c]_{G},$
where $c]_{H}$ and $c]_{G}$ denotes the $H$-orbit and the $G$-orbit through $c\in E_{0}\subset E$, respectively.
Theorem A.0.1.
In the setting above, the map $\E\psi:E_{0}\rightarrow G\backslash$ E is continuous and open surjection. Furthermore, the map $\phi:H\backslash E_{0}\rightarrow G\backslash E$ is a homeomorphism.
Proof.
[…]
Let us write $\eta_{F}$ for the composition of $\eta$ and the projection from $W\times F$ onto $F$. Put $c_{F}:=\eta_{F}(c)\in F$, $c^{\prime}:=\eta(c)=(x,c_{F})\in W\times F$, $V_{x}^{\prime}:=\eta_{F}(V_{x})\subset F$, and $\Omega:=\eta(O.V_{x})\subset W\times F$. We only need to show that $c^{\prime}$ is an interior point of $\Omega$ in $W\times F$. Since $F$ is locally-compact Hausdorff and $V_{x}^{\prime}$ is an open neighborhood of $c_{F}$ in $F$, one can take a compact neighborhood $K$ of $c_{F}$ in $F$ with $K\subset V_{x}^{\prime}$. Let us consider the continuous map
$\Phi:O\times F\rightarrow F,~(g,v)\mapsto\eta_{F}(g^{-1}.\eta^{-1}(g.x,v)).$
Note that for each $v\in F$ and each $g\in O$, the equalities $\Phi(e_{G},v)=v$ and $(g.x,v)=\eta(g.\alpha)$ holds where we define $\alpha$ the unique element in $E_{x}$ with $\eta_{F}(\alpha)=\Phi(g,v)\in F$. By Lemma A.0.2 below, one can find an open neighborhood $O^{\prime}$ of $e_{G}$ in $O$ such that $\Phi(O^{\prime}\times K)\subset V_{x}^{\prime}$. For such $O^{\prime}$, one can see that
$c^{\prime}\in\varpi_{x}(O^{\prime})\times K\subset\Omega.$
Since $\varpi_{x}$ is an open map, $\varpi_{x}(O^{\prime})\times K$ is a neighborhood of $c^{\prime}$ in $W\times F$. This completes the proof. ∎
Lemma A.0.2.
Let $O$, $F$ be both topological spaces, $e\in O$, and $\Phi:O\times F\rightarrow F$ a continuous map with $\Phi(e,v)=v$ for any $v\in F$. Fix $v_{0}\in F$, a compact neighborhood $K$ of $v_{0}$ in $F$, and an open neighborhood $V$ of $v_{0}$ in $F$ with $K\subset V$. Then there exists an open neighborhood $O^{\prime}$ of $e$ in $O$ such that $\Phi(O^{\prime}\times K)\subset V$.
Appendix B Remarks on moduli of geometric structures on homogeneous spaces
Let us introduce the following notion for maps between manifolds equipped with Lie group actions:
Definition B.0.1.
Let $G_{i}$ be a Lie group acting smoothly on a smooth manifold $M_{i}$ for each $i=1,2$. We say that a pair $f:M_{1}\rightarrow M_{2}$ and $\phi:G_{1}\rightarrow G_{2}$ of smooth maps is $(G_{1},G_{2})$-equivariant if $\phi$ is a Lie group homomorphism, and for any $\eta\in G_{1}$ and any $x\in M_{1}$, the equality $f(\eta.x)=\phi(\eta).(f(x))$ holds.
Let us fix a Lie group $G$ and a closed subgroup $H$ of $G$. The homogeneous space corresponding to $(G,H)$ is denoted by $M$. In this section, we put $\mathrm{GRiem}(M)$ and $\mathrm{GStat}(M)$ to the set of all $G$-invariant Riemannian metrics on $M$ and that of all $G$-invariant statistical structures on $M$. Let us define the Lie subgroup
$\mathrm{Aut}_{H}(G):=\left\{\varphi\in\mathrm{Aut}(G)~\middle|~\varphi(H)=H\right\}$
of $\mathrm{Aut}(G)$. For each $\varphi\in\mathrm{Aut}_{H}(G)$, we define a smooth diffeomorphism $\theta_{\varphi}:M\rightarrow M$ by putting
$\theta_{\varphi}(\eta H):=\varphi(\eta)H\quad(\text{for }\eta\in G).$
Then the Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}_{H}(G)$ acts smoothly on $\mathrm{GRiem}(M)$ and $\mathrm{GStat}(M)$ as below:
$(r,\varphi).g$ $:=r\cdot((\theta_{\varphi}^{-1})^{*}g),$
$(r,\varphi).(g,\nabla)$ $:=((r,\varphi).g,(\theta_{\varphi}^{-1})^{*}\nabla).$
In the setting above, the following two theorems hold:
Theorem B.0.2.
Let $g$ and $g^{\prime}$ be both $G$-invariant Riemannian metrics on $M$. Then the following two conditions on $g$ and $g^{\prime}$ are equivalent:
  1. (i)
    The metrics $g$ and $g^{\prime}$ lies in the same $(\mathbb{R}_{>0}\times\mathrm{Aut}_{H}(G))$-orbit in $\mathrm{GRiem}(M)$.
  2. (ii)
    There exists $(G,G)$-equivariant pair $(f:M\rightarrow M,\ \phi:G\rightarrow G)$ of maps such that $\phi\in\mathrm{Aut}(G)$ and $f$ gives an isometry up to scaling between $(M,g)$ and $(M,g^{\prime})$.
Proof of Theorem B.0.2.
First, let us assume that there exists $(r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}_{H}(G)$ such that $g=(r,\varphi).g^{\prime}$ in $\mathrm{GRiem}(M)$. Then one can see that $(\theta_{\varphi}^{-1},\varphi^{-1})$ gives a $(G,G)$-equivariant pair of maps and the pair $(f,r)$ yields an isometry up to scaling between $(M,g)$ and $(M,g^{\prime})$. This proves (i) $\Rightarrow$ (ii).
Conversely, suppose (ii). Fix $(f,\phi)$ as a $(G,G)$-equivariant pair of maps such that $\phi\in\mathrm{Aut}(G)$, $f\in\mathrm{Diff}(M)$, and $r\cdot(f^{\ast}g^{\prime})=g$ ($r>0$). Let us take $\eta_{0}\in G$ with $\eta_{0}H=f(e_{G}H)$ in $M$, where $e_{G}$ denotes the unit of the group $G$. Then one sees that
$f(\eta H)=\phi(\eta)f(e_{G}H)=\phi(\eta)\eta_{0}H\ \ \text{ for any }\eta\in G.$
We define a Lie group automorphism $\varphi$ on $G$ by
$\varphi:G\rightarrow G,~\eta\mapsto\phi^{-1}(\eta_{0}\eta\eta_{0}^{-1}).$
Then $\varphi(H)=H$ since for each $h\in H$, the following equations holds in $M=G/H$:
$(\varphi^{-1}(h))H=(\eta_{0}^{-1}\phi(h)\eta_{0})H=\eta_{0}^{-1}(f(hH))=\eta_{0}^{-1}(f(e_{G}H))=\eta_{0}^{-1}(\eta_{0}H)=e_{G}H.$
Furthermore, one sees that $\theta_{\varphi}^{-1}=\rho^{M}_{\eta_{0}^{-1}}\circ f:M\rightarrow M$, where
$\rho^{M}:G\rightarrow\mathrm{Diff}(M),~\eta\mapsto\rho^{M}_{\eta}$
denotes the $G$-action on $M$. Since the metric $g^{\prime}$ is $G$-invariant, we have
$(r,\varphi).g^{\prime}=r\cdot((\theta_{\varphi}^{-1})^{\ast}g^{\prime})=r\cdot(f^{\ast}((\rho^{M}_{\eta_{0}^{-1}})^{\ast}g^{\prime}))=r\cdot(f^{\ast}g^{\prime})=g.$
This proves (i) $\Leftarrow$ (ii). ∎
Theorem B.0.3.
Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both $G$-invariant statistical structures on $M$. Then the following two conditions on $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ are equivalent:
  1. (i)
    The statistical structures $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ lies in the same $(\mathbb{R}_{>0}\times\mathrm{Aut}_{H}(G))$-orbit in $\mathrm{GStat}(M)$.
  2. (ii)
    There exists a $(G,G)$-equivariant pair $(f:M\rightarrow M,\ \phi:G\rightarrow G)$ of maps such that $\phi\in\mathrm{Aut}(G)$ and $f$ gives a scaling statistical isomorphism between $(M,g,\nabla)$ and $(M,g^{\prime},\nabla^{\prime})$.
Proof of Theorem B.0.3 can be given similarly to the proof of Theorem B.0.2.
Note that Proposition 4.2.2 is a special case of Theorem B.0.3 when $H$ is trivial and $M=G$.
[…]
Let us consider $\mathbb{R}^{n}$ the vector space equipped with the standard inner product $\langle\cdot,\cdot\rangle_{\mathbb{R}^{n}}$. The standard basis on $\mathbb{R}^{n}$ and its dual basis on $(\mathbb{R}^{n})^{\ast}$ will be denoted by $\{e_{1},\dots,e_{n}\}$ and $\{x_{1},\dots,x_{n}\}$, respectively. We denote by $S^{3}((\mathbb{R}^{n})^{\ast})$ the vector space of all symmetric $(0,3)$-tensors on $\mathbb{R}^{n}$, equipped with the $O(n)$-action defined in the usual sense. The $n$-dimensional subspace $V$ of $S^{3}((\mathbb{R}^{n})^{\ast})$ is defined by
$V:=\left\{\sum_{i}a_{i}x_{i}^{3}~\middle|~a_{i}\in\mathbb{R}\right\},$
and we write $\Xi:=O(n).V\subset S^{3}((\mathbb{R}^{n})^{\ast})$. Our concern in this section is to determine the quotient space $O(n)\backslash\Xi$.
Let us define the subspace $V_{+}$ of $V$ by
$V_{+}:=\left\{\sum_{i}\lambda_{i}x_{i}^{3}~\middle|~\lambda_{1}\geq\dots\geq\lambda_{n}\geq 0\right\}.$
The goal of this section is to give a proof of the following theorem:
Theorem C.0.1.
The map
$V_{+}\rightarrow O(n)\backslash\Xi,~\sum_{i}\lambda_{i}x_{i}^{3}\mapsto O(n).\left(\sum_{i}\lambda_{i}x_{i}\right)$
is a homeomorphism.
Remark C.0.2.
We believe that $\Xi$ is a smooth submanifold of the vector space $S^{3}((\mathbb{R}^{n})^{\ast})$ and that the natural $O(n)$-action on $\Xi$ is polar with a section $V$. If this is indeed the case, then Theorem C.0.1 would follow directly from the general theory of polar actions (cf. [11]). Although we have attempted to verify this claim, a proof has not yet been obtained.
Let us consider the standard injective group homomorphism $\mathfrak{S}_{n}\ltimes\{\pm 1\}^{n}$ into $O(n)$, and the image will be denoted by $W\subset O(n)$. Then $V_{+}$ is a complete representative of the $W$-action on $V$.
A key part of Theorem C.0.1 is the following:
Proposition C.0.3.
Let us fix $C,C^{\prime}\in V$ such that both $O(n).C=O(n).C^{\prime}$. Then $W.C=W.C^{\prime}$.
To prove Proposition C.0.3, we give two lemmas below:
Lemma C.0.4.
For each $\lambda\in\mathbb{R}^{n}$, we put
$\mathrm{supp}(\lambda):=\{i\in\{1,\dots,n\}\mid\lambda_{i}\neq 0\}$
and
$\Omega_{\lambda}:=\left\{\sum_{i}a_{i}e_{i}\in\mathbb{R}^{n}~\middle|~a_{i}\neq 0\text{ for }i\in\mathrm{supp}(\lambda),\text{and }a_{i}\lambda_{i}\neq a_{j}\lambda_{j}\text{ for }i,j\in\mathrm{supp}(\lambda)\text{ with }i\neq j\right\}.$
Then $\Omega_{\lambda}$ is open dense in $\mathbb{R}^{n}$.
Proof of Lemma C.0.4.
Recall that a finite intersection of open dense subsets of $\mathbb{R}^{n}$ is again open and dense. Since the conditions $a_{i}\neq 0$ and $a_{i}\lambda_{i}\neq a_{j}\lambda_{j}$ (for $i\neq j$) each define open dense subsets of $\mathbb{R}^{n}$, it follows that $\Omega_{\lambda}$ is open and dense in $\mathbb{R}^{n}$. ∎
Lemma C.0.5.
For each $P\in S^{3}((\mathbb{R}^{n})^{\ast})$ and $u\in\mathbb{R}^{n}$, let us write $\widetilde{K}^{P}_{u}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for the unique linear endomorphism satisfying that
$\langle\widetilde{K}^{P}_{u}(v),w\rangle_{\mathbb{R}^{n}}=\langle P,u\otimes v\otimes w\rangle\quad(\text{for any }v,w\in\mathbb{R}^{n}),$
where $\langle P,u\otimes v\otimes w\rangle$ denotes the pairing of the $(0,3)$-tensor $P$ and the $(3,0)$-tensor $u\otimes v\otimes w$ on $\mathbb{R}^{n}$. Then for each $u\in\mathbb{R}^{n}$, the equality $\widetilde{K}^{P}_{u}=k^{-1}\circ\widetilde{K}^{k.P}_{ku}\circ k$ holds as linear endomorphisms on $\mathbb{R}^{n}$.
Proof.
Let us put $Q=k.P$ and fix any $u,v,w\in\mathbb{R}^{n}$. Then
$\langle\widetilde{K}^{P}_{u}(v),w\rangle_{\mathbb{R}^{n}}$ $=\langle P,u\otimes v\otimes w\rangle$
$=\langle k^{-1}.Q,u\otimes v\otimes w\rangle$
$=\langle Q,ku\otimes kv\otimes kw\rangle$
$=\langle\widetilde{K}^{Q}_{ku}(kv),kw\rangle_{\mathbb{R}^{n}}$
$=\langle k^{-1}\widetilde{K}^{Q}_{ku}(kv),w\rangle_{\mathbb{R}^{n}}$
This proves the claim. ∎
Proof of Proposition C.0.3.
For each $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$, we write $C_{\lambda}=\sum_{i}\lambda_{i}x_{i}^{3}\in V$. Fix $\lambda,\mu\in\mathbb{R}^{n}$ and $k\in O(n)$ with $k.C_{\lambda}=C_{\mu}$. Then our goal is to find $k^{\prime}\in W$ such that $k^{\prime}.C_{\lambda}=C_{\mu}$. By Lemma C.0.4, $\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$ is open dense in $\mathbb{R}^{n}$, and in particular non-empty. Fix $u=\sum_{i}a_{i}e_{i}\in\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$. We write $ku=\sum_{i}b_{i}e_{i}\in\Omega_{\mu}$. Let us define the linear endomorphisms $\widetilde{K}^{C_{\lambda}}_{u}$ and $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathbb{R}^{n}$ as in Lemma C.0.5. Then by the definition, we have $\widetilde{K}^{C_{\lambda}}_{u}(e_{i})=a_{i}\lambda_{i}e_{i}$ and $\widetilde{K}^{C_{\mu}}_{ku}(e_{i})=b_{i}\mu_{i}e_{i}$ for each $i$. Since $u\in\Omega_{\lambda}$, we see that the family
$\{\mathbb{R}e_{i}\}_{i\in\mathrm{supp}(\lambda)}\sqcup\{\mathrm{Span}\{e_{i}\mid\lambda_{i}=0\}\}$
gives the eigenspace decomposition of $\mathbb{R}^{n}$ by the operator $\widetilde{K}^{C_{\lambda}}_{u}$ such that the eigenvalue of $\widetilde{K}^{C_{\lambda}}_{u}$ on $\mathrm{Span}\{e_{i}\mid\lambda_{i}=0\}$ is zero. Furthermore, since $ku\in\Omega_{\mu}$,
$\{\mathbb{R}e_{j}\}_{j\in\mathrm{supp}(\mu)}\sqcup\{\mathrm{Span}\{e_{j}\mid\mu_{j}=0\}\}$
gives the eigenspace decomposition of $\mathbb{R}^{n}$ by the operator $\widetilde{K}^{C_{\mu}}_{ku}$ such that the eigenvalue of $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathrm{Span}\{e_{i}\mid\mu_{i}=0\}$ is zero. By Lemma C.0.5,
$\widetilde{K}^{C_{\lambda}}_{u}=k^{-1}\circ\widetilde{K}^{C_{\mu}}_{ku}\circ k\text{ as }\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}.$
This implies the equations
$\mathrm{Span}\{e_{i}\mid\lambda_{i}=0\}=k^{-1}\mathrm{Span}\{e_{j}\mid\mu_{j}=0\}$
and
$\{\mathbb{R}e_{i}\}_{i\in\mathrm{supp}(\lambda)}=\{\mathbb{R}(k^{-1}e_{j})\}_{j\in\mathrm{supp}(\mu)}.$
In particular, one can find and fix $(\sigma,(\varepsilon_{1},\dots,\varepsilon_{n}))\in\mathfrak{S}_{n}\ltimes\{\pm 1\}^{n}$ satisfying that the equality $ke_{i}=\varepsilon_{i}e_{\sigma(i)}$ holds for each $i\in\mathrm{supp}(\lambda)$. Let us define $k^{\prime}\in W\subset O(n)$ the element corresponding to $(\sigma,(\varepsilon_{1},\dots,\varepsilon_{n}))\in\mathfrak{S}_{n}\ltimes\{\pm 1\}^{n}$ fixed as above. Then for $k_{0}:=k^{\prime-1}k\in O(n)$, we have $k_{0}e_{i}=e_{i}$ for each $i\in\mathrm{supp}(\lambda)$ and $\mathrm{Span}\{e_{i}\mid\lambda_{i}=0\}$ is stable by $k_{0}$. Hence $k_{0}.C_{\lambda}=C_{\lambda}$, and $k^{\prime}.C_{\lambda}=k.C_{\lambda}=C_{\mu}$. This completes the proof. ∎
Proof of Theorem C.0.1.
The continuous map $\psi:V_{+}\rightarrow O(n)\backslash\Xi$ is surjective because any $W$-orbit in $V$ intersects $V_{+}$ and $W\subset O(n)$. The injectivity of $\psi$ follows from Proposition C.0.3 as below: Let us take any two elements $P,Q\in V_{+}$ such that $O(n).P=\psi(P)=\psi(Q)=O(n).Q$. Then by Proposition C.0.3, $W.P=W.Q$. Since $V_{+}$ gives a complete representative of $W$-action on $V$, we have $P=Q$. This proves that the map $\psi$ is injective.
We only need to show that the continuous bijective map $\psi$ is closed. One can see that $V_{+}$ is closed in $S^{3}((\mathbb{R}^{n})^{\ast})$, and hence the inclusion map $V_{+}\rightarrow\Xi$ is closed. Furthermore, the quotient map $\Xi\rightarrow O(n)\backslash\Xi$ is also closed. In fact, in general, for a compact Hausdorff group $K$ and a continuous $K$-action on a locally compact Hausdorff space $Y$, the quotient map $\YY\rightarrow K\backslash Y$ is known to be closed. Thus the composition $\psi:V_{+}\rightarrow O(n)\backslash\Xi$ is also closed. ∎
Reference
  • [1] S. Amari and H. Nagaoka. Methods of information geometry, vol. 191. Amer. Math. Soc., 2000. https://doi.org/10.1090/mmono/191
  • [2] S. Amari. Differential-geometrical methods in statistics, vol. 28 of Lecture Notes in Statistics. Springer-Verlag, New York, 1985. https://doi.org/10.1007/978-1-4612-5056-2
  • [3] A. Andrada, M. L. Barberis, and I. Dotti. Classification of abelian complex structures on 6-dimensional Lie algebras. J. Lond. Math. Soc. (2), 83(1):232–255, 2011. https://doi.org/10.1112/jlms/jdq071
  • [4] N. Ay, J. Jost, H. V. Lê, and L. Schwachhöfer. Information geometry, vol. 64 of Ergeb. Math. Grenzgeb. (3). Springer, Cham, 2017. https://doi.org/10.1007/978-3-319-56478-4
  • [5] L. P. Castellanos Moscoso and H. Tamaru. A classification of left-invariant symplectic structures on some lie groups. Beitr. Algebra Geom., 64(2):471–491, 2023. https://doi.org/10.1007/s13366-022-00643-1
  • [6] A. J. Di Scala. Invariant metrics on the Iwasawa manifold. Q. J. Math., 64(2):555–569, 2013. https://doi.org/10.1093/qmath/has006
  • [7] Y. Fujitani. Information geometry of warped product spaces. Inf. Geom., 6:127–155, 2023. https://doi.org/10.1007/s41884-022-00091-9
  • [8] H. Furuhata. Hypersurfaces in statistical manifolds. Differential Geom. Appl., 27(3):420–429, 2009. https://doi.org/10.1016/j.difgeo.2008.10.019
  • [9] H. Furuhata, J. Inoguchi, and S.-P. Kobayashi. A characterization of the alpha-connections on the statistical manifold of normal distributions. Inf. Geom., 4(1):177–188, 2021. https://doi.org/10.1007/s41884-020-00037-z
  • [10] H. Furuhata and T. Kurose. Hessian manifolds of nonpositive constant hessian sectional curvature. Tohoku Math. J. (2), 65(1):31–42, 2013. https://doi.org/10.2748/tmj/1365452623
  • [11] C. Gorodski. Topics in polar actions. arXiv, arXiv:2208.03577, 2025. https://doi.org/10.48550/arXiv.2208.03577
  • [12] T. Hashinaga, H. Tamaru, and K. Terada. Milnor-type theorems for left-invariant Riemannian metrics on lie groups. J. Math. Soc. Japan, 68(2):669–684, 2016. https://doi.org/10.2969/jmsj/06820669
  • [13] J. Inoguchi. On the statistical Lie groups of normal distributions. Inf. Geom., 7(2):441–447, 2024. https://doi.org/10.1007/s41884-024-00148-x
  • [14] J. Inoguchi and Y. Ohno. Homogeneous statistical manifolds. Inf. Geom., 2025. https://doi.org/10.1007/s41884-025-00172-5
  • [15] S. Kobayashi and K. Nomizu. Foundations of differential geometry. Vol. II, vol. 15 of Intersci. Tracts Pure Appl. Math. Interscience Publ., Wiley, New York-London-Sydney, 1969.
  • [16] S.-P. Kobayashi and Y. Ohno. On a constant curvature statistical manifold. Inf. Geom., 5:31–46, 2022. https://doi.org/10.1007/s41884-022-00065-x
  • [17] S.-P. Kobayashi and Y. Ohno. A characterization of the alpha-connections on the statistical manifold of multivariate normal distributions. Osaka J. Math., 62(2):22, 2025. https://doi.org/10.18910/101132
  • [18] H. Kodama, A. Takahara, and H. Tamaru. The space of left-invariant metrics on a Lie group up to isometry and scaling. Manuscripta Math., 135:229–243, 2011. https://doi.org/10.1007/s00229-010-0419-4
  • [19] Y. Kondo. A classification of left-invariant pseudo-Riemannian metrics on some nilpotent Lie groups. Hiroshima Math. J., 52(3):333–356, 2022. https://doi.org/10.32917/h2021054
  • [20] Y. Kondo and H. Tamaru. A classification of left-invariant Lorentzian metrics on some nilpotent Lie groups. Tohoku Math. J. (2), 75(1):89–117, 2023. https://doi.org/10.2748/tmj.20211122
  • [21] A. Kubo, K. Onda, Y. Taketomi, and H. Tamaru. On the moduli spaces of left-invariant pseudo-Riemannian metrics on Lie groups. Hiroshima Math. J., 46(3):357–374, 2016. https://doi.org/10.32917/hmj/1487991627
  • [22] T. Kurose. Dual connections and affine geometry. Math. Z., 203:115–121, 1990. https://doi.org/10.1007/BF02570725
  • [23] J. Lauret. Degenerations of Lie algebras and geometry of Lie groups. Differential Geom. Appl., 18(2):177–194, 2003. https://doi.org/10.1016/S0926-2245(02)00146-8
  • [24] S. L. Lauritzen. Statistical manifolds. In Differential Geom. Stat. Inference, vol. 10, pp. 163–216, 1987. https://doi.org/10.1214/lnms/1215467061
  • [25] H. Matsuzoe. Statistical manifolds and affine differential geometry. In Probabilistic Approach Geom., vol. 57 of Adv. Stud. Pure Math., pp. 303–321. Math. Soc. Japan, Tokyo, 2010. https://doi.org/10.2969/aspm/05710303
  • [26] S. Mehrshad, B. Najafi, and H. Faraji. Classification of 3-dimensional left-invariant statistical Lie groups and statistical Wallach theorem. Sci. Bull. Politeh. Univ. Buchar. Ser. A Appl. Math. Phys., 86(4):113–126, 2024.
  • [27] J. Milnor. Curvatures of left invariant metrics on Lie groups. Adv. Math., 21(3):293–329, 1976. https://doi.org/10.1016/S0001-8708(76)80002-3
  • [28] K. Nomizu and T. Sasaki. Affine differential geometry: geometry of affine immersions. Cambridge Univ. Press, 1994.
  • [29] B. Opozda. Bochner’s technique for statistical structures. Ann. Global Anal. Geom., 48(4):357–395, 2015. https://doi.org/10.1007/s10455-015-9475-z
  • [30] B. Opozda. A sectional curvature for statistical structures. Linear Algebra Appl., 497:134–161, 2016. https://doi.org/10.1016/j.laa.2016.02.021
  • [31] B. Opozda. Curvature bounded conjugate symmetric statistical structures with complete metric. Ann. Global Anal. Geom., 55(4):687–702, 2019. https://doi.org/10.1007/s10455-019-09647-y
  • [32] S. M. Salamon. Complex structures on nilpotent Lie algebras. J. Pure Appl. Algebra, 157(2–3):311–333, 2001. https://doi.org/10.1016/S0022-4049(00)00033-5
  • [33] H. Shima. Homogeneous hessian manifolds. Ann. Inst. Fourier (Grenoble), 30:91–128, 1980. https://doi.org/10.5802/aif.794
  • [34] H. Shima. Hessian manifolds of constant Hessian sectional curvature. J. Math. Soc. Japan, 47(4):735–753, 1995.
  • [35] H. Shima. The Geometry of Hessian Structures. World Sci., 2007. https://doi.org/10.1142/6241
  • [36] K. Takano. Geodesics on statistical models of the multivariate normal distribution. Tensor (N.S.), 67(2):162–169, 2006.
  • [37] Y. Taketomi. A maximal element of a moduli space of Riemannian metrics. arXiv, arXiv:2210.01483, 2022. https://doi.org/10.48550/arXiv.2210.01483
  • [38] L. Todjihounde. Dualistic structures on warped product manifolds. Differential Geom. Dyn. Syst., 8:278–284, 2006.

### v3

@@
 Primary 53C30, Secondary 53B12, 22E25, 53A15
-The first author is supported by JST SPRING, Grant Number JPMJSP2132.
-The second author is supported by JST SPRING, Grant Number JPMJSP2119.
-The third author is supported by JSPS KAKENHI Grant Number JP24K06714.
-The fourth author is supported by JSPS KAKENHI Grant Numbers 23K22395, 24K21193.
-This work was partly supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165 and Osaka Metropolitan University Strategic Research Promotion Project (Development of International Research Hubs).
-This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
+The first author was supported by JST SPRING, Grant Number JPMJSP2132.
+The second author was supported by JST SPRING, Grant Number JPMJSP2119.
+The third author was supported by JSPS KAKENHI, Grant Number JP24K06714.
+The fourth author was supported by JSPS KAKENHI Grant Numbers 23K22395, 24K21193.
+This work was partially supported by the MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165 and the Osaka Metropolitan University Strategic Research Promotion Project (Development of International Research Hubs).
+This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center at Kyoto University.
 1. Introduction
@@
 We define the moduli space of left-invariant statistical structures on $G$ as the orbit space of the action, and denoted by $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$.
-We will see that two left-invariant statistical structures $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ belongs to an $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$-orbit if and only if there exists a scaling statistical isomorphism compatible with the group structure (see Definition 4.2.1) between statistical Lie groups $(G,g,\nabla)$ and $(G,g^{\prime},\nabla^{\prime})$.
+We will see that two left-invariant statistical structures $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ belongs to an $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$-orbit if and only if there exists a scaling statistical isomorphism compatible with the group structure (see Definition 4.3) between statistical Lie groups $(G,g,\nabla)$ and $(G,g^{\prime},\nabla^{\prime})$.
 In particular, the action of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ preserves the properties of statistical structures, such as dually flatness and conjugate symmetry.
@@
 [22, 17]):
-(1.1) $\mathbb{R}^{n},\quad G_{\mathbb{R}\mathrm{H}^{n}}\ (n\geq 2),\quad H^{3}\times\mathbb{R}^{n-3}\ (n\geq 3).$
+(1.1) $\mathbb{R}^{n},\;G_{\mathbb{R}\mathrm{H}^{n}}\ (n\geq 2),\;H^{3}\times\mathbb{R}^{n-3}\ (n\geq 3).$
 Note that $\mathbb{R}^{n}$ is the abelian Lie group, $G_{\mathbb{R}\mathrm{H}^{n}}$ is the Lie group associated with the real hyperbolic space $\mathbb{R}\mathrm{H}^{n}$ (the solvable part of the Iwasawa decomposition of the identity component $SO^{0}(n,1)$ of $SO(n,1)$, and acts simply-transitively on $\mathbb{R}\mathrm{H}^{n}$), and $H^{3}$ is the three dimensional Heisenberg group.
@@
 If this projection is bijective, we call it a complete set of representatives.
-The notations $V^{+}$, $v_{0}$, and $w_{0}$ appearing in Table 1 will be explained in Theorems 1.0.1, 1.0.2, and 1.0.3, respectively.
+The notations $V^{+}$, $v_{0}$, and $w_{0}$ appearing in Table 1 will be explained in Theorems 1.1, 1.2, and 1.3, respectively.
 Table 1. Sets of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ and $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$ in (1.1)
@@
 The first case is the abelian Lie group $\mathbb{R}^{n}$.
-Theorem 1.0.1.
+Theorem 1.1.
 For the abelian Lie group $\mathbb{R}^{n}$, the following holds::
-  • (1)
+  1. (1)
     All left-invariant statistical structures on $\mathbb{R}^{n}$ are conjugate symmetric.
-  • (2)
+  2. (2)
     The following is a set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(\mathbb{R}^{n})$::
@@
 The second case is the Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ of the real hyperbolic space.
-Theorem 1.0.2.
+Theorem 1.2.
 For the Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$, we denote by
 $v_{0}:=4x_{1}^{3}+6\sum_{i=2}^{n}x_{1}x_{i}^{2}\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast}).$
-  • (1)
+  1. (1)
     $\mathbb{R}v_{0}$ is a complete set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G_{\mathbb{R}\mathrm{H}^{n}})$.
@@
 This implies that there exists a one-dimensional family of left-invariant conjugate symmetric statistical structures on $G_{\mathbb{R}\mathrm{H}^{n}}$.
-  • (2)
+  2. (2)
     $\{v_{0},-v_{0}\}$ is a complete set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G_{\mathbb{R}\mathrm{H}^{n}})$.
@@
 The third case is the Lie group $H^{3}\times\mathbb{R}^{n-3}$, the direct product of the 3-dimensional Heisenberg group and an abelian Lie group. We denote by $\odot$ the symmetric tensor product.
-Theorem 1.0.3.
+Theorem 1.3.
 For the Lie group $H^{3}\times\mathbb{R}^{n-3}$, we denote by
 $w_{0}:=x_{1}x_{1}+x_{2}x_{2}+x_{3}x_{3}\in S^{2}(\mathfrak{h}_{3}^{\ast}).$
-  • (1)
+  1. (1)
     $(\mathbb{R}w_{0}\odot S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast}))\oplus S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$ is a set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(H^{3}\times\mathbb{R}^{n-3})$.
@@
 If $n\geq 4$, then the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(H^{3}\times\mathbb{R}^{n-3})$ is homeomorphic to the orbit space of $(\mathbb{R}w_{0}\odot S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast}))\oplus S^{3}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$ under the action of O(n3)O(n-3).
-  • (2)
+  2. (2)
     There do not exist any left-invariant dually flat structures on $H^{3}\times\mathbb{R}^{n-3}$.
@@
 [16]).
-Corollary 1.0.4 (see also Theorem 6.3.1).
+Corollary 1.4 (see also Theorem 6.8).
 The Amari–Chentsov $\alpha$-connection $\nabla$ on the Takano Gaussian space $(\mathcal{N}_{T},g^{F})$ is characterized by the following two properties:
-  • (1)
+  1. (1)
     The statistical structure $(g^{F},\nabla)$ is conjugate symmetric.
-  • (2)
+  2. (2)
     $\nabla$ is $\mathbb{R}_{>0}\ltimes\mathbb{R}^{n}$-invariant.
@@
 The following relation between $\nabla$ and $\overline{\nabla}$ is well-known.
-Proposition 2.1.1 ([25]).
+Proposition 2.1 ([24]).
 If two of the following four conditions are satisfied, then the remaining two will follow:
@@
 In this subsection, we recall some basic facts about statistical manifolds, see [1, 23] for details.
-In this paper, we say that a triple $(M,g,\nabla)$ is called a statistical manifold if $g$ is a Riemannian metric and $\nabla$ is an affine connection on $M$ satisfying all conditions in Proposition 2.1.1.
+In this paper, we say that a triple $(M,g,\nabla)$ is called a statistical manifold if $g$ is a Riemannian metric and $\nabla$ is an affine connection on $M$ satisfying all conditions in Proposition 2.1.
 For a statistical manifold $(M,g,\nabla)$, the pair $(g,\nabla)$ is called a statistical structure on $M$.
 In particular, $\nabla$ is called a statistical connection on the Riemannian manifold $(M,g)$.
-For each Riemannian metric $g$ on $M$, the pair $(g,\nabla^{g})$ is a statistical structure on $M$ by the definition of the Levi-Civita connection, and $(g,\nabla^{g})$ is called the trivial statistical structure (see [8]).
+For each Riemannian metric $g$ on $M$, the pair $(g,\nabla^{g})$ is a statistical structure on $M$ by the definition of the Levi-Civita connection, and $(g,\nabla^{g})$ is called the trivial statistical structure (cf.
+[8]).
 Thus, statistical manifolds can be considered as a generalization of Riemannian manifolds.
@@
 We define a certain equivalence relation between statistical manifolds as follows.
-Definition 2.2.1.
+Definition 2.2.
 Let $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ be statistical manifolds.
@@
 In particular, in the case $r=1$, the map $f$ is called a statistical isomorphism from $(M,g,\nabla)$ to $(M^{\prime},g^{\prime},\nabla^{\prime})$, and we say that $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ are statistically isomorphic to each other.
-The following proposition will be applied in Theorem 2.3.4.
-Proposition 2.2.2.
+The following proposition will be applied in Theorem 2.9.
+Proposition 2.3.
 Let $(M,g,\nabla)$ and $(M^{\prime},g^{\prime},\nabla^{\prime})$ be statistical manifolds that are statistically isomorphic up to scaling by $(f,r)$. Then the following equality holds:
@@
 In addition, we denote by $S^{3}(T^{\ast}M)$ the set of all symmetric $(0,3)$-tensor fields on $M$. The following fact is folklore known.
-Proposition 2.2.3.
+Proposition 2.4.
 The following maps are both bijective:
@@
 $\mathcal{A}_{\mathrm{Stat}}(M,g)\ni\nabla\longmapsto\nabla g\in S^{3}(T^{\ast}M).$
-By Proposition 2.2.3, a pair consisting of the Riemannian metric $g$ and a cubic form $C\in S^{3}(T^{\ast}M)$, or a pair consisting of $g$ and a difference tensor $K\in\mathcal{K}(M,g)$, can also be called a statistical structure on $M$.
+By Proposition 2.4, a pair consisting of the Riemannian metric $g$ and a cubic form $C\in S^{3}(T^{\ast}M)$, or a pair consisting of $g$ and a difference tensor $K\in\mathcal{K}(M,g)$, can also be called a statistical structure on $M$.
 Thus, we will also call $(g,C)$ and $(g,K)$ statistical structures on $M$ as well.
@@
 For a pair $X,Y$ of vector fields on $M$, the Lie bracket $K_{X},K_{Y}]$ of $K_{X}$ and $K_{Y}$ is defined as $K_{X},K_{Y}](Z)=K_{X}(K_{Y}(Z))-K_{Y}(K_{X}(Z))$.
-The following proposition, proved by Opozda [30, Corollary 3.8], will be applied in Section 5.2:
-Proposition 2.2.4.
+The following proposition, proved by Opozda [29, Corollary 3.8], will be applied in Section 5.2:
+Proposition 2.5.
 Fix a statistical structure $(g,\nabla)$ and a point $p\in M$. We denote by $K$ the difference tensor of $(g,\nabla)$. Then the following two conditions on $(g,\nabla,p)$ are equivalent:
-  1. (1)
+  1. (i)
     $K_{X},K_{Y}]_{p}=0$ for any pair $(X,Y)$ of vector fields on $M$.
-  2. (2)
+  2. (ii)
     There exists an orthonormal basis $e_{1},\dots,e_{n}$ of $(T_{p}M,g_{p})$ and $c_{1},\dots c_{n}\in\mathbb{R}$ such that
@@
 The conjugate symmetry of a statistical structure can be characterized as follows:
-Proposition 2.3.1 (Proposition 2.4 in [16]).
+Proposition 2.6 (Proposition 2.4 in [15]).
 Let $(g,\nabla)$ be a statistical structure on $M$, and let $K$ and $C$ be the difference tensor and cubic form of $(g,\nabla)$ respectively. Then, the following four conditions are equivalent:
@@
     for vector fields $X_{1},X_{2},X_{3}$ and $\sigma\in\mathfrak{S}_{3}$.
-Note that, by Proposition 2.3.1, the following set is a linear subspace of $S^{3}(T^{\ast}M)$:
-(2.3) $\left\{C\in S^{3}(T^{\ast}M)\ |\ (g,C)\text{ is conjugate symmetric}\right\}.$
+Note that, by Proposition 2.6, the following set is a linear subspace of $S^{3}(T^{\ast}M)$:
+(2.3) $\left\{C\in S^{3}(T^{\ast}M)~\middle|~(g,C)\text{ is conjugate symmetric}\right\}.$
 Throughout this paper, for a conjugate symmetric statistical manifold $(M,g,\nabla)$, the $(0,4)$-tensor field $R^{\nabla}_{g}$ is defined by putting
@@
 for vector fields $X,Y,Z,W$ on $M$. Then the following holds:
-Proposition 2.3.2 (cf. [29]).
+Proposition 2.7 (cf. [28]).
 For a conjugate symmetric statistical manifold $(M,g,\nabla)$, the equality below holds:
-(2.4) $R^{\nabla}(X,Y)=R^{\nabla^{g}}(X,Y)+[K_{X},K_{Y}]$
+$R^{\nabla}(X,Y)=R^{\nabla^{g}}(X,Y)+[K_{X},K_{Y}]$
 for all vector fields $X$, $Y$ on $M$. In particular, the $(0,4)$-tensor field $R^{\nabla}_{g}$ satisfies
@@
 (see [21]) or equivalently, $(M,g,\nabla)$ is conjugate symmetric and $\mathrm{Sect}^{\nabla}_{g}(v,w)=k$ for any $p\in M$ and any linearly independent pair $(v,w)$ of vectors in $T_{p}M$.
-Remark 2.3.3.
+Remark 2.8.
 Opozda [28] gave a definition of sectional curvatures for general statistical manifolds.
@@
 It should be emphasized that by the arguments in Section 2.2, the following holds.
-Theorem 2.3.4.
+Theorem 2.9.
 The classes of conjugate symmetric, constant curvature, dually flat, and CHC statistical manifolds are preserved under scaling statistical isomorphisms.
@@
 In this subsection, we introduce some examples of statistical manifolds.
-Example 2.4.1 ($n$-variable normal distribution family).
+Example 2.10 ($n$-variable normal distribution family).
 Let $n\in\mathbb{Z}$ with $n\geq 1$, and $\mathrm{Sym}^{+}(n,\mathbb{R})$ be the space of all positive definite symmetric matrices of order $n$.
 The $n$-variate normal distribution family is a statistical model $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$ with $\mathcal{N}^{n}=\mathrm{Sym}^{+}(n,\mathbb{R})\times\mathbb{R}^{n}$ and
-[…]
+$\mathbf{p}_{N}:\mathcal{N}^{n}\to\mathcal{P}(\mathbb{R}^{n}),~(\Sigma,\mu)\mapsto N(x\mid\Sigma,\mu)dx:=\frac{1}{\sqrt{(2\pi)^{n}\det(\Sigma)}}\exp\bigg(-\frac{1}{2}(x-\mu)^{\mathsf{T}}\Sigma^{-1}(x-\mu)\bigg)dx$
 (see [4] for the definition of statistical model).
@@
 (2.5) $C^{A(\alpha)}_{\hat{\theta}}(X,Y,Z)=\alpha\int_{x\in\mathbb{R}^{n}}(X\log N(x\mid\theta))(Y\log N(x\mid\theta))(Z\log N(x\mid\theta))N(x\mid\hat{\theta})dx\quad(\hat{\theta}\in\mathcal{N}^{n}).$
-Moreover, a statistical connection $\nabla^{A(\alpha)}$ on $(\mathcal{N}^{n},g^{F})$ correspondence to $(g^{F},C^{A(\alpha)})$ in the sense of Proposition 2.2.3 is called the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$.
+Moreover, a statistical connection $\nabla^{A(\alpha)}$ on $(\mathcal{N}^{n},g^{F})$ correspondence to $(g^{F},C^{A(\alpha)})$ in the sense of Proposition 2.4 is called the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$.
 The following fact for the statistical manifold $(\mathcal{N}^{n},g^{F},C^{A(\alpha)})$ is well-known.
-Fact 2.4.2 ([2, 24]).
+Proposition 2.11 ([2, 23]).
 For any $\alpha\in\mathbb{R}$, the statistical manifold $(\mathcal{N}^{n},g^{F},\nabla^{A(\alpha)})$ is conjugate symmetric. Moreover, it is dually flat if and only if $\alpha=\pm 1$.
-Example 2.4.3 (Takano Gaussian space).
+Example 2.12 (Takano Gaussian space).
 Let $n\in\mathbb{Z}$ with $n\geq 1$, $\mathcal{N}_{T}^{n}:=\mathbb{R}_{>0}\times\mathbb{R}^{n}$ and
-$\mathbf{p}_{T}=\mathbf{p}_{N}\circ\iota:\mathcal{N}^{n}_{T}\to\mathcal{P}(\mathbb{R}^{n}),\ \ (\Sigma,\mu)\mapsto N(x\mid\mathrm{diag}(\Sigma,\dots,\Sigma),\mu)dx.$
+$\mathbf{p}_{T}=\mathbf{p}_{N}\circ\iota:\mathcal{N}^{n}_{T}\to\mathcal{P}(\mathbb{R}^{n}),~(\Sigma,\mu)\mapsto N(x\mid\mathrm{diag}(\Sigma,\dots,\Sigma),\mu)dx.$
 Here,
-$\iota:\mathcal{N}^{n}_{T}\hookrightarrow\mathcal{N}^{n},\ \ (\Sigma,\mu)\mapsto(\mathrm{diag}(\Sigma,\dots,\Sigma),\mu).$
+$\iota:\mathcal{N}^{n}_{T}\hookrightarrow\mathcal{N}^{n},~(\Sigma,\mu)\mapsto(\mathrm{diag}(\Sigma,\dots,\Sigma),\mu).$
 Then the triple $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$ is a statistical model in the sense of [4].
-On $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$, the Fisher metric $g^{F}$, the Amari–Chentsov $\alpha$-tensor field $C^{A(\alpha)}$, and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ are defined in the similar way as in Example 2.4.1, respectively.
+On $(\mathcal{N}_{T}^{n},\mathbb{R}^{n},\mathbf{p}_{T})$, the Fisher metric $g^{F}$, the Amari–Chentsov $\alpha$-tensor field $C^{A(\alpha)}$, and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ are defined in the similar way as in Example 2.10, respectively.
 In [7], the Riemannian manifold $(\mathcal{N}_{T}^{n},g^{F})$ is called the Takano Gaussian space, and we use this terminology in this paper as well.
-Remark 2.4.4.
+Remark 2.13.
 For $n=1$, the statistical model $(\mathcal{N}_{T}^{1},\mathbb{R},\mathbf{p}_{T})$ coincides with the univariate normal distribution family $(\mathcal{N}^{1},\mathbb{R},\mathbf{p}_{\mathrm{N}})$.
 For the statistical manifolds $(\mathcal{N}^{n}_{T},g^{F},\nabla^{A(\alpha)})$, the following holds.
-Proposition 2.4.5.
+Proposition 2.14.
 For any $\alpha\in\mathbb{R}$, the statistical manifold $(\mathcal{N}^{n}_{T},g^{F},C^{A(\alpha)})$ has constant curvature (cf.
@@
 For other examples of statistical manifolds obtained from statistical models, see [1, 23] for example.
-Remark 2.4.6.
+Remark 2.15.
 The statistical structure on $\mathcal{N}^{n}$ defined above is related to statistical inference on the multivariate normal distribution family $(\mathcal{N}^{n},\mathbb{R}^{n},\mathbf{p}_{N})$ (see [1] for details).
@@
 We shall introduce some examples of homogeneous statistical manifolds obtained from statistical models. We use the notions introduced in Section 2.4.
-Example 3.1.1.
-The statistical manifold $(\mathcal{N}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.4.1) is an $\mathrm{Aff}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$.
+Example 3.1.
+The statistical manifold $(\mathcal{N}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.10) is an $\mathrm{Aff}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$.
 Note that the isotropy subgroup of the $\mathrm{Aff}(n,\mathbb{R})$-action on $\mathcal{N}^{n}$ at the point $(I_{n},0)$ is $O(n)\ltimes\{0\}$.
 Here, $I_{n}$ denotes the identity matrix of size $n$.
-Example 3.1.2.
+Example 3.2.
 Let $\mathrm{Aff}^{d+}(n,\mathbb{R})\coloneqq\mathbb{R}_{>0}\ltimes\mathbb{R}^{n}$.
-The statistical manifold $(\mathcal{N}_{T}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.4.3) is an $\mathrm{Aff}^{d+}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$.
+The statistical manifold $(\mathcal{N}_{T}^{n},g^{F},\nabla^{A(\alpha)})$ (see Example 2.12) is an $\mathrm{Aff}^{d+}(n,\mathbb{R})$-homogeneous statistical manifold for each $\alpha\in\mathbb{R}$.
 Note that the $\mathrm{Aff}^{d+}(n,\mathbb{R})$-action on $\mathcal{N}_{T}^{n}$ is simply-transitive.
@@
 For each $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$, we also define
-  • $\mathcal{LA}_{\mathrm{Stat}}(G,g):=\{\nabla\in\mathcal{A}_{\mathrm{stat}}(G,g)\mid\text{$\nabla$ is left-invariant}\}\subset\mathcal{A}_{\mathrm{stat}}(G,g)$ and
-  • $\mathcal{LK}(G,g):=\{K\in\mathcal{K}(G,g)\mid\text{$K$ is left-invariant}\}\subset\mathcal{K}(G,g)$.
+  • $\mathcal{LA}_{\mathrm{Stat}}(G,g):=\left\{\nabla\in\mathcal{A}_{\mathrm{stat}}(G,g)~\middle|~\text{$\nabla$ is left-invariant}\right\}\subset\mathcal{A}_{\mathrm{stat}}(G,g)$ and
+  • $\mathcal{LK}(G,g):=\left\{K\in\mathcal{K}(G,g)~\middle|~\text{$K$ is left-invariant}\right\}\subset\mathcal{K}(G,g)$.
 For each $(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)$, we denote by $K^{(g,\nabla)}$ the difference tensor of $(g,\nabla)$.
@@
 where $\langle,\rangle$ denotes the inner product on the vector space $\mathfrak{g}$ corresponding to the left-invariant metric $g$.
-For the sake of computations in sections later, we summarize some correspondences in the form of the following proposition, which is a reformulation of Proposition 2.2.3.
-Proposition 3.2.1.
+For the sake of computations in sections later, we summarize some correspondences in the form of the following proposition, which is a reformulation of Proposition 2.4.
+Proposition 3.3.
 Let $g$ be a left-invariant Riemannian metric on $G$, and $\langle,\rangle$ denotes the inner product on $\mathfrak{g}$ corresponding to $g$. The following maps are bijective:
@@
 Throughout this paper, we also use the following terminologies:
-$\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)$ $\coloneqq\{(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)\mid(g,\nabla)\text{ is conjugate symmetric}\},$
+$\mathcal{L}\mathrm{Stat}^{\mathrm{CS}}(G)$ $\coloneqq\left\{(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)~\middle|~(g,\nabla)\text{ is conjugate symmetric}\right\},$
 […]
 Let us fix $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$ and put:
-  • $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g):=\{C\in S^{3}(\mathfrak{g}^{\ast})\ |\ (g,C)\text{ is conjugate symmetric}\}$,
-  • $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g):=\{C\in S^{3}(\mathfrak{g}^{\ast})\ |\ (g,C)\text{ is dually flat}\}$.
+  • $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g):=\left\{C\in S^{3}(\mathfrak{g}^{\ast})~\middle|~(g,C)\text{ is conjugate symmetric}\right\}$,
+  • $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g):=\left\{C\in S^{3}(\mathfrak{g}^{\ast})~\middle|~(g,C)\text{ is dually flat}\right\}$.
 Note that $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$ is always a linear subspace of $S^{3}(\mathfrak{g}^{\ast})$ but $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g)$ is not.
@@
 The following is a formula for computing $\Gamma^{k}_{ij}$ from the structure constants.
-Proposition 3.3.1.
+Proposition 3.4.
 $\Gamma_{ij}^{k}=\frac{1}{2}(a_{ij}^{k}+a_{ki}^{j}+a_{kj}^{i})$.
@@
 The following is a formula for computing the tensor $\nabla^{g}(x_{i_{1}}x_{i_{2}}x_{i_{3}})\in S^{3}(\mathfrak{g}^{\ast})\otimes\mathfrak{g}^{\ast}$.
-Proposition 3.3.2.
+Proposition 3.5.
 Let $1\leq i_{1},i_{2},i_{3}\leq n$. Then the tensor $\nabla^{g}(x_{i_{1}}x_{i_{2}}x_{i_{3}})\in S^{3}(\mathfrak{g}^{\ast})\otimes\mathfrak{g}^{\ast}$ can be written as
@@
 $R^{C}_{g}(X,Y,Z,W)=g(R^{\nabla^{(g,C)}}(X,Y)Z,W),$
-where $\nabla^{(g,C)}$ denotes the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.2.1.
-By Proposition 2.3.2, we shall regard the left-invariant $(0,4)$-tensor field $R^{C}_{g}$ on $G$ as an element of $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$, where $\wedge^{2}\mathfrak{g}^{\ast}$ denotes the space of alternating $2$-tensors on $\mathfrak{g}^{\ast}$, and $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ the space of symmetric $2$-tensors on $(\wedge^{2}\mathfrak{g}^{\ast})$.
+where $\nabla^{(g,C)}$ denotes the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.3.
+By Proposition 2.7, we shall regard the left-invariant $(0,4)$-tensor field $R^{C}_{g}$ on $G$ as an element of $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$, where $\wedge^{2}\mathfrak{g}^{\ast}$ denotes the space of alternating $2$-tensors on $\mathfrak{g}^{\ast}$, and $(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ the space of symmetric $2$-tensors on $(\wedge^{2}\mathfrak{g}^{\ast})$.
 Let us put $\omega_{ij}:=x_{i}\otimes x_{j}-x_{j}\otimes x_{i}$ for each $i,j$.
@@
 Throughout this paper, for each $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$, we denote the $(0,4)$-tensor $R_{g}^{C}-R_{g}^{0}\in(\wedge^{2}\mathfrak{g}^{\ast})\odot(\wedge^{2}\mathfrak{g}^{\ast})$ by $K^{(g,C)},K^{(g,C)}]$, or simply $[K,K]$ when the context is clear.
-Then by Proposition 2.3.2,
+Then by Proposition 2.7,
 $K,K](X,Y,Z,W)=\langle[K_{X},K_{Y}]Z,W\rangle$
@@
 By direct calculation, we obtain the following:
-Theorem 3.3.3.
+Theorem 3.6.
 Let $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$. For each $u=1,\dots,n$, we define the matrix $K_{u}\in M(n,\mathbb{R})$ by putting
@@
 Then
-$[K^{(g,C)},K^{(g,C)}]=2\sum_{i<j,k<l,(i,j)<(k,l)}[K_{l},K_{k}]_{ij}(\omega_{ij}\odot\omega_{kl})+\sum_{i<j}[K_{j},K_{i}]_{ij}(\omega_{ij}\odot\omega_{ij}),$
+$K^{(g,C)},K^{(g,C)}]=2\sum_{i<j,k<l,(i,j)<(k,l)}[K_{l},K_{k}]_{ij}(\omega_{ij}\odot\omega_{kl})+\sum_{i<j}[K_{j},K_{i}]_{ij}(\omega_{ij}\odot\omega_{ij}),$
 where $K_{u},K_{v}]:=K_{u}K_{v}-K_{v}K_{u}$ denotes the Lie bracket of the matrices $K_{u}$ and $K_{v}$.
@@
 Throughout this paper, we say that two Riemannian manifolds $(M,g)$ and $(M^{\prime},g^{\prime})$ are isometric up to scaling if there exist a scalar $r>0$ and a diffeomorphism $f:M\to M^{\prime}$ such that $r\cdot f^{*}g^{\prime}=g$.
-As is well-known, the Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ acts smoothly on the space of left-invariant Riemannian metrics $\mathop{\widetilde{\mathfrak{M}}}(G)$ (see Section  3.2 for more details on $\mathop{\widetilde{\mathfrak{M}}}(G)$) by
-(4.1) $(r,\varphi).g\coloneqq r\cdot(\varphi^{-1})^{\ast}g=r\cdot g(\varphi^{-1}_{\ast}\ \cdot\ ,\varphi^{-1}_{\ast}\ \cdot\ )\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),\ g\in\mathop{\widetilde{\mathfrak{M}}}(G)).$
+As is well-known, the Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ acts smoothly on the space of left-invariant Riemannian metrics $\mathop{\widetilde{\mathfrak{M}}}(G)$ (see Section 3.2 for more details on $\mathop{\widetilde{\mathfrak{M}}}(G)$) by
+(4.1) $(r,\varphi).g\coloneqq r\cdot(\varphi^{-1})^{\ast}g=r\cdot g(\varphi^{-1}_{\ast}~\cdot~,\varphi^{-1}_{\ast}~\cdot~)\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),~g\in\mathop{\widetilde{\mathfrak{M}}}(G)).$
 If left-invariant Riemannian metrics $g_{1}$ and $g_{2}$ are related by the group action in (4.1), then the Riemannian manifolds $(G,g_{1})$ and $(G,g_{2})$ are isometric up to scaling, see Appendix B for details.
@@
 The Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g})$ also acts smoothly on $\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ as below:
-(4.2) $(r,\varphi).\langle,\rangle\coloneqq r\cdot\langle\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ \rangle\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g}),\ \langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})).$
+(4.2) $(r,\varphi).\langle,\rangle\coloneqq r\cdot\langle\varphi^{-1}~\cdot~,\varphi^{-1}~\cdot~\rangle\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g}),~\langle,\rangle\in\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})).$
 Note that if $G$ is connected and simply-connected, then $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$ are isomorphic to each other as Lie groups, and furthermore, the homeomorphic correspondence $\mathop{\widetilde{\mathfrak{M}}}(G)\to\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ is equivariant regarding the actions of $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathfrak{g})$.
 Thus the moduli space $\mathfrak{PM}(G)$ is homeomorphic to the quotient space $(\mathbb{R}_{>0}\times\mathrm{Aut}(\mathfrak{g}))\backslash\mathop{\widetilde{\mathfrak{M}}}(\mathfrak{g})$ in such a situation.
-Remark 4.1.1.
+Remark 4.1.
 In [17], for a Lie group $G$, the orbit space of $\mathop{\widetilde{\mathfrak{M}}}(G)$ under a certain action of the group $\mathbb{R}^{\times}\times\mathrm{Aut}(G)$ is referred to as the moduli space of left-invariant Riemannian metrics on $G$.
@@
 The connected and simply-connected Lie group with $\mathfrak{PM}(G)$ being a singleton have been classified by [22].
-Proposition 4.1.2 ([23]; see also [18]).
+Proposition 4.2 ([22]; see also [17]).
 Let $G$ be a connected and simply-connected Lie group which $\mathfrak{PM}(G)$ be a singleton. Then, $G$ is isomorphic to one of the following Lie groups:
@@
 The Lie group $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ acts smoothly on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ (see Section 3.2 for the notation of $\mathop{\mathcal{L}\mathrm{Stat}}(G)$) by
-(4.4) $(r,\varphi).(g,\nabla)\coloneqq(r\cdot(\varphi^{-1})^{\ast}g,(\varphi^{-1})^{\ast}\nabla)\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),\ (g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)).$
+(4.4) $(r,\varphi).(g,\nabla)\coloneqq(r\cdot(\varphi^{-1})^{\ast}g,(\varphi^{-1})^{\ast}\nabla)\quad((r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}(G),~(g,\nabla)\in\mathop{\mathcal{L}\mathrm{Stat}}(G)).$
 We mentioned the smooth structure on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ in Section 3.2 as the trivial vector bundle $\mathop{\mathcal{L}\mathrm{Stat}}(G)\cong\mathop{\widetilde{\mathfrak{M}}}(G)\times S^{3}(\mathfrak{g}^{\ast})$.
@@
 We shall introduce the following notion:
-Definition 4.2.1.
+Definition 4.3.
 Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both left-invariant statistical structures on $G$.
-We say that a scaling statistical isomorphism $(f,r)$ from $(G,g,\nabla)$ to $(G,g^{\prime},\nabla^{\prime})$ (see Definition 2.2.1 for the notation) is compatible with the group structure if there exists $\varphi\in\mathrm{Aut}(G)$ such that $f(gx)=\varphi(g)f(x)$ for all $g,x\in G$.
+We say that a scaling statistical isomorphism $(f,r)$ from $(G,g,\nabla)$ to $(G,g^{\prime},\nabla^{\prime})$ (see Definition 2.2 for the notation) is compatible with the group structure if there exists $\varphi\in\mathrm{Aut}(G)$ such that $f(gx)=\varphi(g)f(x)$ for all $g,x\in G$.
 Then the next proposition holds:
-Proposition 4.2.2.
+Proposition 4.4.
 Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both left-invariant statistical structures on $G$.
 Then $(g,\nabla)\approx(g^{\prime},\nabla^{\prime})$ on $\mathop{\mathcal{L}\mathrm{Stat}}(G)$ holds if and only if there exists a scaling statistical isomorphism compatible with the group structure between $(G,g,\nabla)$ and $(G,g^{\prime},\nabla^{\prime})$.
-The proof of Proposition 4.2.2 is postponed to Appendix B (as Theorem B.0.3).
-As a corollary to Proposition 4.2.2, one sees that the group action (4.4) preserves properties such as conjugate symmetry, constant curvature, dually flatness, and CHC for left-invariant statistical structures on $G$ (see Section 2.3 for the definitions of these classes).
-Definition 4.2.3.
+The proof of Proposition 4.4 is postponed to Appendix B (as Theorem B.3).
+As a corollary to Proposition 4.4, one sees that the group action (4.4) preserves properties such as conjugate symmetry, constant curvature, dually flatness, and CHC for left-invariant statistical structures on $G$ (see Section 2.3 for the definitions of these classes).
+Definition 4.5.
 […]
 It should be remarked that the moduli spaces, equipped with the quotient topology, might not be Hausdorff in general.
-In fact, the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ is not Hausdorff if $G=\mathbb{R}^{n}$ (see Remark 5.1.6).
+In fact, the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(G)$ is not Hausdorff if $G=\mathbb{R}^{n}$ (see Remark 5.6).
 We also define the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{CS}}}(G)$ of left-invariant conjugate symmetric statistical structure on $G$ and the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G)$ of left-invariant dually flat structure on $G$ by the quotient space as follows:
@@
 $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g)$] at each point $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$ (see Section 3.2 for the definitions of $S^{3}_{\mathrm{CS}}(\mathfrak{g}^{\ast},g)$ and $S^{3}_{\mathrm{DF}}(\mathfrak{g}^{\ast},g)$).
-By the general theory of the topology of equivariant fiber bundles on homogeneous spaces (cf. Theorem A.0.1 in Appendix A), we obtain the following:
-Proposition 4.3.1.
+By the general theory of the topology of equivariant fiber bundles on homogeneous spaces (cf. Theorem A.1 in Appendix A), we obtain the following:
+Proposition 4.6.
 Let $G$ be a Lie group with $\mathfrak{PM}(G)\cong\{\ast\}$ (see Section 4.1 for the details). Fix $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$. Then the following holds:
@@
     $F_{g}:S^{3}(\mathfrak{g}^{\ast})\to\mathop{\mathcal{ML}\mathrm{Stat}}(G),~C\mapsto(\mathbb{R}_{>0}\times\mathrm{Aut}(G)).(g,\nabla^{(g,C)})$
-    is continuous, open and surjective. Here, $\nabla^{(g,C)}$ is the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.2.1.
+    is continuous, open and surjective. Here, $\nabla^{(g,C)}$ is the left-invariant statistical connection corresponding to $(g,C)$ in the sense of Proposition 3.3.
   2. (2)
@@
 Here, $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$ denotes the isotropy subgroup of $\mathbb{R}_{>0}\times\mathrm{Aut}(G)$ with respect to the action (4.1) at the point $g\in\mathop{\widetilde{\mathfrak{M}}}(G)$, and the Lie group $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$ acts on $S^{3}(\mathfrak{g}^{\ast})$ as below:
-(4.6) $(r,\varphi).C\coloneqq r\cdot C(\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ ,\varphi^{-1}\ \cdot\ )\quad((r,\varphi)\in(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g},\ C\in S^{3}(\mathfrak{g}^{\ast})),$
+(4.6) $(r,\varphi).C\coloneqq r\cdot C(\varphi^{-1}~\cdot~,\varphi^{-1}~\cdot~,\varphi^{-1}~\cdot~)\quad((r,\varphi)\in(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g},~C\in S^{3}(\mathfrak{g}^{\ast})),$
 where $\varphi^{-1}\cdot$ denotes the induced action of $\varphi^{-1}\in\mathrm{Aut}(G)$ on $\mathfrak{g}$.
@@
 We also note that for the isotropy subgroup $(\mathbb{R}_{>0}\times\mathrm{Aut}(G))^{g}$, the following is known:
-Proposition 4.3.2 ([37], Lemma 4.2).
+Proposition 4.7 (cf. Lemma 4.2 in Taketomi (2022), arXiv:2210.01483).
 Assume that $G$ is connected and simply-connected (and thus $\mathrm{Aut}(G)$ can be identified with $\mathrm{Aut}(\mathfrak{g})$).
@@
 5. The case of $\mathbb{R}^{n}$
-In this section, we consider the commutative Lie group $\mathbb{R}^{n}$ with $n\geq 1$. Then $\mathfrak{PM}(\mathbb{R}^{n})\cong\{\ast\}$ (see Proposition 4.1.2).
+In this section, we consider the commutative Lie group $\mathbb{R}^{n}$ with $n\geq 1$. Then $\mathfrak{PM}(\mathbb{R}^{n})\cong\{\ast\}$ (see Proposition 4.2).
 Throughout this section, we denote the Lie algebra of $\mathbb{R}^{n}$ by $\mathfrak{g}_{\mathbb{R}^{n}}$.
@@
 The goal of this section is to give a proof of the following theorem:
-Theorem 5.1.1.
+Theorem 5.1.
 The following hold.
@@
 […]
-As a corollary to Theorem 5.1.1 (2), we have the following proposition:
-Proposition 5.1.2.
+As a corollary to Theorem 5.1 (2), we have the following proposition:
+Proposition 5.2.
 Let $g$ be a left-invariant Riemannian metric on $\mathbb{R}^{n}$ with $n\geq 2$.
@@
 Conversely, we suppose that $(g,\nabla)$ is a left-invariant dually flat structure on $\mathbb{R}^{n}$ of CHC, and let us prove that $\nabla=\nabla^{g}$.
-By Proposition 2.3.2 and $R^{\nabla}=R^{\nabla^{g}}=0$, we have $K_{X},K_{Y}]=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}$.
+By Proposition 2.7 and $R^{\nabla}=R^{\nabla^{g}}=0$, we have $K_{X},K_{Y}]=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}$.
 Here, $K$ is the difference tensor of $(g,\nabla)$.
 We also note that $(\mathbb{R}^{n},g)$ has the constant sectional curvature 0.
-Thus by [34, Corollary 2] and [10, Proposition 3.1], for each $X,Y,Z\in\mathfrak{g}_{\mathbb{R}^{n}}$, the equality below holds:
+Thus by [33, Corollary 2] and [10, Proposition 3.1], for each $X,Y,Z\in\mathfrak{g}_{\mathbb{R}^{n}}$, the equality below holds:
 (5.1) $K(K(Z,X),Y)=0\text{ for }X,Y,Z\in\mathfrak{g}_{\mathbb{R}^{n}}.$
-Take the orthonormal basis $e_{1},\dots,e_{n}$ of $\mathfrak{g}_{\mathbb{R}^{n}}$ as in Theorem 5.1.1 (2) and write $C=\sum_{i}\lambda_{i}x_{i}^{3}$.
+Take the orthonormal basis $e_{1},\dots,e_{n}$ of $\mathfrak{g}_{\mathbb{R}^{n}}$ as in Theorem 5.1 (2) and write $C=\sum_{i}\lambda_{i}x_{i}^{3}$.
 Note that $K(e_{i},e_{i})=\lambda_{i}e_{i}$ for each $i$.
@@
 Note that for each $a\in\mathbb{R}$, $at^{3}$ defines a left-invariant symmetric $(0,3)$-tensor fields on $\mathbb{R}$, and $(\mathbb{R},g^{E_{1}},at^{3})$ defines a one-dimensional dually flat statistical manifold.
-Theorem 5.1.3.
+Theorem 5.3.
 For each $(g,C)\in\mathcal{L}\mathrm{Stat}^{\mathrm{DF}}(\mathbb{R}^{n})$, there exists $(a_{1},\dots,a_{n})\in\mathbb{R}^{n}$ such that $(\mathbb{R}^{n},g,C)$ is statistically isomorphic to $\prod_{i=1}^{n}(\mathbb{R},g^{E_{1}},a_{i}t^{3})$, where $\prod_{i=1}^{n}(\mathbb{R},g^{E_{1}},a_{i}t^{3})$ denotes the direct product of the family $\{(\mathbb{R},g^{E_{1}},a_{i}t^{3})\}_{i}$ of one-dimensional dually flat statistical manifolds in the sense of Section 2.2.
@@
 The open domain $\Omega_{c}$ of the Euclidean space $\mathbb{R}^{n}$ is defined by
-$\Omega_{c}:=\{(x_{1},\dots,x_{n})\mid x_{i}>0\text{ for each }i\leq p\},$
+$\Omega_{c}:=\left\{(x_{1},\dots,x_{n})~\middle|~x_{i}>0\text{ for each }i\leq p\right\},$
 and $\nabla_{c}$ denotes the Euclidean connection on $\Omega_{c}$.
@@
 $\psi_{c_{i}}(x_{i}):=\begin{cases}\frac{2}{c_{i}}\log(\frac{c_{i}}{2}x_{i})\text{ if }i\leq p\\ x_{i}\text{ if }p<i.\end{cases}$
-Corollary 5.1.4.
+Corollary 5.4.
 For each $(g,\nabla)\in\mathcal{L}\mathrm{Stat}^{\mathrm{DF}}(\mathbb{R}^{n})$, there exists $c\in\{\lambda\in\mathbb{R}^{n}\mid\lambda_{1}\geq\dots\lambda_{n}\geq 0\}$ such that the Hessian manifold $(\mathbb{R}^{n},g,\nabla)$ is statistically isomorphic to the Hessian manifold $(\Omega_{c},g_{c},\nabla_{c})$ up to scaling.
 Proof.
-By Proposition 4.1.2, and Theorem 5.1.1, there exists an orthogonal basis $\mathcal{B}=\{e_{1},\dots,e_{n}\}$ on $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c=(c_{1},\dots,c_{n})\in\{\lambda\in\mathbb{R}^{n}\mid\lambda_{1}\geq\dots\lambda_{n}\geq 0\}$ such that by putting $C=\sum_{i}c_{i}x_{i}^{3}$, the dually flat statistical manifold $(\mathbb{R}^{n},g,\nabla)$ is statistically isomorphic to $(\mathbb{R}^{n},g^{E},C)$ up to scaling.
+By Proposition 4.2, and Theorem 5.1, there exists an orthogonal basis $\mathcal{B}=\{e_{1},\dots,e_{n}\}$ on $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c=(c_{1},\dots,c_{n})\in\{\lambda\in\mathbb{R}^{n}\mid\lambda_{1}\geq\dots\lambda_{n}\geq 0\}$ such that by putting $C=\sum_{i}c_{i}x_{i}^{3}$, the dually flat statistical manifold $(\mathbb{R}^{n},g,\nabla)$ is statistically isomorphic to $(\mathbb{R}^{n},g^{E},C)$ up to scaling.
 Then, it can be directly verified that the diffeomorphism $\phi_{c}$ gives a scaling statistical isomorphism from the dually flat statistical manifold $(\Omega_{c},g_{c},\nabla_{c})$ onto $(\mathbb{R}^{n},g^{E},C)$.
 ∎
-Remark 5.1.5.
-It should be noted that, in the context of Hessian geometry, the claim of Corollary 5.1.4 can essentially be found in [33, 35], where it appears as the classification of homogeneous Hessian domains associated with commutative normal Hessian algebras.
-Remark 5.1.6.
+Remark 5.5.
+It should be noted that, in the context of Hessian geometry, the claim of Corollary 5.4 can essentially be found in [32, 34], where it appears as the classification of homogeneous Hessian domains associated with commutative normal Hessian algebras.
+Remark 5.6.
 The moduli space $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(\mathbb{R}^{n})$ is not Hausdorff.
@@
 5.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $\mathbb{R}^{n}$
-In this subsection, we give a proof of Theorem 5.1.1.
-Proof of (1) in Theorem 5.1.1.
-By Proposition 4.1.2, our goal is to show that $S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)=S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$.
+In this subsection, we give a proof of Theorem 5.1.
+Proof of (1) in Theorem 5.1.
+By Proposition 4.2, our goal is to show that $S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)=S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$.
 Fix any $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$.
 Since $\nabla^{g^{E}}_{Y}X=0$ for any $X,Y\in\mathfrak{g}_{\mathbb{R}^{n}}^{\ast}$, we have $\nabla^{g^{E}}C=0$.
-This proves that $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)$ by Proposition 2.3.1.
+This proves that $C\in S^{3}_{\mathrm{CS}}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast},\langle,\rangle)$ by Proposition 2.6.
 ∎
-Proof of (2) in Theorem 5.1.1.
+Proof of (2) in Theorem 5.1.
 Let $C\in S^{3}(\mathfrak{g}_{\mathbb{R}^{n}}^{\ast})$.
-Throughout this proof, the difference tensor of the left-invariant statistical structure $(\langle,\rangle,C)$ is denoted by $K$, and the affine connection by $\nabla$ in the sense of Proposition 3.2.1.
+Throughout this proof, the difference tensor of the left-invariant statistical structure $(\langle,\rangle,C)$ is denoted by $K$, and the affine connection by $\nabla$ in the sense of Proposition 3.3.
 Moreover, we denote the $(1,3)$-curvature tensor of $\nabla$ by $R$, and the $(1,3)$-curvature tensor of Levi-Civita connection of $g^{E}$ by $R_{g^{E}}\equiv 0$.
 Proof of (ii) $\Rightarrow$ (i):
-Assume (ii). Then by Proposition 3.2.1, the difference tensor $K$ satisfies
+Assume (ii). Then by Proposition 3.3, the difference tensor $K$ satisfies
 (5.2) $K(e_{i},e_{j})=-2\delta_{ij}\lambda_{i}e_{i}\quad(i,j=1,\dots,n),$
-where $\delta_{ij}$ denotes the Kronecker delta.
-We aim to show that $R$ coincides with 0.
-By combining $R_{g^{E}}\equiv 0$, Theorem 5.1.1 (1) with the equality Proposition 2.3.2, we see the following.
+where $\delta_{ij}$ denotes the Kronecker delta. We aim to show that $R$ coincides with 0. By combining $R_{g^{E}}\equiv 0$, Theorem 5.1 (1) with the equality Proposition 2.7, we see the following.
 $R(e_{i},e_{j})=[K_{e_{i}},K_{e_{j}}]\quad(i,j=1,\dots,n).$
-By Equation (5.2) and Proposition 2.2.4, we see that
+By Equation (5.2) and Proposition 2.5, we see that
 $R(e_{i},e_{j})=0\quad(i,j=1,\dots,n).$
@@
 Proof of (i) $\Rightarrow$ (ii):
-Assume (i). Recall that any dually flat structure is conjugate symmetric. Thus, by Proposition 2.3.2 and $R_{g^{E}}\equiv 0$, we have the equality
+Assume (i). Recall that any dually flat structure is conjugate symmetric. Thus, by Proposition 2.7 and $R_{g^{E}}\equiv 0$, we have the equality
 $R(X,Y)=[K_{X},K_{Y}]$
@@
 holds for all $u$, $v\in\mathfrak{g}_{\mathbb{R}^{n}}$.
-By Proposition 2.2.4, there exists an orthonormal basis $e_{1},\dots,e_{n}$ of $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c_{1},\dots c_{n}\in\mathbb{R}$ such that
+By Proposition 2.5, there exists an orthonormal basis $e_{1},\dots,e_{n}$ of $(\mathfrak{g}_{\mathbb{R}^{n}},\langle,\rangle)$ and $c_{1},\dots c_{n}\in\mathbb{R}$ such that
 $K(e_{i},e_{i})=c_{i}e_{i},\quad K(e_{i},e_{j})=0$
@@
 This completes the proof.
-Proofs of (3) and (4) in Theorem 5.1.1.
+Proofs of (3) and (4) in Theorem 5.1.
 The claim
 […]
-follows from Proposition 4.3.1, Proposition 4.3.2 (1) and Theorem 5.1.1 (1). The claim
+follows from Proposition 4.6, Proposition 4.7 (1) and Theorem 5.1 (1). The claim
 […]
-is obtained by applying Proposition 4.3.1, Proposition 4.3.2 (1), Theorem 5.1.1 (2), and Theorem C.0.1 in Appendix C. ∎
+is obtained by applying Proposition 4.6, Proposition 4.7 (1), Theorem 5.1 (2), and Theorem C.1 in Appendix C. ∎
 6. The case of $G_{\mathbb{R}\mathrm{H}^{n}}$
 In this section, we consider the connected and simply-connected Lie group $G_{\mathbb{R}\mathrm{H}^{n}}$ with $n\geq 2$ (see Section 4.1 for the definition of $G_{\mathbb{R}\mathrm{H}^{n}}$).
-Then $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$ (see Proposition 4.1.2).
+Then $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$ (see Proposition 4.2).
 Throughout this section, we denote the Lie algebra of $G_{\mathbb{R}\mathrm{H}^{n}}$ by $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$.
@@
 The goal of this section is to give a proof of the following theorem:
-Theorem 6.1.1.
+Theorem 6.1.
 Let us define
@@
 $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong\{C^{\alpha}\mid\alpha\in\pm 1\}\cong\{\pm 1\}.$
-Remark 6.1.2.
-The dually flat structure $(g,C^{1})$ can be found in [8, Example 2.3], as an example of Hessian structures of CHC $4$ on hyperbolic spaces.
-Notice that its dual structure $(\langle,\rangle,C^{-1})$ on $G_{\mathbb{R}\mathrm{H}^{n}}$ is dually flat but not CHC.
-As supplementary remarks to Theorem 6.1.1, we will also discuss the following points in Section 6.3 (see Theorem 6.3.1):
-Proposition 6.1.3.
+Remark 6.2.
+The dually flat structure $(g,C^{-1})$ can be found in [8, Example 2.3], as an example of Hessian structures of CHC $4$ on hyperbolic spaces.
+Notice that its dual structure $(g,C^{1})$ on $G_{\mathbb{R}\mathrm{H}^{n}}$ is dually flat but not CHC.
+As supplementary remarks to Theorem 6.1, we will also discuss the following points in Section 6.3 (see Theorem 6.8):
+Proposition 6.3.
 For any $\alpha\in\mathbb{R}$, the left-invariant statistical structure $(g,C^{\alpha})$ has constant curvature.
@@
 6.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $G_{\mathbb{R}\mathrm{H}^{n}}$
-In this subsection, we give a proof of Theorem 6.1.1.
+In this subsection, we give a proof of Theorem 6.1.
 We shall define the subgroup $L$ of $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})$ by
 $L:=\mathrm{Aut}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cap O(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}},\langle,\rangle)\subset GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}).$
-Since $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$, by applying Proposition 4.3.1 and 4.3.2 (2), we have
+Since $\mathfrak{PM}(G_{\mathbb{R}\mathrm{H}^{n}})\cong\{\ast\}$, by applying Propositions 4.6 and 4.7 (2), we have
 $\mathop{\mathcal{ML}\mathrm{Stat}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong L\backslash S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast}),$
@@
 $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(G_{\mathbb{R}\mathrm{H}^{n}})$ $\cong L\backslash S^{3}_{\mathrm{DF}}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast},\langle,\rangle).$
-Therefore Theorem 6.1.1 follows from the three propositions below.
-Proposition 6.2.1.
+Therefore Theorem 6.1 follows from the three propositions below.
+Proposition 6.4.
 We shall identify $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})$ with $GL(n,\mathbb{R})$ by considering matrices of linear transformations on $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ with respect to the canonical basis.
@@
 Furthermore, $C^{\alpha}\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})$ is $L$-invariant for each $\alpha\in\mathbb{R}$.
-Proposition 6.2.2.
+Proposition 6.5.
 Let $C\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})\cong\mathrm{Homog}_{3}(x_{1},\dots,x_{n})$.
@@
 Furthermore, $\nabla^{g}C^{\alpha}=-6\alpha(\sum_{i\geq 2}x_{i}^{2})^{2}\in S^{4}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})$.
-Proposition 6.2.3.
+Proposition 6.6.
 $R^{C^{\alpha}}_{g}=(1-\alpha^{2})R^{0}_{g}$ for each $\alpha\in\mathbb{R}$.
-First, we shall give a proof of Proposition 6.2.1:
-Proof of Proposition 6.2.1.
+First, we shall give a proof of Proposition 6.4:
+Proof of Proposition 6.4.
 Under the identification $GL(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})\cong GL(n,\mathbb{R})$, it is well-known that the automorphism group $\mathrm{Aut}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}})$ can be written as
@@
 is also $L$-invariant. ∎
-In order to prove Proposition 6.2.2, we prepare the following lemma:
-Lemma 6.2.4.
+In order to prove Proposition 6.5, we prepare the following lemma:
+Lemma 6.7.
 Let us put $x_{ijkl}:=x_{i}\otimes x_{j}\otimes x_{k}\otimes x_{l}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 4}$ for each $1\leq i,j,k,l\leq n$. Then
@@
 hold.
-Proof of Lemma 6.2.4.
+Proof of Lemma 6.7.
 By definition, the structure constants of $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}$ with respect to the canonical basis are given by $a_{1t}^{t}=1$, $a_{t1}^{t}=-1$ ($t\geq 2$), with all other $a_{ij}^{k}$ equal to zero.
-Then, by applying Proposition 3.3.1, the generalized Christoffel symbols can be computed as
+Then, by applying Proposition 3.4, the generalized Christoffel symbols can be computed as
 $\Gamma_{t1}^{t}$ $=-1\quad(t\geq 2),$
 $\Gamma_{tt}^{1}$ $=1\quad(t\geq 2),$
-with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 6.2.4 can be obtained by applying Proposition 3.3.2. ∎
-Let us give a proof of Proposition 6.2.2.
-Proof of Proposition 6.2.2.
+with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 6.7 can be obtained by applying Proposition 3.5. ∎
+Let us give a proof of Proposition 6.5.
+Proof of Proposition 6.5.
 First, let us assume that $\nabla^{g}C$ is symmetric. Put
@@
 and $C=\sum_{u,v,s}C_{uvs}(x_{uvs})$, where we put $x_{uvs}:=x_{u}\otimes x_{v}\otimes x_{s}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 3}$ and $x_{uvst}:=x_{u}\otimes x_{v}\otimes x_{s}\otimes x_{t}\in(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{n}}^{\ast})^{\otimes 4}$.
-Then by Lemma 6.2.4,
+Then by Lemma 6.7,
 $0$ $=(\nabla^{g}C)_{i11i}-(\nabla^{g}C)_{i1i1}=-C_{111}+\frac{2}{3}C_{1ii}\quad(\text{for }i\geq 2),$
@@
 $0$ $=(\nabla^{g}C)_{ijk1}-(\nabla^{g}C)_{ij1k}=-\frac{1}{6}C_{ijk}\quad(\text{for }2\leq i<j<k).$
-Thus $C=C_{111}(x_{111}+\frac{3}{2}\sum_{i\geq 2}x_{1}x_{i}^{2})=C^{\alpha}$ for $\alpha=C_{111}/4$. Conversely, by Lemma 6.2.4, one can see that
+Thus $C=C_{111}(x_{111}+\frac{3}{2}\sum_{i\geq 2}x_{1}x_{i}^{2})=C^{\alpha}$ for $\alpha=C_{111}/4$. Conversely, by Lemma 6.7, one can see that
 $\nabla^{g}C^{\alpha}$ $=-2\alpha\sum_{i\geq 2}\sum_{t\geq 2}(x_{tiit}+x_{itit}+x_{iitt})$
@@
 and in particular, $\nabla^{g}C^{\alpha}$ is symmetric. ∎
-Finally, we shall prove Proposition 6.2.3:
-Proof of Proposition 6.2.3.
-Fix $\alpha\in\mathbb{R}$, and put $C:=C^{\alpha}$. Recall that $R^{0}_{g}=\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$. Thus by applying Proposition 2.3.2, our goal is to show that
+Finally, we shall prove Proposition 6.6:
+Proof of Proposition 6.6.
+Fix $\alpha\in\mathbb{R}$, and put $C:=C^{\alpha}$. Recall that $R^{0}_{g}=\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$. Thus by applying Proposition 2.7, our goal is to show that
 $K,K]=[K^{(g,C)},K^{(g,C)}]=(-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}$
@@
 $=16\alpha^{2}(E_{uv}-E_{vu}).$
-Therefore by Theorem 3.3.3,
-$[K,K]$ $=(-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}=(-\alpha^{2})R^{0}_{g}.$
+Therefore by Theorem 3.6,
+$K,K]=(-\alpha^{2})\sum_{1\leq i<j}\omega_{ij}\odot\omega_{ij}=(-\alpha^{2})R^{0}_{g}.$
 This completes the proof. ∎
@@
 In this subsection, we present a characterization of the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ on $(\mathcal{N}_{T}^{m},\mathbb{R}^{m},\mathbf{p}_{T})$ (see Sections 2.4 and 3.1 for notation used in this subsection).
-Theorem 6.3.1.
+Theorem 6.8.
 Let $\nabla$ be a statistical connection on the Takano Gaussian space $(\mathcal{N}_{T}^{m},g^{F})$. Then, the following three conditions on $(g^{F},\nabla)$ are equivalent:
-  1. (1)
+  1. (i)
     $\nabla$ is $\mathrm{Aff}^{d+}(m,\mathbb{R})$-invariant and $(g^{F},\nabla)$ is conjugate symmetric.
-  2. (2)
+  2. (ii)
     $\nabla$ is $\mathrm{Aff}^{d+}(m,\mathbb{R})$-invariant and $(g^{F},\nabla)$ has constant curvature.
-  3. (3)
+  3. (iii)
     $\nabla=\nabla^{A(\alpha)}$ holds for some $\alpha\in\mathbb{R}$.
@@
 One sees that the Lie group $G_{\mathbb{R}\mathrm{H}^{m+1}}$ is isomorphic to $\mathrm{Aff}^{d+}(m,\mathbb{R})$.
-Moreover, $\mathrm{Aff}^{d+}(m,\mathbb{R})$ has a simply-transitive action on $\mathcal{N}_{T}^{m}$ under which the Fisher metric $g^{F}$ and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ on $\mathcal{N}_{T}^{m}$ are invariant, see Example 3.1.2.
-Thus, by Theorem 6.1.1 and Proposition 2.4.5, we obtain the claim.
+Moreover, $\mathrm{Aff}^{d+}(m,\mathbb{R})$ has a simply-transitive action on $\mathcal{N}_{T}^{m}$ under which the Fisher metric $g^{F}$ and the Amari–Chentsov $\alpha$-connection $\nabla^{A(\alpha)}$ on $\mathcal{N}_{T}^{m}$ are invariant, see Example 3.2.
+Thus, by Theorem 6.1 and Proposition 2.14, we obtain the claim.
 ∎
 In [16], the Amari–Chentsov $\alpha$-connection on the $m$-variate normal distribution family $(\mathcal{N}^{m},\mathbb{R}^{m},\mathbf{p}_{N})$ is characterized in the following sense:
-Proposition 6.3.2 ([17]).
+Proposition 6.9 ([16]).
 Let
@@
 $\mathrm{Aff}^{s}(m,\mathbb{R}):=R\ltimes\mathbb{R}^{m}$, and $\nabla$ be an statistical connection on $(\mathcal{N}^{m},g^{F})$. Then, the following two conditions on $(g^{F},\nabla)$ are equivalent:
-  1. (1)
+  1. (i)
     $\nabla$ is $\mathrm{Aff}^{s}(m,\mathbb{R})$-invariant and $(g^{F},\nabla)$ is conjugate symmetric.
-  2. (2)
+  2. (ii)
     The affine connection $\nabla$ is the Amari–Chentsov $\alpha$-connection for some $\alpha\in\mathbb{R}$.
-Theorem 6.3.1 provides a characterization similar to the above Proposition 6.3.2 for the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}_{T}^{m},\mathbb{R}^{m},\mathbf{p}_{T})$.
+Theorem 6.8 provides a characterization similar to the above Proposition 6.9 for the Amari–Chentsov $\alpha$-connection on $(\mathcal{N}_{T}^{m},\mathbb{R}^{m},\mathbf{p}_{T})$.
 Let us consider the case where $m=2$.
@@
 Note that $\mathrm{Aff}^{d+}(1,\mathbb{R})\cong G_{\mathbb{R}\mathrm{H}^{2}}$.
-By Theorem 6.1.1, for the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(\mathrm{Aff}^{d+}(1,\mathbb{R}))$, the following theorem holds.
-Theorem 6.3.3.
+By Theorem 6.1, for the moduli space $\mathop{\mathcal{ML}\mathrm{Stat}}(\mathrm{Aff}^{d+}(1,\mathbb{R}))$, the following theorem holds.
+Theorem 6.10.
 The following is a complete set of representatives for $\mathop{\mathcal{ML}\mathrm{Stat}}(\mathrm{Aff}^{d+}(1,\mathbb{R}))$.
-$S:=\left\{(a,b,c,d)\mid b>0\right\}\ \cup\ \left\{(a,0,c,d)\mid d\geq 0\right\}\subset S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}^{\ast}).$
+$S:=\left\{(a,b,c,d)~\middle|~b>0\right\}\ \cup\ \left\{(a,0,c,d)~\middle|~d\geq 0\right\}\subset S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}^{\ast}).$
 Here, $(a,b,c,d):=ax_{1}^{3}+bx_{1}^{2}x_{2}+cx_{1}x_{2}^{2}+dx_{2}^{3}\in S^{3}(\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}^{\ast})$, and $\{x_{1},x_{2}\}$ is the dual basis of the canonical basis on $\mathfrak{g}_{\mathbb{R}\mathrm{H}^{2}}$ (see Equation (6.1)).
@@
 In this section, we consider the product Lie group of the $3$-dimensional Heisenberg group $H^{3}$ and the commutative Lie group $\mathbb{R}^{n-3}$, where $n\geq 3$.
-Then $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$ (see Proposition 4.1.2).
+Then $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$ (see Proposition 4.2).
 Throughout this section, we denote the Lie algebra of $H^{3}$ by $\mathfrak{h}_{3}$.
@@
 The goal of this section is to give a proof of the following theorem:
-Theorem 7.1.1.
+Theorem 7.1.
 Let us define
@@
 $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(H^{3}\times\mathbb{R}^{n-3})$ $=\emptyset.$
-Remark 7.1.2.
+Remark 7.2.
 For the case of $H^{3}$, by Inoguchi and third author [13], there is no non-trivial left-invariant conjugate symmetric statistical structure on $H^{3}$, that is $S^{3}_{\mathrm{CS}}(\mathfrak{h}_{3}^{\ast},g)=\{0\}$ holds for any left-invariant Riemannian metric $g\in\mathop{\widetilde{\mathfrak{M}}}(H^{3})$.
@@
 That is, $\mathcal{L}\mathrm{Stat}^{\mathrm{DF}}(H^{3})=\emptyset$.
-Remark 7.1.3.
+Remark 7.3.
 One can also see that there is no left-invariant constant curvature statistical structure on the $3$-dimensional Heisenberg group $H^{3}$.
@@
 According to [26], any left-invariant Riemannian metric on a non-abelian nilpotent Lie group is not Einstein, and hence not of constant curvature.
-As supplementary remarks to Theorem 7.1.1, we will also discuss the following points:
-Proposition 7.1.4.
+As supplementary remarks to Theorem 7.1, we will also discuss the following points:
+Proposition 7.4.
 Any left-invariant conjugate symmetric statistical structure on $H^{3}\times\mathbb{R}^{n}$ does not have constant curvature.
 7.2. Left-invariant conjugate-symmetric and dually flat statistical structures on $H^{3}\times\mathbb{R}^{n-3}$
-In this subsection, we give proofs of Theorem 7.1.1 and Proposition 7.1.4.
+In this subsection, we give proofs of Theorem 7.1 and Proposition 7.4.
 We shall define the subgroup $L^{\prime}$ of $GL(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})$ by
 $L^{\prime}:=\mathrm{Aut}(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})\cap O(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}},\langle,\rangle)\subset GL(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}).$
-Since $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$, by applying Proposition 4.3.1 and 4.3.2 (2), we have
+Since $\mathfrak{PM}(H^{3}\times\mathbb{R}^{n-3})\cong\{\ast\}$, by applying Propositions 4.6 and 4.7 (2), we have
 $\mathop{\mathcal{ML}\mathrm{Stat}}(H^{3}\times\mathbb{R}^{n-3})$ $\cong L^{\prime}\backslash S^{3}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast}),$
@@
 $\mathop{\mathcal{ML}\mathrm{Stat}^{\mathrm{DF}}}(H^{3}\times\mathbb{R}^{n-3})$ $\cong L^{\prime}\backslash S^{3}_{\mathrm{DF}}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast},\langle,\rangle).$
-Therefore Theorem 7.1.1 and Proposition 7.1.4 follows from the three propositions below.
-Proposition 7.2.1.
+Therefore Theorem 7.1 and Proposition 7.4 follows from the three propositions below.
+Proposition 7.5.
 We shall identify $GL(\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})$ with $GL(n,\mathbb{R})$ by considering matrices of linear transformations on $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ with respect to the canonical basis.
@@
 Furthermore, $w_{0}p_{1}\in S^{3}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast})$ is $S(O(2)\times O(1))$-invariant for any $p_{1}\in S^{1}(\mathfrak{g}_{\mathbb{R}^{n-3}}^{\ast})$.
-Proposition 7.2.2.
+Proposition 7.6.
 Let $C\in S^{3}((\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}})^{\ast})\cong\mathrm{Homog}_{3}(x_{1},\dots,x_{n})$.
@@
 Furthermore, $\nabla^{g}(w_{0}p_{1}+p_{3})=0$.
-Proposition 7.2.3.
+Proposition 7.7.
 For each
@@
 $R^{C}_{g}(e_{1},e_{3},e_{3},e_{1})$ $=\frac{1}{4}+\frac{1}{36}\sum_{k\geq 4}c_{k}^{2}.$
-First, we shall give a proof of Proposition 7.2.1:
-Proof of Proposition 7.2.1.
+First, we shall give a proof of Proposition 7.5:
+Proof of Proposition 7.5.
 Recall that the automorphism group of $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ has been known (see e.g.
@@
 ∎
-In order to prove Proposition 7.2.2, we prepare the following lemma:
-Lemma 7.2.4.
+In order to prove Proposition 7.6, we prepare the following lemma:
+Lemma 7.8.
 The following equalities hold.
@@
 $\nabla^{g}(x_{i}x_{j}x_{k})$ $=0.$
-Proof of Lemma 7.2.4.
+Proof of Lemma 7.8.
 By definition, the structure constants of $\mathfrak{h}_{3}\oplus\mathfrak{g}_{\mathbb{R}^{n-3}}$ with respect to the canonical basis are given by $a_{12}^{3}=1$, $a_{21}^{3}=-1$, with all other $a_{ij}^{k}$ equal to zero.
-Then, by applying Proposition 3.3.1, the generalized Christoffel symbols can be computed as
+Then, by applying Proposition 3.4, the generalized Christoffel symbols can be computed as
 $\Gamma_{12}^{3}$ $=\Gamma_{23}^{1}=\Gamma_{32}^{1}=\tfrac{1}{2},$
 $\Gamma_{13}^{2}$ $=\Gamma_{21}^{3}=\Gamma_{31}^{2}=-\tfrac{1}{2},$
-with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 7.2.4 can be obtained by applying Proposition 3.3.2. ∎
-Let us give a proof of Proposition 7.2.2.
-Proof of Proposition 7.2.2.
+with all other $\Gamma_{ij}^{k}$ equal to zero. The claims of Lemma 7.8 can be obtained by applying Proposition 3.5. ∎
+Let us give a proof of Proposition 7.6.
+Proof of Proposition 7.6.
 First, let us assume that $\nabla^{g}C$ is symmetric. Put
 $\nabla^{g}C=\sum_{u,v,s,t}((\nabla^{g}C)_{uvst})(x_{u}\otimes x_{v}\otimes x_{s}\otimes x_{t})$
-and $C=\sum_{u,v,s}C_{uvs}(x_{u}\otimes x_{v}\otimes x_{s})$. Then by Lemma 7.2.4, the following claims hold for each $i,j\geq 4$:
+and $C=\sum_{u,v,s}C_{uvs}(x_{u}\otimes x_{v}\otimes x_{s})$. Then by Lemma 7.8, the following claims hold for each $i,j\geq 4$:
   1. (1)
@@
 $C=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(c_{4}x_{4}+\dots c_{n}x_{n})+\sum_{4\leq i\leq j\leq k\leq n}C_{ijk}x_{i}x_{j}x_{k}.$
-Then, it immediately follows from Lemma 7.2.4 that $\nabla^{g}C=0$. ∎
-Finally, we shall prove Proposition 7.2.3:
-Proof of Proposition 7.2.3.
+Then, it immediately follows from Lemma 7.8 that $\nabla^{g}C=0$. ∎
+Finally, we shall prove Proposition 7.7:
+Proof of Proposition 7.7.
 Let
@@
 Recall that $R^{0}_{g}=\frac{3}{4}\omega_{12}\odot\omega_{12}-\frac{1}{4}\omega_{13}\odot\omega_{13}-\frac{1}{4}\omega_{23}\odot\omega_{23}$ (see [26]).
-Thus by applying Proposition 2.3.2, our goal is to show that
+Thus by applying Proposition 2.7, our goal is to show that
 (7.2) $[K,K](e_{1},e_{2},e_{2},e_{1})$ $=[K^{(g,C)},K^{(g,C)}](e_{1},e_{2},e_{2},e_{1})=\frac{1}{36}\sum_{k\geq 4}c_{k}^{2},$
@@
 $K,K]=2\sum_{i<j,k<l,(i,j)<(k,l)}[K_{l},K_{k}]_{ij}(\omega_{ij}\odot\omega_{kl})+\sum_{i<j}[K_{j},K_{i}]_{ij}(\omega_{ij}\odot\omega_{ij})$
-(cf. Theorem 3.3.3). Thus
+(cf. Theorem 3.6). Thus
 $[K,K](e_{1},e_{2},e_{2},e_{1})=-[K,K](e_{1},e_{2},e_{1},e_{2})$ $=-[K_{2},K_{1}]_{12}=[K_{1},K_{2}]_{12},$
@@
 Appendix A Remarks on topologies of equivariant fiber bundles
-In this section, we recall a theorem (Theorem A.0.1) concerning topologies of equivariant fiber bundles, which is applied in Section 4.3.
+In this section, we recall a theorem (Theorem A.1) concerning topologies of equivariant fiber bundles, which is applied in Section 4.3.
 Let $E,X,F$ be all locally-compact Hausdorff spaces, and $\pi:E\rightarrow X$ is an topological $F$-bundle.
@@
 where $c]_{H}$ and $c]_{G}$ denotes the $H$-orbit and the $G$-orbit through $c\in E_{0}\subset E$, respectively.
-Theorem A.0.1.
+Theorem A.1.
 In the setting above, the map $\E\psi:E_{0}\rightarrow G\backslash$ E is continuous and open surjection. Furthermore, the map $\phi:H\backslash E_{0}\rightarrow G\backslash E$ is a homeomorphism.
@@
 Note that for each $v\in F$ and each $g\in O$, the equalities $\Phi(e_{G},v)=v$ and $(g.x,v)=\eta(g.\alpha)$ holds where we define $\alpha$ the unique element in $E_{x}$ with $\eta_{F}(\alpha)=\Phi(g,v)\in F$.
-By Lemma A.0.2 below, one can find an open neighborhood $O^{\prime}$ of $e_{G}$ in $O$ such that $\Phi(O^{\prime}\times K)\subset V_{x}^{\prime}$.
+By Lemma A.2 below, one can find an open neighborhood $O^{\prime}$ of $e_{G}$ in $O$ such that $\Phi(O^{\prime}\times K)\subset V_{x}^{\prime}$.
 For such $O^{\prime}$, one can see that
@@
 Since $\varpi_{x}$ is an open map, $\varpi_{x}(O^{\prime})\times K$ is a neighborhood of $c^{\prime}$ in $W\times F$. This completes the proof. ∎
-Lemma A.0.2.
+Lemma A.2.
 Let $O$, $F$ be both topological spaces, $e\in O$, and $\Phi:O\times F\rightarrow F$ a continuous map with $\Phi(e,v)=v$ for any $v\in F$.
@@
 Let us introduce the following notion for maps between manifolds equipped with Lie group actions:
-Definition B.0.1.
+Definition B.1.
 Let $G_{i}$ be a Lie group acting smoothly on a smooth manifold $M_{i}$ for each $i=1,2$.
@@
 In the setting above, the following two theorems hold:
-Theorem B.0.2.
+Theorem B.2.
 Let $g$ and $g^{\prime}$ be both $G$-invariant Riemannian metrics on $M$. Then the following two conditions on $g$ and $g^{\prime}$ are equivalent:
@@
     There exists $(G,G)$-equivariant pair $(f:M\rightarrow M,\ \phi:G\rightarrow G)$ of maps such that $\phi\in\mathrm{Aut}(G)$ and $f$ gives an isometry up to scaling between $(M,g)$ and $(M,g^{\prime})$.
-Proof of Theorem B.0.2.
+Proof of Theorem B.2.
 First, let us assume that there exists $(r,\varphi)\in\mathbb{R}_{>0}\times\mathrm{Aut}_{H}(G)$ such that $g=(r,\varphi).g^{\prime}$ in $\mathrm{GRiem}(M)$.
@@
 Then one sees that
-$f(\eta H)=\phi(\eta)f(e_{G}H)=\phi(\eta)\eta_{0}H\ \ \text{ for any }\eta\in G.$
+$f(\eta H)=\phi(\eta)f(e_{G}H)=\phi(\eta)\eta_{0}H~\text{ for any }\eta\in G.$
 We define a Lie group automorphism $\varphi$ on $G$ by
@@
 This proves (i) $\Leftarrow$ (ii). ∎
-Theorem B.0.3.
+Theorem B.3.
 Let $(g,\nabla)$ and $(g^{\prime},\nabla^{\prime})$ be both $G$-invariant statistical structures on $M$.
@@
     There exists a $(G,G)$-equivariant pair $(f:M\rightarrow M,\ \phi:G\rightarrow G)$ of maps such that $\phi\in\mathrm{Aut}(G)$ and $f$ gives a scaling statistical isomorphism between $(M,g,\nabla)$ and $(M,g^{\prime},\nabla^{\prime})$.
-Proof of Theorem B.0.3 can be given similarly to the proof of Theorem B.0.2.
-Note that Proposition 4.2.2 is a special case of Theorem B.0.3 when $H$ is trivial and $M=G$.
+Proof of Theorem B.3 can be given similarly to the proof of Theorem B.2.
+Note that Proposition 4.4 is a special case of Theorem B.3 when $H$ is trivial and $M=G$.
 […]
@@
 The goal of this section is to give a proof of the following theorem:
-Theorem C.0.1.
+Theorem C.1.
 The map
@@
 is a homeomorphism.
-Remark C.0.2.
+Remark C.2.
 We believe that $\Xi$ is a smooth submanifold of the vector space $S^{3}((\mathbb{R}^{n})^{\ast})$ and that the natural $O(n)$-action on $\Xi$ is polar with a section $V$.
-If this is indeed the case, then Theorem C.0.1 would follow directly from the general theory of polar actions (cf.
-[11]).
+If this is indeed the case, then Theorem C.1 would follow directly from the general theory of polar actions (cf.
+Gorodski (2022), arXiv:2208.03577).
 Although we have attempted to verify this claim, a proof has not yet been obtained.
@@
 Then $V_{+}$ is a complete representative of the $W$-action on $V$.
-A key part of Theorem C.0.1 is the following:
-Proposition C.0.3.
+A key part of Theorem C.1 is the following:
+Proposition C.3.
 Let us fix $C,C^{\prime}\in V$ such that both $O(n).C=O(n).C^{\prime}$. Then $W.C=W.C^{\prime}$.
-To prove Proposition C.0.3, we give two lemmas below:
-Lemma C.0.4.
+To prove Proposition C.3, we give two lemmas below:
+Lemma C.4.
 For each $\lambda\in\mathbb{R}^{n}$, we put
-$\mathrm{supp}(\lambda):=\{i\in\{1,\dots,n\}\mid\lambda_{i}\neq 0\}$
+$\mathrm{supp}(\lambda):=\left\{i\in\{1,\dots,n\}~\middle|~\lambda_{i}\neq 0\right\}$
 and
@@
 Then $\Omega_{\lambda}$ is open dense in $\mathbb{R}^{n}$.
-Proof of Lemma C.0.4.
+Proof of Lemma C.4.
 Recall that a finite intersection of open dense subsets of $\mathbb{R}^{n}$ is again open and dense.
@@
 ∎
-Lemma C.0.5.
+Lemma C.5.
 For each $P\in S^{3}((\mathbb{R}^{n})^{\ast})$ and $u\in\mathbb{R}^{n}$, let us write $\widetilde{K}^{P}_{u}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for the unique linear endomorphism satisfying that
@@
 This proves the claim. ∎
-Proof of Proposition C.0.3.
+Proof of Proposition C.3.
 For each $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$, we write $C_{\lambda}=\sum_{i}\lambda_{i}x_{i}^{3}\in V$.
@@
 Then our goal is to find $k^{\prime}\in W$ such that $k^{\prime}.C_{\lambda}=C_{\mu}$.
-By Lemma C.0.4, $\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$ is open dense in $\mathbb{R}^{n}$, and in particular non-empty.
+By Lemma C.4, $\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$ is open dense in $\mathbb{R}^{n}$, and in particular non-empty.
 Fix $u=\sum_{i}a_{i}e_{i}\in\Omega_{\lambda}\cap k^{-1}\Omega_{\mu}$.
 We write $ku=\sum_{i}b_{i}e_{i}\in\Omega_{\mu}$.
-Let us define the linear endomorphisms $\widetilde{K}^{C_{\lambda}}_{u}$ and $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathbb{R}^{n}$ as in Lemma C.0.5.
+Let us define the linear endomorphisms $\widetilde{K}^{C_{\lambda}}_{u}$ and $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathbb{R}^{n}$ as in Lemma C.5.
 Then by the definition, we have $\widetilde{K}^{C_{\lambda}}_{u}(e_{i})=a_{i}\lambda_{i}e_{i}$ and $\widetilde{K}^{C_{\mu}}_{ku}(e_{i})=b_{i}\mu_{i}e_{i}$ for each $i$.
@@
 gives the eigenspace decomposition of $\mathbb{R}^{n}$ by the operator $\widetilde{K}^{C_{\mu}}_{ku}$ such that the eigenvalue of $\widetilde{K}^{C_{\mu}}_{ku}$ on $\mathrm{Span}\{e_{i}\mid\mu_{i}=0\}$ is zero.
-By Lemma C.0.5,
+By Lemma C.5,
 $\widetilde{K}^{C_{\lambda}}_{u}=k^{-1}\circ\widetilde{K}^{C_{\mu}}_{ku}\circ k\text{ as }\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}.$
@@
 ∎
-Proof of Theorem C.0.1.
+Proof of Theorem C.1.
 The continuous map $\psi:V_{+}\rightarrow O(n)\backslash\Xi$ is surjective because any $W$-orbit in $V$ intersects $V_{+}$ and $W\subset O(n)$.
-The injectivity of $\psi$ follows from Proposition C.0.3 as below: Let us take any two elements $P,Q\in V_{+}$ such that $O(n).P=\psi(P)=\psi(Q)=O(n).Q$.
-Then by Proposition C.0.3, $W.P=W.Q$.
+The injectivity of $\psi$ follows from Proposition C.3 as below: Let us take any two elements $P,Q\in V_{+}$ such that $O(n).P=\psi(P)=\psi(Q)=O(n).Q$.
+Then by Proposition C.3, $W.P=W.Q$.
 Since $V_{+}$ gives a complete representative of $W$-action on $V$, we have $P=Q$.
@@
   • [6] A. J. Di Scala. Invariant metrics on the Iwasawa manifold. Q. J. Math., 64(2):555–569, 2013. https://doi.org/10.1093/qmath/has006
-  • [7] Y. Fujitani. Information geometry of warped product spaces. Inf. Geom., 6:127–155, 2023. https://doi.org/10.1007/s41884-022-00091-9
+  • [7] Y. Fujitani. Information geometry of warped product spaces. Inf. Geom., 6(1):127–155, 2023. https://doi.org/10.1007/s41884-022-00091-9
   • [8] H. Furuhata. Hypersurfaces in statistical manifolds. Differential Geom. Appl., 27(3):420–429, 2009. https://doi.org/10.1016/j.difgeo.2008.10.019
@@
   • [10] H. Furuhata and T. Kurose. Hessian manifolds of nonpositive constant Hessian sectional curvature. Tohoku Math. J. (2), 65(1):31–42, 2013. https://doi.org/10.2748/tmj/1365452623
-  • [11] C. Gorodski. Topics in polar actions. arXiv, arXiv:2208.03577, 2025. https://doi.org/10.48550/arXiv.2208.03577
   • [11] T.
@@
   • [12] J. Inoguchi. On the statistical Lie groups of normal distributions. Inf. Geom., 7(2):441–447, 2024. https://doi.org/10.1007/s41884-024-00148-x
-  • [14] J. Inoguchi and Y. Ohno. Homogeneous statistical manifolds. Inf. Geom., 2025. https://doi.org/10.1007/s41884-025-00172-5
+  • [13] J. Inoguchi and Y. Ohno. Homogeneous statistical manifolds. Inf. Geom., 8(2):285–341, 2025. https://doi.org/10.1007/s41884-025-00172-5
   • [14] S. Kobayashi and K. Nomizu. Foundations of Differential Geometry. Vol. II, vol. 15 of Intersci. Tracts Pure Appl. Math. Interscience Publ., Wiley, New York-London-Sydney, 1969.
-  • [16] S.-P. Kobayashi and Y. Ohno. On a constant curvature statistical manifold. Inf. Geom., 5:31–46, 2022. https://doi.org/10.1007/s41884-022-00065-x
+  • [15] S.-P. Kobayashi and Y. Ohno. On a constant curvature statistical manifold. Inf. Geom., 5(1):31–46, 2022. https://doi.org/10.1007/s41884-022-00065-x
   • [16] S.-P.
@@
 Osaka J.
-Math., 62(2):22, 2025.
+Math., 62(2):329–349, 2025.
 https://doi.org/10.18910/101132
@@
 The space of left-invariant metrics on a Lie group up to isometry and scaling.
-Manuscripta Math., 135:229–243, 2011.
+Manuscripta Math., 135(1-2):229–243, 2011.
 https://doi.org/10.1007/s00229-010-0419-4
@@
 https://doi.org/10.32917/hmj/1487991627
-  • [22] T. Kurose. Dual connections and affine geometry. Math. Z., 203:115–121, 1990. https://doi.org/10.1007/BF02570725
+  • [21] T. Kurose. Dual connections and affine geometry. Math. Z., 203(1):115–121, 1990. https://doi.org/10.1007/BF02570725
   • [22] J. Lauret. Degenerations of Lie algebras and geometry of Lie groups. Differential Geom. Appl., 18(2):177–194, 2003. https://doi.org/10.1016/S0926-2245(02)00146-8
@@
   • [26] J. Milnor. Curvatures of left invariant metrics on Lie groups. Adv. Math., 21(3):293–329, 1976. https://doi.org/10.1016/S0001-8708(76)80002-3
-  • [28] K. Nomizu and T. Sasaki. Affine differential geometry: geometry of affine immersions. Cambridge Univ. Press, 1994.
+  • [27] K. Nomizu and T. Sasaki. Affine Differential Geometry, vol. 111 of Cambridge Tracts in Mathematics. Cambridge Univ. Press, Cambridge, 1994.
   • [28] B. Opozda. Bochner’s technique for statistical structures. Ann. Global Anal. Geom., 48(4):357–395, 2015. https://doi.org/10.1007/s10455-015-9475-z
@@
   • [30] B. Opozda. Curvature bounded conjugate symmetric statistical structures with complete metric. Ann. Global Anal. Geom., 55(4):687–702, 2019. https://doi.org/10.1007/s10455-019-09647-y
-  • [32] S. M. Salamon. Complex structures on nilpotent Lie algebras. J. Pure Appl. Algebra, 157(2–3):311–333, 2001. https://doi.org/10.1016/S0022-4049(00)00033-5
-  • [33] H. Shima. Homogeneous hessian manifolds. Ann. Inst. Fourier (Grenoble), 30:91–128, 1980. https://doi.org/10.5802/aif.794
+  • [31] S. M. Salamon. Complex structures on nilpotent Lie algebras. J. Pure Appl. Algebra, 157(2-3):311–333, 2001. https://doi.org/10.1016/S0022-4049(00)00033-5
+  • [32] H. Shima. Homogeneous Hessian manifolds. Ann. Inst. Fourier (Grenoble), 30(3):91–128, 1980. https://doi.org/10.5802/aif.794
   • [33] H. Shima. Hessian manifolds of constant Hessian sectional curvature. J. Math. Soc. Japan, 47(4):735–753, 1995.
-  • [35] H. Shima. The Geometry of Hessian Structures. World Sci., 2007. https://doi.org/10.1142/6241
+  • [34] H. Shima. The Geometry of Hessian Structures. World Sci., Singapore, 2007. https://doi.org/10.1142/6241
   • [35] K. Takano. Geodesics on statistical models of the multivariate normal distribution. Tensor (N.S.), 67(2):162–169, 2006.
-  • [37] Y. Taketomi. A maximal element of a moduli space of Riemannian metrics. arXiv, arXiv:2210.01483, 2022. https://doi.org/10.48550/arXiv.2210.01483
   • [36] L. Todjihounde. Dualistic structures on warped product manifolds. Differential Geom. Dyn. Syst., 8:278–284, 2006.